### v1

# QPanda3: A High-Performance Software-Hardware Collaborative Framework for Large-Scale Quantum-Classical Computing Integration

Abstract
QPanda3 is a high-performance quantum programming framework that enhances quantum computing efficiency through optimized circuit compilation, an advanced instruction stream format (OriginBIS), and hardware-aware execution strategies. These engineering optimizations significantly improve both processing speed and system performance, addressing key challenges in the NISQ era. A core innovation, OriginBIS, accelerates encoding speeds by up to 86.9× compared to OpenQASM 2.0, while decoding is 35.6× faster, leading to more efficient data handling, reduced memory overhead, and improved communication efficiency. This directly enhances the execution of quantum circuits, making large-scale quantum simulations more feasible. Comprehensive benchmarking demonstrates QPanda3’s superior performance: quantum circuit construction is 20.7× faster, execution speeds improve by 3.4×, and transpilation efficiency increases by 14.97× over Qiskit. Notably, in compiling a 118-qubit W-state circuit on a 2D-grid topology, QPanda3 achieves an unprecedented 869.9× speedup, underscoring its ability to handle complex quantum workloads at scale. By combining high-speed quantum processing with a modular and extensible software architecture, QPanda3 provides a practical bridge between today’s NISQ devices and future fault-tolerant quantum computing. It facilitates real-world applications in financial modeling, materials science, and combinatorial optimization, while its robust and scalable design supports industrial adoption and cloud-based deployment.
Index Terms:
Quantum Computing, Software-Hardware Collaborative, High-Performance, Quantum Circuit Compilation, Intermediate Representation
Figure 1: Comparison of average time utilized for various quantum circuit processing tasks between QPanda3 and Qiskit. The tasks include circuit compilation on different topologies (all-to-all, square, heavy-hex, and linear), circuit manipulation, and circuit construction. The results indicate that QPanda3 demonstrates significantly lower compilation time across all topologies compared to Qiskit, while the time difference in manipulation and construction tasks is relatively smaller.
I Introduction
Quantum computing, as a revolutionary computational paradigm, has garnered widespread attention from both academia and industry due to its exceptional parallel computing capabilities and potential to solve problems intractable for classical computers. In recent years, significant advancements in quantum hardware have propelled quantum computing from theoretical exploration to practical applications. However, to fully harness its potential, robust software frameworks are essential to bridge the gap between users and quantum hardware. The effectiveness of quantum computing software frameworks depends on several key factors, including user accessibility, computational efficiency, and seamless hardware integration. Against this backdrop, QPanda3 [1], a modern quantum programming framework, has undergone extensive optimization to enhance usability, quantum circuit compilation efficiency, and program transmission performance, thereby driving quantum computing toward broader applications.
Currently, quantum computing is undergoing a transition from theoretical research to engineering practice. Despite the challenges faced by existing quantum hardware, such as noise, decoherence, and limited qubit scalability, engineering optimizations are steadily improving its usability, enabling Noisy Intermediate-Scale Quantum (NISQ) devices to demonstrate computational advantages over classical computers for specific problems. This engineering evolution not only determines the practical applications of NISQ devices but also influences the feasibility of large-scale fault-tolerant quantum computing in the future. Thus, enhancing the usability, stability, and execution efficiency of quantum computing has become a central challenge in its research and application. Key approaches to addressing these challenges include compilation optimization, instruction stream design, and software-hardware co-optimization.
As a high-performance quantum programming framework, QPanda3 adopts an engineering-driven approach to enhance the practical usability of quantum computing. Its core objectives include:
  • Enhancing the computational capabilities of existing quantum devices—Through efficient compilation techniques, optimized qubit mapping strategies, and quantum gate compression algorithms, QPanda3 enables NISQ devices to execute target algorithms more efficiently, achieving superior computational performance for specific tasks.
  • Optimizing software-hardware co-design in quantum computing—Unlike classical computing, quantum computing requires software to have deep awareness of hardware characteristics. QPanda3 employs adaptive qubit mapping and hardware-aware compilation strategies to ensure optimal execution across different quantum processors, thereby improving overall computational fidelity.
  • Advancing quantum computational advantages—Although some NISQ devices have demonstrated computational superiority for certain problems, achieving large-scale quantum computing necessitates further engineering breakthroughs. QPanda3 adopts a modular architecture, allowing seamless integration with future fault-tolerant quantum computing systems, thereby reducing the transition cost from NISQ to next-generation quantum technologies.
The intermediate representation (IR) of quantum programs serves as a crucial bridge between high-level quantum programming languages and low-level hardware instructions, influencing program readability, portability, and execution efficiency. However, traditional quantum IR solutions often suffer from poor readability and low transmission efficiency, making it difficult for researchers to understand the mapping between high-level algorithms and low-level hardware execution. Furthermore, the lack of a unified IR standard among mainstream quantum programming languages hinders cross-platform algorithm benchmarking and experimental comparisons.
To address these issues, QPanda3 introduces a novel IR framework featuring OriginIR and an efficient binary instruction set OriginBIS to enhance program readability, transmission efficiency, and cross-platform compatibility. In classical-quantum hybrid computing and multi-device collaborative computing scenarios, existing IR solutions are typically optimized for specific hardware architectures, limiting their portability across different quantum platforms. By optimizing its IR design, QPanda3 enables efficient migration across diverse quantum platforms, thereby improving task allocation flexibility and execution efficiency.
Efficient quantum circuit compilation is crucial for transforming abstract quantum algorithms into hardware-executable instructions, encompassing quantum gate decomposition, qubit mapping, and circuit optimization. Building upon its predecessor, QPanda2, QPanda3 introduces multiple optimizations that significantly enhance compilation efficiency. Benchmarking results indicate that QPanda3 surpasses QPanda2 and other mainstream quantum computing frameworks in circuit compilation speed, dramatically improving quantum algorithm execution efficiency. In the NISQ era, where hardware constraints are prevalent, optimizing compilation efficiency is critical for maximizing computational utility.
As a quantum programming framework focused on software-hardware co-optimization, QPanda3 excels in cloud-based quantum computing environments. Given the scarcity and high cost of quantum computing resources, centralized cloud deployment is often adopted to enhance accessibility and resource utilization. Additionally, cloud-based quantum computing facilitates the establishment of standardized software-hardware environments, allowing users to conduct experiments and development without direct access to quantum hardware. Moreover, the cloud computing model fosters ecosystem collaboration in quantum computing, simplifying resource sharing and algorithm optimization.
To achieve optimal quantum execution, a comprehensive understanding of quantum hardware characteristics—such as qubit connectivity, error rates, and coherence times—is essential. QPanda3 integrates device-aware compilation techniques to minimize circuit depth and gate operations, thereby improving computational efficiency. Furthermore, QPanda3 provides device-specific quantum program analysis tools to help users assess execution performance, identify potential bottlenecks, and further optimize quantum programs. Through software-hardware co-optimization, QPanda3 enables quantum programs to adapt to different hardware architectures, achieving more efficient computation.
Experimental results demonstrate that QPanda3 significantly outperforms existing quantum programming frameworks across multiple performance metrics. For example:
  • Compilation efficiency: Experiments on Benchpress[2], in a fully connected topology, QPanda3’s compilation speed is, on average, 9.46× faster than Qiskit[3], with peak acceleration reaching 77.76×. In a square topology, its average compilation speed surpasses Qiskit[3] by 24.6×, with a peak of 869.9×. For heavy-hexagon topologies, QPanda3 achieves an average 15.1× speedup, with a peak of 332.0×. In linear topology, the average speedup is 10.71×, with peak acceleration reaching 868.7×.
  • IR transmission efficiency: The encoding speed of OriginBIS is 18.7× faster than OriginIR and 86.9× faster than OpenQASM 2.0 [4]. The decoding speed is 18.9× faster than OriginIR and 35.6× faster than OpenQASM 2.0.
The content of this paper is organized as follows: Firstly, we review the related work on quantum computing compilation frameworks. Secondly, we provide an overview of the overall architecture and core features of QPanda3. Then, we successively introduce the representation method and transmission mechanism of quantum programs, device-based quantum circuit optimization and compilation strategies, and device-based quantum program analysis, delving into the three key components of QPanda3. Finally, we validate the performance advantages of QPanda3 through experiments.
II Related Work
II-A Intermediate Representation of Quantum Programs
The Intermediate Representation (IR) of quantum programs serves as a crucial bridge connecting high-level quantum algorithm descriptions with underlying hardware implementations in quantum computing. IR provides descriptions of quantum programs at various levels of abstraction, each suited for the design, processing, transmission, and execution of quantum programs. Lower-level abstractions tend to be closer to the machine instruction format of quantum computing devices, while higher-level abstractions are more easily understood by humans. Representing quantum programs based on the quantum circuit model is a common approach, and most IRs of quantum programs are built upon this foundation. Even for IRs of quantum programs that include classical instructions and complex control flows, they are typically extensions based on the quantum circuit model.
Quantum Assembly Language (QASM)[5] is an important form of intermediate representation for quantum programs, primarily used to provide a low-level abstraction of quantum programs. It is important to note that QASM for quantum computing differs significantly from assembly languages in classical computer architectures. Although QASM is a low-level representation, it still requires the device to dynamically manage quantum registers and classical registers (logical-to-physical mapping) rather than directly operating on hardware registers.
Intermediate representations of quantum programs that resemble high-level programming languages offer good user readability. However, such representations tend to be relatively complex. For example, with the increasing complexity of quantum computing systems and application demands, IBM has continuously extended QASM to develop OpenQASM 2.0 and OpenQASM 3.0. As widely adopted versions, OpenQASM 2.0[4] and OpenQASM 3.0[6] not only extend quantum gate operations and qubit management but also introduce more complex syntactic structures and functional features, such as classical control flow and modular programming support. They are more akin to the style of the classical programming language C compared to the original QASM. F-QASM[7] extends QASM with feedback instructions, enhancing the efficiency of implementing measurement-based branch and loop statements. Similar intermediate representations include Scaffold[8], QCL[9], Quipper[10], ProjectQ[11], and QIR[12]. These types of intermediate representations for quantum programs sacrifice portability due to their complexity, hindering comparative experimental research by researchers.
Device-oriented intermediate representations of quantum programs exhibit strong dependencies on the specific device. OpenPulse[13], Pulser[14], and JaqalPaw[15] are examples of intermediate representations tailored for pulse-based devices, offering fine-grained control over quantum devices. QuMIs[16] is an intermediate representation designed for distributed device control. eQASM[17], on the other hand, is a low-level intermediate representation that is directly linked to binary machine instructions. Clearly, these intermediate representations contain a significant amount of device-specific information within the code describing quantum programs, which provides precise control over the execution details of quantum programs but also reduces their device independence.
The intermediate representation of quantum programs can also serve as a transmission protocol for transferring quantum programs between devices. Currently, there is relatively little research discussing this topic. It is important to note that this type of protocol differs from NetQASM[18] and InQuIR[19]. NetQASM and InQuIR are quantum network-oriented intermediate representations based on quantum communication protocols, used to control the devices and processes involved in quantum communication.
II-B Compilation, Optimization, Qubit Mapping, and Routing of Quantum Circuits
Quantum logic circuits can represent the quantum computing component of quantum programs. The quantum circuit model can also depict the actual physical operation steps in a quantum processor. The quantum software stack provides high-level programming languages for designing the former. The latter typically corresponds to a sequence of machine instructions that directly operate the quantum device. Quantum circuit compilation is the process of converting quantum logic circuits into sequences of machine instructions. The transformation from an initial quantum logic circuit to a sequence of machine instructions involves multiple structural conversions of the quantum circuit. These conversions include steps such as unitary matrix decomposition, qubit mapping, qubit routing, optimization, and compilation into a sequence of machine instructions.
Quantum computing software such as Qiskit and QASMTrans[20] provides modern support for quantum circuit compilation. Paulihedral[21] is a compiler designed specifically for the VQE algorithm. Similarly, application-oriented compilers also include Twoqan[22], which is dedicated to QAOA circuits. CaQR focuses on the generation of dynamic circuits[23], is a compiler specifically for pulse-based devices. AutoComm[24] and QuComm[25], on the other hand, provide support for distributed devices.
Qubit mapping and routing are essential steps for quantum logic circuits to be executed by quantum processors. Qubit mapping assigns corresponding physical qubit resources to each logical qubit, while qubit routing ensures the connectivity constraints between physical qubits. Optimization aims to obtain compiled circuits with good computational performance. Sabre[26] proposes a method to minimize the number of ancillary qubits. TOQM[27], on the other hand, aims to reduce the depth of the translated circuits[28], among others, investigates optimization schemes related to quantum gate aggregation.
II-C Quantum Program Profiling
Quantum program profiling is of great significance for fully harnessing the potential of quantum computing. With the development of quantum software, quantum programs often contain not only pure quantum gate operations but also other classical instruction execution processes. Therefore, quantum program profiling involves the analysis of quantum circuits and operational subprocesses.
Quantum computing performance analysis benchmarks can be utilized for quantum circuit analysis. Currently, numerous proposed benchmarks are employed for quantum computing performance analysis, primarily focusing on evaluating the performance of quantum processors when executing specific circuits. These metrics include quantum volume[29], Q-score[30], quantum LINPACK[31], and quantum process tomography[32], as well as using quantum applications like VQE[33] and QAOA[34] to analyze the performance of quantum processors. SupermarQ further proposes multiple analysis benchmarks. Clearly, these benchmarks also reflect the performance of the quantum circuits used for testing on the execution device. The benchmarks proposed by SupermarQ[35], such as inter-qubit communication volume, critical depth, coherence rate, parallelism, and qubit activity, can be calculated based on available device information and quantum circuit structure, and thus can also be applied to device-performance-oriented quantum circuit analysis.
Ideas from classical program analysis can be applied to analyze the operational processes in quantum programs. Qprof[36], based on gprof[37], can be used for quantum program analysis to obtain metrics such as subprocess invocation rates, qubit occupancy rates, and time consumption. In this paper, QPanda3 is introduced, which applies process analysis methods to the interconnectivity analysis of quantum logic gates and qubit utilization analysis.
III Overview of QPanda3
QPanda3 is an advanced quantum programming framework that provides comprehensive support for users in quantum computing across both software and hardware aspects. As illustrated in Figure 2, QPanda3 not only offers abstractions for quantum programs and computing devices but also provides multiple related components and tools.
III-A Oriented Towards Quantum Computing
The core of quantum computing software lies in quantum programs, for which QPanda3 provides multi-level abstractions. Specifically, QPanda3 describes quantum programs using a quantum gate-based circuit model, while distinguishing between circuits without measurements and those with measurements to emphasize the difference between measurement operations and quantum gate operations. This distinction is directly reflected in QProg and QCircuit, which are part of the high-level programming language mechanism provided by QPanda3. The high-level programming language mechanism of QPanda3 uses instruction sequences within Python scripts to represent quantum programs. Users can design such quantum programs using Python and the APIs provided by QPanda3. Although high-level programming languages like Python are more user-friendly, they also present issues such as complex mapping relationships with hardware machine instructions and low transmission efficiency. QPanda3 offers a streamlined and efficiently transmittable intermediate representation of quantum programs. Furthermore, QPanda3 can compile these programs into a machine-level instruction stream format for quantum programs.
QPanda3 abstracts quantum computing hardware to meet users’ needs for executing quantum programs and conducting hardware-software co-design. In terms of computing resources, QPanda3 abstracts real quantum processors, quantum simulators, and classical processors, thereby facilitating users to efficiently utilize these resources to complete computational tasks. Regarding the physical attributes of quantum systems, QPanda3 provides noise simulation and Hamiltonian simulation, supporting users in conducting related experimental research.
Centered around quantum programs, QPanda3 offers a rich set of components. A significant portion of these components is designed to meet the needs of hardware-software co-design, including modules specifically for quantum circuit optimization and compilation, modules tailored for device-based quantum program analysis, and application interfaces that facilitate local-cloud collaboration. Other components provide various supports for quantum programs, primarily including a quantum program translation module to ensure program portability, a quantum program visualization module to enhance design efficiency, and a variational quantum circuit module for batch generation of circuits and post-processing.
Figure 2: Framework
QPanda3 aims to fully satisfy users’ requirements for quantum computing-related work and therefore provides numerous useful tools. These tools pertain to quantum information, operators, and testing convenience. The tools related to quantum information offer multiple representations of quantum states and quantum channels, along with several information analysis utilities. The tools concerning operators provide corresponding transformation and simulation tools. The testing tools include interfaces for randomly generating quantum circuits, which can produce a large number of random circuits based on different configurations, thereby meeting users’ diverse testing needs.
III-B From QPanda2
QPanda3 fully leverages the successful experience of QPanda2 in terms of design and functionality. It continues the support for the Python language, enabling researchers and developers to quickly implement quantum algorithms using Python’s concise syntax and rich ecosystem. This design not only lowers the threshold for quantum programming but also enhances development efficiency, making it particularly suitable for beginners in quantum computing and cross-disciplinary researchers. QPanda3 introduces OriginIR as an efficient intermediate representation for quantum programs and provides conversion tools between OriginIR and OpenQASM. This design facilitates seamless migration of quantum programs across different frameworks and hardware platforms, while also providing convenience for the optimization and analysis of quantum circuits. QPanda3 supports multiple types of quantum simulators (such as full amplitude simulators and partial amplitude simulators), quantum noise simulation, Hamiltonian simulation, and management of classical registers. This comprehensive hardware abstraction capability allows researchers to verify and optimize quantum algorithms in various scenarios. To reduce the learning curve for users, QPanda3 retains the quantum circuit construction methods consistent with QPanda2. Users can quickly get started with familiar interfaces and syntax, while enjoying the performance improvements and functional extensions offered by the new framework.
III-C Enhanced Performance
QPanda3 demonstrates significant performance improvements, with notable advantages primarily manifesting in efficient quantum program transmission and high-performance quantum circuit compilation.
(1)Efficient Quantum Program Transmission
QPanda3 has designed a set of efficient protocols specifically for quantum program transmission, aiming to optimize the process of transmitting quantum programs between terminal and cloud devices, as well as among other devices. By standardizing the representation format and transmission process of quantum programs, this protocol significantly enhances the efficiency of users submitting tasks to quantum computing devices. Specifically, the protocol supports encapsulating and transmitting quantum programs in a compact intermediate representation form, avoiding the inefficiencies caused by non-uniform formats or excessive redundant information in traditional transmission methods. Additionally, the protocol employs efficient compression and encoding techniques, effectively reducing the amount of data during transmission, thereby saving communication bandwidth between devices and accelerating interaction speed. In practical applications, this protocol is not only suitable for task submission between user terminals and cloud-based quantum computing devices but also for program sharing and collaborative computing among different quantum computing devices. For example, in distributed quantum computing scenarios, multiple quantum devices can quickly exchange quantum programs and data through this protocol, achieving more efficient parallel computing and resource utilization. Furthermore, the design of this protocol takes into account the compatibility and scalability of quantum programs, adapting to the evolution needs of future quantum hardware and algorithms.
(2) Efficient Quantum Circuit Optimization and Compilation
The implementation and optimization of quantum algorithms still face numerous challenges, particularly in the construction and compilation of quantum circuits. As an advanced quantum computing framework, QPanda3 is dedicated to addressing these issues and providing efficient quantum circuit compilation tools. Through optimized implementation, QPanda3 significantly accelerates the process of constructing quantum logic circuits from Python code, enhancing the development efficiency of quantum algorithms. Secondly, the framework introduces the Transpiler lass, which enables device-oriented quantum circuit compilation. It dynamically adjusts compilation strategies based on the characteristics of the target hardware, thereby improving the execution efficiency and portability of quantum programs. Furthermore, QPanda3 optimizes the quantum circuit compilation process from multiple aspects, significantly enhancing the efficiency of quantum circuit compilation. With efficient quantum circuit compilation technology, QPanda3 provides strong support for the implementation and optimization of quantum algorithms.
III-D Software-Hardware Co-Design Oriented
QPanda3 is a modern quantum computing software development kit oriented towards software-hardware co-design. This is manifested in various key scenarios including quantum program design, quantum program transmission, quantum circuit compilation, quantum program analysis, and quantum program execution. As illustrated in Figure 3, QPanda3, running on both terminal and cloud devices, fully meets the critical requirements of quantum computing based on software-hardware co-design.
Figure 3: Workflow
(1) Cloud Execution and Inter-device Transmission of Quantum Programs
Due to the unique characteristics of quantum computers, the paradigm of terminal design for quantum programming coupled with cloud execution has been widely accepted and has found numerous practical applications. A hallmark of this paradigm is the separation of devices used for program design from those used for program execution, with the two interconnected and collaborating through communication links. Quantum programs are transmitted from the design end to the computation end via these communication links. QPanda3 is not merely a tool for quantum programming; rather, it is a collaborative tool that adheres to the aforementioned paradigm. On one hand, QPanda3 provides convenient access to the quantum cloud, allowing quantum programs designed on QPanda3 to be directly executed on various computing devices within the quantum cloud. On the other hand, to optimize the efficiency of quantum program transmission between devices, QPanda3 has designed a compact intermediate representation for quantum programs. Additionally, QPanda3 offers tools for converting between quantum programs expressed in high-level programming languages and this intermediate representation. Furthermore, serving as an interactive interface for users to access quantum computing devices, QPanda3 enables users to obtain relevant information about the computing devices.
(2) Hardware-Software Co-Design for Optimized Compilation of Quantum Circuits
Quantum circuit compilation is the process of transforming quantum logic circuits into sequences of machine instructions. This process, which is a software-based procedure, necessitates a comprehensive consideration of the actual connection topology of physical qubits and the physical operations that the device employs to control these qubits, with such operations typically corresponding to a limited set of quantum gates. Clearly, quantum circuit compilation represents a typical scenario of hardware-software co-design. The extent of various optimizations applied during the quantum circuit compilation process influences the degree to which quantum programs utilize device performance, and directly impacts the correctness of quantum computation results. QPanda3 decouples key steps in quantum circuit compilation, including qubit mapping, qubit routing, and machine instruction generation, thereby enabling users to compile and generate machine instruction sequences that can be efficiently executed on specified devices, based on different physical qubit layouts and native quantum gate sets. During the circuit compilation process, QPanda3 implements multiple software-level optimization strategies, some of which accelerate compilation speed, while others enhance the execution performance of the compiled circuits.
(3) Hardware-Software Co-Design for Quantum Program Analysis and Design
The goal of quantum programming is to develop quantum programs that can solve target problems, fully utilize computational resources, and exhibit acceptable quantum computing performance. For the quantum computing component, the key to quantum programming lies in obtaining efficient quantum logic circuits.
QPanda3 enables users to design quantum programs using a high-level programming language (Python) and a user-friendly intermediate representation (OriginIR). QPanda3 provides a variety of quantum simulators, which can be deployed on the quantum cloud as services for users. These simulators are also built into QPanda3, allowing users to run them using the computing resources of their personal computers (CPUs, GPUs). Quantum simulators on the cloud and on personal computers can be used to verify whether the theoretical execution results of quantum programs match expectations, thereby guiding the further design of quantum programs. Collaborating with local or cloud computing resources to validate the correctness of quantum programs represents a classic hardware-software co-design scenario. There may be multiple implementations of quantum programs that solve the same problem, and different quantum programs may exhibit varying performance on the same device. It is of great significance to select an appropriate implementation (or subroutine) of a quantum program that is tailored to the target problem and the available computing device. Such task- and device-adapted quantum programs can obtain more accurate results more efficiently, aligning more closely with the goals of quantum programming. This type of quantum program design relies on device-based quantum program analysis. QPanda3 provides this analysis capability, facilitating users in designing more efficient quantum programs. Specifically, QPanda3 offers quantum program analysis tools that are oriented towards quantum computing performance.
IV Representation and Transmission of Quantum Programs
IV-A User-Friendly Intermediate Representation – OriginIR
The significant role of high-level intermediate representations of quantum programs in the quantum software stack cannot be overstated. QPanda3 employs OriginIR as a lower-level representation of quantum programs compared to high-level programming languages.
IV-A1 Compatibility with QPanda2
QPanda3 largely retains all the features of OriginIR from QPanda2, while also introducing appropriate extensions. Specifically, the syntax for declaring classical bits, declaring quantum bits, separating adjacent instruction statements, representing block statement scopes, and declaring quantum logic gates and quantum-related operations remains consistent with QPanda2. QPanda3 also supports all the quantum logic gates available in QPanda2. OriginIR instruction strings generated by QPanda2 can be correctly parsed and processed by QPanda3. Conversely, OriginIR instruction strings generated by QPanda3, which do not include auxiliary information such as comments, can also be correctly parsed and processed by QPanda2. OriginIR enables seamless portability of quantum programs between QPanda2 and QPanda3.
IV-A2 Geared Towards Researchers
(1) Complexity and Conversion
The QASM series represents intermediate representations of quantum programs. To investigate the performance of quantum programs, researchers often test and compare the same quantum program on different platforms. QASMBench[38], based on OpenQASM-2, serves as a relevant benchmark for quantum computing-related tests and has been widely adopted in numerous research studies.
Unlike the universality of the quantum circuit model, intermediate representations (IRs) of quantum programs exhibit significant variations due to differences in quantum computing devices, abstraction levels, and syntactic structures. Even IRs of quantum programs targeting the same quantum computing device and operating at similar abstraction levels can still have considerable discrepancies. Taking the QASM series as an example, OpenQASM-2 and OpenQASM-3 differ in many syntactic aspects, to the extent that even Qiskit itself requires separate conversion tools for OpenQASM-2 and OpenQASM-3. When other software platforms provide auxiliary conversion tools, they often only support a subset of the main content within the QASM series. The more complex the syntactic rules of an intermediate representation of a quantum program are, the greater the difficulty in converting it to other quantum program IRs. This limits researchers’ ability to use quantum program IRs for porting and experimental research of quantum programs across multiple platforms. This issue has given rise to research on more universal quantum program IRs, such as cQASM[39].
OriginIR offers a perspective for reducing the difficulties associated with intermediate conversions of quantum programs, from the standpoint of low complexity. OriginIR provides a concise representation of quantum programs based on a quantum logic gate model, with minimal details unrelated to the circuit. Furthermore, OriginIR adopts a unified format for describing classical registers, as opposed to the multiple formats supported by the QASM series. Additionally, the syntactic rules of OriginIR are straightforward. These characteristics enable researchers to achieve the transplantation of quantum programs from OriginIR through simple processing.
To meet the needs of researchers, QPanda3 provides a tool for converting QASM to OriginIR. It should be noted that QPanda3 only supports a subset of the syntactic rules of OpenQASM-2. However, this tool is already capable of handling the majority of use cases in QASMBench.
(2) Readability
To enhance the readability of OriginIR, compared to QPanda2, QPanda3 allows users to freely add line comments and block comments to OriginIR instruction strings, and also permits appropriate indentation of OriginIR line instructions using spaces. Although these comments and indentations are filtered out during the process of parsing OriginIR instructions in QPanda3, the information in the comments provides important reference for users, and the indentation significantly improves the readability of the OriginIR instruction sequence. The improvement in readability will greatly facilitate researchers in conducting quantum computing-related research based on OriginIR.
IV-B Intermediate Representation for Quantum Program Transmission - OriginBIS
Quantum program transmission is a crucial component of the quantum computing software stack. Although user-friendly intermediate representations (IRs) such as OriginIR can be used to transmit quantum programs in their entirety, these IRs carry a considerable amount of redundant information. Aspects like readability are unnecessary for mere inter-machine quantum program transmission, and excessive redundant information significantly reduces the efficiency of quantum program transmission and increases unnecessary transmission costs. Compressing user-friendly IRs like OriginIR before transmission is an optional solution to reduce quantum program transmission costs. However, this approach adds compression and decompression steps, increasing the time cost of the conversion process. QPanda3 has designed a binary instruction stream (BIS), named OriginBIS, as an intermediate representation for quantum program transmission. This IR operates at the binary machine instruction level, enabling efficient quantum program transmission without adding unnecessary conversion time costs.
OriginBIS employs a stream-based approach for organization and transmission. On one hand, consecutive quantum program instructions are concatenated in data packets and transmitted over communication links in a streaming manner. On the other hand, data within individual quantum program instructions is concatenated in data packets using a streaming approach. This stream-based method allows OriginBIS to provide efficient binary representations tailored to different quantum program instructions and also supports the application of specialized optimization techniques for data transmission in quantum programs, thereby enhancing the efficiency of quantum program transmission.
OriginBIS employs a classification-based binary instruction format alignment scheme to enhance the speed at which devices process OriginBIS instructions. Specifically, OriginBIS first classifies the various instructions in a quantum program based on their function, the type and number of associated data. Then, it designs a binary instruction format with a unified length and bit fields for each category of instructions. This provides efficient support for the instruction dispatch and execution of quantum program instructions within the device. The classification strategy allows OriginBIS to provide adaptive and efficient binary instruction formats for complex instruction types, and the fact that instructions within the same category share the same format enables efficient processing of each category of instructions by the device. In this way, OriginBIS effectively balances the richness of quantum program instructions and the efficiency of device processing of these instructions.
In some quantum program transmission scenarios, the time cost of transmission may be a primary concern. To address this, OriginBIS incorporates an adaptive compression scheme. This scheme is a lightweight data compression approach based on variable-length integer encoding. It applies limited compression to the quantum program.
V Optimized Compilation of Quantum Circuits
The compilation process of quantum circuits involves the transformation from high-level abstract descriptions to hardware-executable instructions. It encompasses the entire process from initialization to final translation, with each step closely interconnected, collectively forming a complete, efficient, and reliable compilation chain. Achieving high-quality execution of quantum circuits and obtaining desirable results relies on effective adaptation and utilization of quantum computing resources. High-quality compilation of quantum circuits is a crucial approach to breaking through the performance bottlenecks in their execution. Efficient compilation of quantum circuits aims to reduce the time and economic costs associated with processing large-scale quantum circuits. The practical application of quantum software is inseparable from efficient quantum circuit compilation schemes.
V-A Transpiler of QPanda3
QPanda3 features a device-oriented modern quantum circuit transpiler. Significant differences exist among various quantum processors. Regardless of the disparities in implementation technologies, there are substantial variations in the topological structures of physical qubits among different quantum processors. Not only do different product series exhibit marked differences, but different versions within the same series may also vary due to factors such as the number of qubits. Additionally, dynamic allocation of physical qubit resources can lead to changes in the topological structure of available physical qubits on the same quantum processor. These factors pose challenges for quantum circuit transpilation. Therefore, modern quantum circuit transpilers need to generate high-quality transpiled circuits that are tailored to the current resource status of the quantum computing device. The quantum circuit transpiler built into QPanda3 takes the topological structure of available physical qubits of the quantum computing device as an input parameter and transpiles a specific quantum logic circuit into a sequence of machine instructions that can be efficiently executed on the corresponding device.
QPanda3 decouples the key steps in the internal implementation of its quantum circuit transpiler, thereby enhancing the flexibility for functional extensions of the compiler module. These key steps include preprocessing, qubit mapping, qubit routing, optimization, and machine instruction generation. The preprocessing step, as the starting point of the compilation process, is responsible for initializing the quantum circuit, including identifying the qubits, quantum gates, and their dependencies within the circuit, laying the foundation for subsequent operations. The objective of qubit mapping is to map logical qubits to the available physical qubits in the quantum processor. This process involves matching the quantum circuit with the corresponding topological structure of the physical qubits. After qubit mapping, the allocated physical qubits may not satisfy the connectivity constraints originally imposed by the quantum circuit. Qubit routing, as the immediate subsequent step to qubit mapping, is specifically designed to address this issue. Optimization is a crucial part of the quantum circuit compilation process, and reasonable optimization strategies can improve the speed and quality of compilation to varying degrees. Unlike other steps that have a strict sequential order among them, optimization may occur throughout various other steps, enhancing compilation efficiency from different aspects. Some targeted and independent optimization strategies can also be abstracted into a separate step. QPanda3 employs such a strategy, using OptimizationPass for abstraction and management, thereby facilitating multiple optimizations before qubit mapping and after qubit routing.
To promote the development of quantum computing-related research and industries, QPanda3 provides external access to its built-in transpiler through dedicated interfaces. These interfaces enable users to perform device-based quantum circuit compilation tasks. Specifically, users are required to generate the edge set of an undirected graph that corresponds to the topological structure of the available physical qubits. Apart from a few specific constraints that need to be satisfied, users can complete this step based on the information of any quantum computing device. Subsequently, users can utilize the interfaces provided by QPanda3 to generate the corresponding machine instruction sequences based on the designed quantum programs and the edge set data. Quantum programs designed using QPanda3 can seamlessly leverage these interfaces. For quantum programs developed on other platforms, users can employ the intermediate representation conversion tool provided by QPanda3 to facilitate quantum program migration. Furthermore, to further enhance user convenience, QPanda3 offers an interface for randomly generating the edge set of an undirected graph. This interface can also generate classic square, fully connected, and linear topological structures based on the number of qubits.
To achieve efficient quantum circuit compilation, QPanda3 has introduced several improvements over its predecessor, QPanda2. In addition to making the built-in transpiler accessible externally for the first time, QPanda3 has also optimized other components used in quantum program design and the quantum circuit compilation process. QPanda3 has undergone a comprehensive upgrade in its internal conversion mechanisms. During the construction of quantum circuits, extensive internal data structure and algorithm conversions are often required to ensure the correctness and executability of the circuits. QPanda3 employs more efficient data structures and algorithms, optimizing these internal conversion processes and significantly enhancing the speed of circuit construction. Whether it is the merging of quantum gates, simplification of circuits, or allocation of resources, QPanda3 can accomplish these tasks at a faster pace, thereby improving overall compilation efficiency.
V-B Starting with an Example
This section demonstrates the use of the built-in transpiler in QPanda3 through a simple example. In the following explanation, we will also introduce the basic steps for designing quantum programs using QPanda3, as well as useful tools for quantum circuit visualization and topology data generation.
(1) A Quantum Circuit Compilation Task
Figure 4, subgraph (a) illustrates the topology of available qubits on a certain quantum computing device using an undirected graph, with the corresponding physical qubits labeled as $q_{0}$, $q_{1}$, and $q_{2}$. Subgraph (b) depicts a quantum logic circuit. The compilation task in this section is to obtain the machine instruction sequence for this circuit that adapts to the given topology. Subgraph (c) presents the visualization result of the circuit in subgraph (b) using QPanda3. It is worth noting that the CZ gate is a symmetric gate.
(a) Qubits’ Topology
(b) Quantum Logic Circuit
(c) Visualization Result of The Circuit
Figure 4: Quantum circuit compilation task
QPanda3 uses ”pyqpanda3” as its Python package name, which differs from the package name ”pyqpanda” used in QPanda2. This allows users to import both the ”pyqpanda3” and ”pyqpanda” packages simultaneously, thereby ensuring compatibility with different versions. It should be noted that QPanda3 shares most of its naming conventions with QPanda2, so care should be taken to avoid issues related to name conflicts during the import process.
The core components of QPanda3 are exported by the package pyqpanda3.core. These core components include, but are not limited to, natively supported quantum logic gates such as RZ, X1, and SWAP. They also encompass the abstract object QProg for quantum programs. draw_qprog and set_print_options are two interfaces related to quantum circuit visualization tools. Specifically, set_print_options is used to control the number of decimal places displayed for gate parameters when visualizing parametric gates, while draw_qprog is used for visualizing quantum circuits.
The components related to the built-in circuit transpiler in QPanda3 are exported by pyqpanda3.transpilation. The Transpiler class is used for managing circuit compilation. The generate_topology function is used for randomly generating topology data.
from pyqpanda3.core import QProg, draw_qprog, RZ, X1, CZ, SWAP, set_print_options
from pyqpanda3.transpilation import Transpiler, generate_topology
(2) Constructing Quantum Logic Circuits
prog_1 = QProg()
prog_1 << RZ(3,1.57) << RZ(4,3.14) << RZ(5,6.28) << CZ(3,4) << CZ(4,5) << CZ(3,5)
set_print_options(2)
[…]
The output is
Figure 5: Result of Constructing Quantum Logic Circuits
(3) Setting the Topology of Physical Qubits
This line of code defines a two-dimensional Python list. This list stores the edge set of an undirected graph, which can represent the topology of the currently available physical qubits on a quantum computing device. Users can also generate similar data using the generate_topology interface.
topo = [[0,1],[1,2]]
(4) Compiling with Different Optimization Levels
The transpile method of the Transpiler object is used to perform the quantum circuit compilation step. Thanks to the generality of the quantum circuit model, a quantum circuit composed of quantum gates and other operations that correspond one-to-one with machine instructions can be mapped to an executable sequence of machine instructions. This brings many conveniences, including the ability to observe the results of quantum circuit compilation using quantum circuit visualization methods.
Due to the dynamic nature of the quantum circuit compilation process, different but equivalent compiled circuits may be generated each time. The output content is extensive and will not be shown here. In the subsequent sections, some key output results of this program will be extracted and analyzed.
transpiler = Transpiler()
prog_level_0 = transpiler.transpile(prog_1, topo, {}, 0)
[…]
V-C Qubit Mapping and Routing
(1) Qubit Mapping
Mapping logical qubits in a quantum circuit to physical qubits, taking into account the topology and connectivity of the quantum hardware, to ensure the executability of the circuit.
An example of a compilation result from the code in Section 5.2 is shown in Figure 6. Based on the parameters of the first three RZ gates added to the quantum circuit, it can be determined that logical qubit $q_{5}$ is mapped to physical qubit $q_{0}$, logical qubit $q_{4}$ is mapped to physical qubit $q_{1}$, and logical qubit $q_{3}$ is mapped to physical qubit $q_{2}$.
Figure 6: Qubit Mapping Result
(2) Qubit Routing
Inserting necessary SWAP gates between physical qubits to address connectivity limitations in quantum hardware and ensure that quantum gates can be executed smoothly on physical qubits in the predetermined order.
In the quantum logic circuit example in Section 5.2, there are direct CZ gate connections between each pair of the three logical qubits, including a direct connection between logical qubit $q_{5}$ and logical qubit $q_{3}$. However, in the aforementioned qubit mapping result, physical qubit $q_{0}$ and physical qubit $q_{1}$ are not directly connected, which violates the corresponding connectivity constraint.
In the optimization compilation at level 0 in QPanda3, the SWAP gate is implemented equivalently using the following sub-circuit.
Figure 7: Qubit Routing Result
Comparing the two figures below, it can be observed that the compilation effect is equivalent to using SWAP gates to satisfy the connectivity constraints. The equivalent circuit corresponding to the aforementioned compilation result is shown in the figure below. Clearly, this circuit is equivalent to the original quantum logic circuit. It should be noted that the example uses gates such as RZ and CZ, which are not decomposed or transformed during the preprocessing stage.
Figure 8: Using SWAP Gates to Satisfy The Connectivity Constraints
It should be noted that there are multiple ways to perform layout and routing for this quantum logic circuit. When compiling this quantum logic circuit, QPanda3 may not always use this specific layout and routing scheme, as other layout and routing solutions can still satisfy the target requirements.
V-D Optimization
QPanda3 adopts a strategy that combines local and global optimization. Local optimization primarily focuses on optimizing small regions within the quantum circuit, such as merging adjacent quantum gates and eliminating redundant operations, to reduce the complexity and depth of the circuit. Global optimization, on the other hand, considers the optimization of the entire quantum circuit by adjusting the order of quantum gates, reallocating resources, and other means to further improve the execution efficiency of the circuit.
V-D1 Sabre
Sabre[26] is an advanced quantum circuit compilation algorithm. This algorithm incorporates a decay effect and employs a bidirectional heuristic search based on swaps. It achieves a good trade-off between the depth of the quantum circuit and the number of gates. QPanda3 has further optimized the implementation of Sabre and extensively applies it in the compilation process. Sabre plays a crucial role in optimization, layout, and routing, and through efficient algorithm design, it can handle complex quantum circuits, improving the accuracy and efficiency of compilation. The Sabre algorithm is closely integrated with the characteristics of quantum hardware, enabling the compiled circuit to better adapt to the hardware’s execution environment, thereby ensuring efficient execution of the algorithm.
V-D2 Independent Optimization Steps
This step involves isolating certain optimization strategies as a standalone procedure, intended for pre-layout and post-routing optimization. The optimization strategies concerned encompass merging adjacent quantum gates, eliminating redundant operations, simplifying complex gate sequences, and so forth. These measures aim to reduce the depth and complexity of the circuit, thereby enhancing execution efficiency. The following two simplistic examples illustrate the optimization effects. It is noteworthy that the examples utilize gates such as RZ and CZ, which are not decomposed or transformed during the preprocessing stage.
(1) Example 1 - Merging Adjacent Gates
This circuit presents a scenario that is amenable to optimization, where consecutive RZ gates can be consolidated. This scenario can be further generalized to the optimization of various types of consecutive single-gate operations. In this instance, when QPanda3 compiles the circuit at optimization level 0, no optimization strategies are employed. Consequently, the result retains two consecutive RZ gate operations, consistent with the quantum logic circuit as depicted in the figure. However, when QPanda3 compiles the circuit at optimization levels 1 and 2, it merges these two consecutive RZ gates into a single RZ gate. The optimized circuit necessitates only the time required to execute one RZ gate, whereas the pre-optimized circuit consumes approximately twice the duration.
Figure 9: Merging Adjacent Gates
(2) Example 2 - Eliminating Redundant Gates
This circuit showcases a scenario that can be optimized, where consecutive SWAP gates exist and can be eliminated. This scenario can be extended to the optimization of various types of consecutive two-qubit gate operations. When QPanda3 performs quantum circuit compilation at optimization level 2, adjacent SWAP gates are removed. The elimination of redundant gates effectively enhances the utilization of quantum bits.
Figure 10: Eliminating Redundant Gates
V-D3 Pre-placement Optimization and Post-routing Optimization
Prior to the placement stage, QPanda3 conducts a series of preprocessing optimizations on the quantum circuit. These optimization operations aim to alleviate the burden on subsequent placement and routing processes, thereby improving compilation efficiency. For instance, by merging adjacent quantum gates, the number of gates that need to be considered during the placement stage can be reduced. Additionally, by eliminating ineffective operations, the occupation of quantum hardware resources can be minimized. After the routing stage is completed, QPanda3 further optimizes the circuit to enhance its execution efficiency even more.
VI Device-Based Quantum Program Analysis
Whether a quantum program can be executed efficiently on a specific computing device and yield desirable results determines its suitability for performance-sensitive tasks and its advantage in terms of computational resource costs. Efficient utilization of device resources in quantum programs relies on device-oriented fine-tuning. On one hand, skillful program design based on programming experience can be employed; on the other hand, purposeful adjustments can be made according to the runtime performance of the quantum program on the computing device. This section focuses on the latter, introducing the tools provided by QPanda3 for device-based quantum program analysis.
VI-A Runtime State Information and Quantum Program Analysis
In classical computing, runtime state information of various hardware components can be directly obtained through dedicated counters, registers, and debug-level interfaces during program execution. Due to the unique nature of quantum mechanics, however, observing a quantum processor without interrupting its operation poses challenges. Typically, one cannot read information from physical qubits as one would from bits in a classical register without causing the program to halt. Nevertheless, in the NISQ (Noisy Intermediate-Scale Quantum) era, a necessary condition for quantum computing is the ability to drive the directed evolution of qubits externally. Information such as the type and duration of externally initiated drive operations can clearly be acquired and fed back to the software layer. Additionally, statistical information on certain runtime-inaccessible parameters can be obtained through multiple trials. This quantum device information can be published as product specification parameters on the one hand, and re-collected as needed on the other. In the context of quantum program analysis, the former is suitable for analyzing standard product specifications, while the latter is applicable to analyzing the current state of the device.
Software is an essential component for program analysis. Statistical information about the hardware can be used as input data for direct analysis by the software. Alternatively, runtime-related information from the hardware can be collected within the software itself to analyze the program’s performance on the current hardware during the present time period.
QPanda3 is a modern software tool designed to serve quantum computing. It provides device-oriented quantum program analysis tools. The open interfaces allow users to perform device-specific quantum program analysis based on the specification parameters and runtime statistical information of quantum computing devices. Integrated within the quantum cloud, QPanda3 operates in high synergy with quantum computers to meet the demands of quantum program analysis for various objectives. Currently, QPanda3 offers quantum circuit analysis tools tailored to device performance and quantum program performance analysis tools based on workflow and device information.
VI-B Quantum Circuit Analysis for Device Performance
VI-B1 Overview
QPanda3 extends the quantum computing performance benchmarks to the quantum circuit analysis context. Specifically, QPanda3 incorporates key benchmarks proposed by SupermarQ[35], including Program Communication, Critical-Depth, Entanglement-Ratio, Parallelism, and Liveness. These benchmarks serve to measure the quantum computing performance of quantum circuits on specific devices, providing guidance for users to evaluate and improve quantum programs.
When utilizing the open interfaces provided by QPanda3, users are required to supply the compiled circuit as the object of analysis. The compiled circuit can be obtained through the circuit compilation tools introduced in Section 5.
QPanda3 packages and returns the quantum program analysis results based on the aforementioned benchmarks in the form of Python lists. This approach facilitates users in collecting corresponding result data when analyzing a large number of circuits. Additionally, QPanda3 provides corresponding visualization tools.
Here, an example is provided to illustrate how to use QPanda3 to obtain the performance metrics of a quantum program, including Program Communication, Critical-Depth, Entanglement-Ratio, Parallelism, and Liveness. QPanda3 offers quantum program analysis functionality through the draw_circuit_features interface in the pyqpanda3.profiling package. The following code constructs a simple quantum circuit, which is managed using a QCircuit object. Finally, the draw_circuit_features function is utilized to perform the analysis.
from pyqpanda3.profiling import draw_circuit_features
from pyqpanda3.core import QCircuit, H, CNOT, X
circuit = QCircuit(4)
circuit << H(0) << CNOT(0, 1) << CNOT(1, 2) << X(3)
draw_circuit_features(circuit, True)
The corresponding visualization results are shown in Figure 11.
Figure 11: Performance Metrics of A Quantum Program
VI-C Performance Analysis of Quantum Programs Based on Flow and Device Information
Inspired by classical program analysis methods, QPanda3 provides users with a tool for analyzing quantum programs based on flow and device information. This tool assists users in analyzing the utilization rates and execution time percentages of various quantum logic gates within their respective processes in quantum programs, providing a basis for users to evaluate and improve their quantum programs. This tool is implemented based on gprof. Next, we will first demonstrate the usage of relevant interfaces through a code example. Then, centering around this example, we will introduce the differences between this feature of QPanda3 and classical program analysis, as well as the other software components of this functionality.
VI-C1 An Example
QPanda3 provides the functionality for performance analysis of quantum programs based on flow and device information through the draw_circuit_profile interface in the pyqpanda3.profiling package. In the following code, two quantum circuits, cir1 and cir2, are first constructed, and then the draw_circuit_profile function is used to analyze the circuit cir2. QPanda3 abstracts circuits consisting solely of quantum gates using the QCircuit class. Adding one QCircuit object to another represents the addition of a new quantum sub-circuit. The device information used includes the quantum gates supported by the quantum processor employed and the average execution time of these gates on the quantum processor. This information is passed as the second parameter to the draw_circuit_profile function in the form of a Python dictionary.
from pyqpanda3.profiling import draw_circuit_profile
from pyqpanda3.core import QCircuit, H, CNOT, X
cir1 = QCircuit(4)
cir1 << H(0) << CNOT(0, 1) << CNOT(1, 2) << X(3) << X(3)
cir2 = QCircuit(4)
cir2 << X(1) << cir1
draw_circuit_profile(cir2, {’CNOT’: 200, ’H’: 60, ’X’: 30}, True)
In this code, the third parameter of the draw_circuit_profile function controls whether visualization is performed. The corresponding visualization result is shown in Figure 12.
Figure 12: Profile of Circuit
VI-C2 Comparison with Classical Program Analysis
Quantum sub-circuits bear similarities to called functions (subroutines) in classical programs. Both represent the totality of program operations within a certain period of time. The former represents all quantum logic gate operations within a period, while the latter represents all classical program operations within a period. Both can also contain other sequences of operations. A quantum sub-circuit can contain other sub-circuits, and a called function (subroutine) can call other functions (subroutines). Therefore, both can be effectively described using directed graphs. Function (subroutine) calls can be accurately described using edges in a directed graph. Similarly, the inclusion relationships between sub-circuits can also be accurately described using edges in a directed graph. From the example above, it can be seen that QPanda3 accurately describes the relationships between sub-circuits and between circuits and quantum gates using a directed graph. In Figure 11, QCircuit_0 corresponds to the first instantiated QCircuit object, cir1, and QCircuit_1 corresponds to the second instantiated QCircuit object, cir2. The arrow from QCircuit_1 to QCircuit_0 indicates that the quantum circuit corresponding to QCircuit_1 (cir2) has the quantum circuit corresponding to QCircuit_0 (cir1) as a direct sub-circuit. The arrow from QCircuit_0 to H indicates that the quantum circuit corresponding to QCircuit_0 (cir1) directly contains the H gate. Obviously, the H gate in the diagram is also a sub-circuit of QCircuit_1 (cir2). For clarity, the term ”direct” is used here to specifically refer to relationships like those between QCircuit_0 and QCircuit_1 (or between QCircuit_0 and H).
During execution, there is a fundamental distinction between quantum sub-circuits and called functions (subroutines) in classical programs. In a classical subroutine, each instruction reutilizes the same hardware computing resources, such as the program instruction pointer register, stack register, and program status word register. The program status word register is employed to facilitate subroutine calls in classical programs. However, to prevent the collapse of the quantum system, quantum computing processors that handle entanglement will map all operations in a quantum circuit onto the quantum processor in a single step. That is, sub-circuits like cir1 are fully expanded when cir2 is executed, rather than being invoked and executed through instruction jumps and instruction stack maintenance as in classical programs.
Flow-based quantum program analysis holds significant importance. In the actual execution of quantum circuits, sub-circuits are unfolded, yet their utilization brings great convenience during the design phase of quantum logic circuits. On the one hand, some sub-circuits are adaptations of classical subroutines. On the other hand, some sub-circuits are sequences of quantum gates designed for specific functions. Through flow-based and device-informed quantum program performance analysis, users can ascertain the resource usage of various quantum sub-circuits and quantum gates during execution, thereby adjusting the implementation schemes of corresponding sub-circuits accordingly.
VI-C3 Output Analysis Results
The analysis results provided by draw_circuit_profile offer the frequency and time consumption percentage of each sub-circuit and quantum gate within the quantum circuit. The percentages in Figure 11 represent the corresponding time consumption percentages, and the ”Nx” notation indicates the frequency value N. When multiple arrows converge on a single node, the respective usage frequencies and time consumption percentages are summarized using a summation approach.
QPanda3 also supports outputting results in the form of gprof-style reports. These outputs provide more specific and detailed analysis data. Figure 13 presents partial results from the aforementioned example code.
Figure 13: Gprof-Style Report
(a) IRs’ Encode Time
(b) IRs’ Decode Time
(c) IRs’ Post-encoding Size
(d) OriginBIS’ Encode Time
(e) OriginBIS’ Decode Time
(f) Visualization of the circuit
Figure 14: For Circuit Count
VII Experiments
VII-A Performance Experiments of OriginBIS
QPanda3 has specifically designed OriginBIS for quantum program transmission. A typical application scenario involves users designing quantum programs on local devices and executing them using abundant computing resources on quantum clouds. In this scenario, quantum programs exist as memory objects in high-level programming languages on the source device. After transmission, they exist as directly executable machine instructions on the target device. This scenario comprises three stages: firstly, the source device converts the quantum program in the form of a high-level programming language memory object into an intermediate representation of the quantum program; secondly, the intermediate representation is transmitted to the target device using a communication link; and finally, the target device converts the intermediate representation into directly executable machine instructions. It should be noted that the scope of the concept of quantum program transmission may vary depending on specific business requirements. If the business only requires the quantum program on the source device to appear on the target device, then transmitting the data storing the intermediate representation of the quantum program from the source device to the target device completes the quantum program transmission, which only includes the second stage. If the business requires quantum program transmission from a memory object on the source device to executable machine instructions on the target device, then the quantum program transmission should encompass all three stages. For convenience of description, in the experiments of this paper, the quantum program transmission process refers to the second stage, the first stage is termed the Encode stage, and the third stage is termed the Decode stage. The experiments in this paper primarily focus on testing and comparing the performance of OriginBIS across these three stages.
This paper tests the performance of OriginBIS in quantum program transmission and quantum program format conversion through multiple comparative experiments. The experimental results demonstrate that OriginBIS significantly improves the processing efficiency of the corresponding processes.
(a) IRs’ Encode Time
(b) IRs’ Decode Time
(c) IRs’ Post-encoding Size
(d) OriginBIS’ Encode Time
(e) OriginBIS’ Decode Time
(f) OriginBIS’ Post-encoding Size
Figure 15: For Circuit Depth
(a) IRs’ Encode Time
(b) IRs’ Decode Time
(c) IRs’ Post-encoding Size
(d) OriginBIS’ Encode Time
(e) OriginBIS’ Decode Time
(f) OriginBIS’ Post-encoding Size
Figure 16: For Qubit Count
VII-A1 Evaluation Metrics
(1) Encode Time and Decode Time
For the Encode and Decode stages, the speed at which the device completes the entire processing operations of the respective stages reflects the processing efficiency of the corresponding processes. Considering that it is difficult to quantify the amount of processing operations performed by the device into specific numerical values, this paper uses the time taken to complete the same conversion task to reflect the impact of the intermediate representation on processing efficiency during the Encode and Decode stages. Specifically, for the same conversion task, the longer the time taken, the lower the processing efficiency; the shorter the time taken, the higher the processing efficiency.
(2) Post-encoding Size
Unlike the Encode and Decode stages, where time is used as a metric, this paper employs the data size of the quantum program after encoding (Post-encoding Size) to measure the impact of the intermediate representation on transmission efficiency during the quantum program transmission stage. The quantum program transmission stage constitutes a classical communication process, where latency is a crucial performance evaluation metric. However, in practical communication environments, latency is influenced by various factors, including the bandwidth and distance of physical communication links and devices, as well as the classical network protocols used for data transmission. Quantum program transmission primarily involves transmitting the intermediate representation of the quantum program as pure data. The data size of the intermediate representation is the only component of the quantum program intermediate representation that is directly related to communication latency, excluding the influences of physical communication links, devices, and classical network protocols. Specifically, for the same quantum program transmission task, a larger Post-encoding Size results in lower transmission efficiency, while a smaller Post-encoding Size leads to higher transmission efficiency.
VII-A2 Experimental Setup
This paper conducts comparative experiments from three aspects: the number of circuits, circuit depth, and the number of qubits. Batch transmission of a large number of circuits is one of the real and essential practical application requirements. Circuit depth is a crucial factor that affects the execution results of quantum circuits. The number of qubits measures the amount of quantum computing resources occupied by quantum circuits. These three metrics serve as important references for users when submitting quantum programs and for devices when processing quantum programs for execution. The experiments in this paper are carried out using the scheme shown in Table I. For the experiment on the number of quantum circuits, the number of quantum circuits is taken as the independent variable, while the circuit depth is controlled at a fixed value of 500 and the number of qubits is controlled at a fixed value of 72. For the experiment on circuit depth, the circuit depth is taken as the independent variable, with the number of circuits controlled at a fixed value of 500 and the number of qubits controlled at a fixed value of 72. For the experiment on the number of qubits, the number of qubits is taken as the independent variable, with the circuit depth controlled at a fixed value of 500 and the number of circuits controlled at a fixed value of 500. It should be noted that the parameters related to circuits are not limited to the number of circuits, circuit depth, and the number of qubits. Moreover, there may be numerical constraints among these parameters. For example, when the number of quantum gates is fixed, there is a certain negative correlation between circuit depth and the number of qubits. To minimize the influence of other parameters and the relationships among them, this paper adopts a randomized technical approach, generating random circuits based on the conventions shown in Table I to test the performance of quantum program intermediate representations in quantum program transmission. Relevant experiments are conducted on OriginBIS, OriginIR, and QASM to compare the quantum program transmission performance of these intermediate representations.
TABLE I: Experimental Setup For Quantum Program Transmission
Experiment Circuit Count Circuit Depth Qubit Count
For Circuit Count Independent Variable 500 72
For Circuit Depth 500 Independent Variable 72
For Qubit Count 500 500 Independent Variable
VII-A3 Experimental Results and Analysis
(1) For Circuit Count
The experimental results for the number of circuits are shown in Figure 14. The grouped columns chart in subplots (a), (b), and (c) demonstrates that OriginBIS is significantly smaller than OriginIR and QASM in terms of Encode Time, Decode Time, and Post-encoding Size. This indicates that in these experiments, OriginBIS outperforms OriginIR and QASM in the efficiency of quantum program transmission. Through simple numerical estimation, it can be found that for Encode Time, OriginIR is approximately 17 times that of OriginBIS, and QASM is approximately 80 times that of OriginBIS; for Decode Time, OriginIR is approximately 20 times that of OriginBIS, and QASM is approximately 40 times that of OriginBIS; for Post-encoding Size, OriginIR is approximately 3.5 times that of OriginBIS, and QASM is approximately 6.5 times that of OriginBIS. The point-line charts in subplots (d), (e), and (f) show that for OriginBIS, Encode Time, Decode Time, and Post-encoding Size increase with the growth in the number of circuits.
(2) For Circuit Depth
The experimental results regarding circuit depth are presented in Figure 15. The grouped columns chart in subplots (a), (b), and (c) demonstrate that OriginBIS exhibits significantly lower values in terms of Encode Time, Decode Time, and Post-encoding Size compared to both OriginIR and QASM. This indicates that in these experiments, OriginBIS outperforms OriginIR and QASM in terms of efficiency for quantum program transmission. Upon simple numerical estimation, it can be observed that for Encode Time, OriginIR is approximately 18 times that of OriginBIS, while QASM is about 88 times; for Decode Time, OriginIR is roughly 21 times that of OriginBIS, and QASM is approximately 39 times; for Post-encoding Size, OriginIR is around 3.5 times that of OriginBIS, and QASM is about 6.5 times. The point-line charts in subplots (d), (e), and (f) reveal that for OriginBIS, Encode Time, Decode Time, and Post-encoding Size increase with the growth of circuit depth.
(a) IRs’ Encode Time
(b) IRs’ Decode Time
(c) IRs’ Post-encoding Size
(d) OriginBIS’ Encode Time
(e) OriginBIS’ Decode Time
(f) OriginBIS’ Post-encoding Size
Figure 17: For Gate Count
(3) For Qubit Count
The experimental results concerning the number of qubits are illustrated in Figure 16. The grouped columns chart in subplots (a), (b), and (c) indicates that OriginBIS exhibits significantly lower Encode Time and Decode Time compared to both OriginIR and QASM. Additionally, for qubit counts of 50 or more, OriginBIS also demonstrates lower Post-encoding Time compared to OriginIR and QASM. This suggests that in the majority of the samples tested, OriginBIS outperforms OriginIR and QASM in terms of efficiency for quantum program transmission. Excluding the 90-qubit data in subplot (b) and the 40-bit data in subplot (c), upon simple numerical estimation, it can be observed that for Encode Time, OriginIR is approximately 21 times that of OriginBIS, while QASM is about 93 times; for Decode Time, OriginIR is roughly 20 times that of OriginBIS, and QASM is approximately 39 times; for Post-encoding Size, OriginIR is around 3.5 times that of OriginBIS, and QASM is about 6.5 times. The point-line charts in subplots (d), (e), and (f) reveal that for OriginBIS, there is a trend of increasing Encode Time and Decode Time with the growth in the number of qubits, albeit with considerable data fluctuation. However, Post-encoding Time consistently increases with the increase in the number of qubits.
(4) Gate Count
During the experiments examining circuit count, circuit depth, and qubit count, it was observed that the number of quantum gates varies with changes in these parameters. Figure 18 illustrates the trend in the number of quantum gates, demonstrating a strong positive linear correlation with the value of the third parameter when the values of any two of these three quantities are held constant.
The number of gates is also an important metric for evaluating quantum programs. By combining the gate count data from Figure 18 with the corresponding Encode Time, Decode Time, and Post-encoding Size data, the results presented in Figure 17 were obtained. Subplots (a), (b), and (c) indicate that the Encode Time, Decode Time, and Post-encoding Size for OriginBIS are significantly lower than those for OriginIR and QASM. Subplots (d), (e), and (f) reveal that there is a trend of increasing Encode Time and Decode Time with the growth in the number of quantum gates, albeit with noticeable data fluctuations. However, Post-encoding Time increases with the increase in the number of quantum gates, exhibiting a strong linear characteristic.
The experimental results concerning the number of qubits are illustrated in Figure 16. The grouped columns chart in subplots (a), (b), and (c) indicates that OriginBIS exhibits significantly lower Encode Time and Decode Time compared to both OriginIR and QASM. Additionally, for qubit counts of 50 or more, OriginBIS also demonstrates lower Post-encoding Time compared to OriginIR and QASM. This suggests that in the majority of the samples tested, OriginBIS outperforms OriginIR and QASM in terms of efficiency for quantum program transmission. Excluding the 90-qubit data in subplot (b) and the 40-qubit data in subplot (c), upon simple numerical estimation, it can be observed that for Encode Time, OriginIR is approximately 21 times that of OriginBIS, while QASM is about 93 times; for Decode Time, OriginIR is roughly 20 times that of OriginBIS, and QASM is approximately 39 times; for Post-encoding Size, OriginIR is around 3.5 times that of OriginBIS, and QASM is about 6.5 times. The point-line charts in subplots (d), (e), and (f) reveal that for OriginBIS, there is a trend of increasing Encode Time and Decode Time with the growth in the number of qubits, albeit with considerable data fluctuation. However, Post-encoding Time consistently increases with the increase in the number of qubits.
Figure 18: Gate Count for Circuit Count, Circuit Depth and QBit Count
Figure 19: Experimental Results About Circuit Compilation
VII-B Compilation Efficiency Comparison Experiment
VII-B1 Experiment Setup
In this paper, a comparative experiment related to compilation efficiency was conducted on multiple mainstream quantum programming frameworks based on Benchpress. Benchpress is an evaluation suite for multi-quantum computing software development kits, containing thousands of test cases. It allows for the uniform testing of multiple quantum software packages across various performance and functional indicators, with the results reflecting the cost of processing quantum circuits on quantum computing devices using different software packages. All relevant experiments for each software package mentioned in the Benchpress literature were replicated in this study, with the addition of tests for QPanda3. All software packages tested are listed in Table II. It should be noted that a timeout limit of 180 seconds was imposed for executing each test case.
The testing of these software packages was conducted in the same software and hardware environments. The environmental parameters used in our experiments are presented in Table III.
TABLE II: SDKs
SDK Version
amazon-braket-sdk(braket) 1.91.0
bqskit 1.2.0
cirq 1.4.1
pytket(tket) 2.0.1
qiskit 1.4.2
qiskit_ibm_transpiler 0.10.3
pystaq 3.5
pyqpanda3 0.2.0
TABLE III: Software and Hardware Environments
Type Info
CPU Intel(R) Xeon(R) Platinum 8336C CPU @ 2.30GHz
Memory 2.0TB
Operating System Ubuntu 20.04.5 LTS
Python 3.11.11
VII-B2 Evaluation Metrics
This paper adopts the evaluation metrics introduced by Benchpress. They are Standard Pytest Output Type, circuit construction time, manipulate time, and circuit compilation time.
VII-B3 Experimental Results And Analysis
TABLE IV: Experimental Results Evaluated Based on The Criterion of Standard Pytest Output Type
SDK Passed Failed XFailed Skipped Total
braket 7 2 0 1057 1066
cirq 10 2 0 1054 1066
qiskit 1044 0 0 22 1066
QPanda3 1016 28 0 22 1066
In our experiments, the results of the test_status_counts are presented in Table IV. It was observed that many SDKs failed to satisfy the requirements of numerous test cases in Benchpress. Specifically, bqskit was only capable of performing tests related to circuit construction. pytket and pystaq exhibited compatibility issues with the environments of other SDKs, making it impossible to conduct fair testing and comparison using a unified test suite and environment. Furthermore, qiskit_ibm_transpiler demonstrated strong dependency on the IBM computing platform, rendering it untestable under our experimental environment. As evident from Table IV, both braket and cirq skipped a significant number of test cases. According to the Benchpress literature, qiskit skipped 22 test cases and passed 1,044 test cases. Tabel IV indicates that our experiments actually tested multiple SDKs using 1,066 test cases. The performance of qiskit in our experiments was consistent with the reports in the Benchpress literature. QPanda successfully passed $95.3\%$ of all test cases, with the number of skipped cases equal to that of qiskit. This table showcases the excellent versatility of QPanda3 within the unified test suite, Benchpress.
The performance of each SDK in circuit construction is illustrated in Figure 20. It can be observed that QPanda3, qiskit, and cirq passed all seven test cases. For the five test cases, namely test_multi_control_circuit, test_bigint_qasm2_import, test_param_circSU2_100_build, test_clifford_build, and test_QV100_build, QPanda3 demonstrated the shortest execution time. For the test case test_param_circSU2_100_bind, QPanda3’s execution time was almost on par with qiskit, ranking as the second fastest. In the test case test_QV100_qasm2_import, although QPanda3 lagged behind qiskit, its execution time was significantly less than that of cirq and bqskit. Despite QASM not being the native intermediate representation for quantum programs in QPanda3, it still exhibited a circuit construction speed close to that of qiskit in these two test cases. This suggests that the QASM conversion tool in QPanda3 may be highly efficient, or that QPanda3’s circuit construction efficiency is so high that it can offset any inefficiencies in QASM conversion. This figure highlights the advantages of QPanda3 in circuit construction.
Figure 20: Experimental Results About Circuit Construction(shorter is better)
Figure 21: Experimental Results About Circuit Manipulation(shorter is better)
The Benchpress literature highlights that Manipulate is a crucial metric for evaluating SDKs. As shown in Figure 21, only QPanda3 and qiskit passed all four test cases. QPanda3 demonstrated the shortest execution time in three tests: test_multi_control_decompose, test_QV100_basis_change, and test_random_clifford_decompose. For the test_DTC100_twirling, QPanda3’s execution time was only slightly longer than that of qiskit, with a minimal difference between the two. This figure indicates that QPanda3 has a clear advantage in the Manipulate metric.
In our experiments, only QPanda3 and qiskit passed the majority of the test cases related to compilation. We have plotted the corresponding experimental results into multiple scatter plots, as shown in Figure 19, which illustrate the differences between QPanda3 and qiskit in terms of 2Q Gate Count, 2Q Gate Depth, and compilation time for various topological structures. In each plot, the blue dashed line represents the point where the values for QPanda3 and qiskit are equal for the corresponding metric. Points above the blue dashed line indicate that the corresponding value for QPanda3 is less than that for qiskit. As observed from Figure 19, for both 2Q Gate Count and 2Q Gate Depth, the majority of the scatter points are densely distributed along or very close to the blue dashed line, suggesting that the performance of QPanda3 and qiskit is essentially the same. However, regarding compilation time, all scatter points in the four subplots of Figure 19 are located above the blue dashed line, indicating that QPanda3 exhibits higher compilation efficiency than qiskit. This characteristic is not limited by topological structure and is effective for a large number of different circuits. Additionally, it can be observed that only a very few test cases result in compilation times exceeding 10 seconds when using QPanda3. These observations collectively demonstrate the high compilation efficiency of QPanda3.
VIII Conclusion
This paper focuses on efficient compilation and execution in quantum computing, introducing OriginIR and OriginBIS as intermediate representations to enhance programmability, transmission efficiency, and execution performance. Experimental results demonstrate that OriginBIS significantly outperforms OpenQASM 2.0 in encoding and decoding speed as well as information capacity. Additionally, QPanda3 surpasses Qiskit in quantum circuit construction, operation execution, and compilation speed, exhibiting exceptional performance improvements, particularly in large-scale quantum circuit compilation tasks. These advancements lay a solid foundation for the engineering applications of future quantum computing.
IX Future Work
The quantum programming framework QPanda3 is poised to enhance its capabilities for future ”quantum-HPC-AI” integrated computing by focusing on several key advancements. Building on its existing strength of compiling quantum circuits 14.97 times faster than Qiskit in Benchpress tests, QPanda3 will develop hybrid task schedulers to enable seamless coordination between quantum computations and classical supercomputing resources. It aims to integrate AI-driven optimization tools, such as AutoML modules, to automatically reduce quantum gate counts and circuit depths while mitigating noise impacts. The framework will expand its hybrid programming interfaces to support classical HPC technologies like MPI and CUDA, facilitating efficient quantum-classical algorithm interoperability. Additionally, QPanda3 plans to implement a unified resource management platform for dynamic allocation of quantum hardware, supercomputing clusters, and AI accelerators. To strengthen ecosystem integration, it will promote standardization through open-source collaboration, enabling deeper compatibility with classical AI libraries (e.g., TensorFlow) while maintaining its low-latency compilation advantages. These upgrades will position QPanda3 as a core enabler for cross-paradigm applications in finance, drug discovery, and materials science.

### v2

@@
 Abstract
-QPanda3 is a high-performance quantum programming framework that enhances quantum computing efficiency through optimized circuit compilation, an advanced instruction stream format (OriginBIS), and hardware-aware execution strategies.
-These engineering optimizations significantly improve both processing speed and system performance, addressing key challenges in the NISQ era.
-A core innovation, OriginBIS, accelerates encoding speeds by up to 86.9× compared to OpenQASM 2.0, while decoding is 35.6× faster, leading to more efficient data handling, reduced memory overhead, and improved communication efficiency.
-This directly enhances the execution of quantum circuits, making large-scale quantum simulations more feasible.
-Comprehensive benchmarking demonstrates QPanda3’s superior performance: quantum circuit construction is 20.7× faster, execution speeds improve by 3.4×, and transpilation efficiency increases by 14.97× over Qiskit.
-Notably, in compiling a 118-qubit W-state circuit on a 2D-grid topology, QPanda3 achieves an unprecedented 869.9× speedup, underscoring its ability to handle complex quantum workloads at scale.
-By combining high-speed quantum processing with a modular and extensible software architecture, QPanda3 provides a practical bridge between today’s NISQ devices and future fault-tolerant quantum computing.
-It facilitates real-world applications in financial modeling, materials science, and combinatorial optimization, while its robust and scalable design supports industrial adoption and cloud-based deployment.
+In emerging quantum–classical integration applications, the classical time cost—especially from compilation and protocol-level communication—often exceeds the execution time of quantum circuits themselves, posing a severe bottleneck to practical deployment.
+To overcome these limitations, QPanda3 has been extensively optimized as a high-performance quantum programming framework tailored for the demands of the NISQ era and quantum–classical hybrid workflows.
+It features optimized circuit compilation, a custom binary instruction stream (OriginBIS), and hardware-aware execution strategies to significantly reduce latency and communication overhead.
+OriginBIS achieves up to 86.9$\times$ faster encoding and 35.6$\times$ faster decoding than OpenQASM 2.0, addressing critical bottlenecks in hybrid quantum systems.
+Benchmarks show 10.7$\times$ compilation speedup and up to 597$\times$ acceleration in compiling large-scale circuits (e.g., a 118-qubit W-state) compared to Qiskit.
+n high-performance simulation, QPanda3 excels in variational quantum algorithms, achieving up to 26$\times$ faster gradient computation than Qiskit, with minimal time-complexity growth across circuit depths.
+These capabilities make QPanda3 well-suited for scalable quantum algorithm development in finance, materials science, and combinatorial optimization, while supporting industrial deployment and cloud-based execution in quantum–classical hybrid computing scenarios.
 Index Terms:
-Quantum Computing, Software-Hardware Collaborative, High-Performance, Quantum Circuit Compilation, Intermediate Representation
+Quantum Computing, Software-Hardware Collaborative, High-Performance, Quantum-classical Integration, Quantum Circuit Compilation, Intermediate Representation
 Figure 1: Comparison of average time utilized for various quantum circuit processing tasks between QPanda3 and Qiskit.
@@
 I Introduction
-Quantum computing, as a revolutionary computational paradigm, has garnered widespread attention from both academia and industry due to its exceptional parallel computing capabilities and potential to solve problems intractable for classical computers.
-In recent years, significant advancements in quantum hardware have propelled quantum computing from theoretical exploration to practical applications.
-However, to fully harness its potential, robust software frameworks are essential to bridge the gap between users and quantum hardware.
-The effectiveness of quantum computing software frameworks depends on several key factors, including user accessibility, computational efficiency, and seamless hardware integration.
-Against this backdrop, QPanda3 [1], a modern quantum programming framework, has undergone extensive optimization to enhance usability, quantum circuit compilation efficiency, and program transmission performance, thereby driving quantum computing toward broader applications.
-Currently, quantum computing is undergoing a transition from theoretical research to engineering practice.
-Despite the challenges faced by existing quantum hardware, such as noise, decoherence, and limited qubit scalability, engineering optimizations are steadily improving its usability, enabling Noisy Intermediate-Scale Quantum (NISQ) devices to demonstrate computational advantages over classical computers for specific problems.
-This engineering evolution not only determines the practical applications of NISQ devices but also influences the feasibility of large-scale fault-tolerant quantum computing in the future.
-Thus, enhancing the usability, stability, and execution efficiency of quantum computing has become a central challenge in its research and application.
-Key approaches to addressing these challenges include compilation optimization, instruction stream design, and software-hardware co-optimization.
-As a high-performance quantum programming framework, QPanda3 adopts an engineering-driven approach to enhance the practical usability of quantum computing. Its core objectives include:
-  • Enhancing the computational capabilities of existing quantum devices—Through efficient compilation techniques, optimized qubit mapping strategies, and quantum gate compression algorithms, QPanda3 enables NISQ devices to execute target algorithms more efficiently, achieving superior computational performance for specific tasks.
-  • Optimizing software-hardware co-design in quantum computing—Unlike classical computing, quantum computing requires software to have deep awareness of hardware characteristics.
-QPanda3 employs adaptive qubit mapping and hardware-aware compilation strategies to ensure optimal execution across different quantum processors, thereby improving overall computational fidelity.
-  • Advancing quantum computational advantages—Although some NISQ devices have demonstrated computational superiority for certain problems, achieving large-scale quantum computing necessitates further engineering breakthroughs.
-QPanda3 adopts a modular architecture, allowing seamless integration with future fault-tolerant quantum computing systems, thereby reducing the transition cost from NISQ to next-generation quantum technologies.
-The intermediate representation (IR) of quantum programs serves as a crucial bridge between high-level quantum programming languages and low-level hardware instructions, influencing program readability, portability, and execution efficiency.
-However, traditional quantum IR solutions often suffer from poor readability and low transmission efficiency, making it difficult for researchers to understand the mapping between high-level algorithms and low-level hardware execution.
-Furthermore, the lack of a unified IR standard among mainstream quantum programming languages hinders cross-platform algorithm benchmarking and experimental comparisons.
-To address these issues, QPanda3 introduces a novel IR framework featuring OriginIR and an efficient binary instruction set OriginBIS to enhance program readability, transmission efficiency, and cross-platform compatibility.
-In classical-quantum hybrid computing and multi-device collaborative computing scenarios, existing IR solutions are typically optimized for specific hardware architectures, limiting their portability across different quantum platforms.
-By optimizing its IR design, QPanda3 enables efficient migration across diverse quantum platforms, thereby improving task allocation flexibility and execution efficiency.
-Efficient quantum circuit compilation is crucial for transforming abstract quantum algorithms into hardware-executable instructions, encompassing quantum gate decomposition, qubit mapping, and circuit optimization.
-Building upon its predecessor, QPanda2, QPanda3 introduces multiple optimizations that significantly enhance compilation efficiency.
-Benchmarking results indicate that QPanda3 surpasses QPanda2 and other mainstream quantum computing frameworks in circuit compilation speed, dramatically improving quantum algorithm execution efficiency.
-In the NISQ era, where hardware constraints are prevalent, optimizing compilation efficiency is critical for maximizing computational utility.
-As a quantum programming framework focused on software-hardware co-optimization, QPanda3 excels in cloud-based quantum computing environments.
-Given the scarcity and high cost of quantum computing resources, centralized cloud deployment is often adopted to enhance accessibility and resource utilization.
-Additionally, cloud-based quantum computing facilitates the establishment of standardized software-hardware environments, allowing users to conduct experiments and development without direct access to quantum hardware.
-Moreover, the cloud computing model fosters ecosystem collaboration in quantum computing, simplifying resource sharing and algorithm optimization.
-To achieve optimal quantum execution, a comprehensive understanding of quantum hardware characteristics—such as qubit connectivity, error rates, and coherence times—is essential.
-QPanda3 integrates device-aware compilation techniques to minimize circuit depth and gate operations, thereby improving computational efficiency.
-Furthermore, QPanda3 provides device-specific quantum program analysis tools to help users assess execution performance, identify potential bottlenecks, and further optimize quantum programs.
-Through software-hardware co-optimization, QPanda3 enables quantum programs to adapt to different hardware architectures, achieving more efficient computation.
-Experimental results demonstrate that QPanda3 significantly outperforms existing quantum programming frameworks across multiple performance metrics. For example:
-  • Compilation efficiency: Experiments on Benchpress[2], in a fully connected topology, QPanda3’s compilation speed is, on average, 9.46× faster than Qiskit[3], with peak acceleration reaching 77.76×.
-In a square topology, its average compilation speed surpasses Qiskit[3] by 24.6×, with a peak of 869.9×.
-For heavy-hexagon topologies, QPanda3 achieves an average 15.1× speedup, with a peak of 332.0×.
-In linear topology, the average speedup is 10.71×, with peak acceleration reaching 868.7×.
-  • IR transmission efficiency: The encoding speed of OriginBIS is 18.7× faster than OriginIR and 86.9× faster than OpenQASM 2.0 [4].
-The decoding speed is 18.9× faster than OriginIR and 35.6× faster than OpenQASM 2.0.
-The content of this paper is organized as follows: Firstly, we review the related work on quantum computing compilation frameworks.
-Secondly, we provide an overview of the overall architecture and core features of QPanda3.
-Then, we successively introduce the representation method and transmission mechanism of quantum programs, device-based quantum circuit optimization and compilation strategies, and device-based quantum program analysis, delving into the three key components of QPanda3.
-Finally, we validate the performance advantages of QPanda3 through experiments.
+Quantum computing, as a transformative computational paradigm, has attracted significant attention from both academia and industry due to its exceptional parallelism and its potential to tackle problems intractable for classical computers.
+In recent years, substantial advances in quantum hardware have propelled the field from theoretical exploration into practical application.
+However, fully realizing the potential of quantum computing requires robust software frameworks that bridge the gap between end-users and quantum hardware.
+The effectiveness of such frameworks hinges on several critical factors, including user accessibility, computational efficiency, and seamless hardware integration.
+In this context, QPanda3 [1], a modern quantum programming framework, has undergone extensive engineering optimizations to improve usability, circuit compilation efficiency, and program transmission performance—accelerating the practical deployment of quantum computing across diverse application domains.
+Quantum computing is currently undergoing a transition from theoretical exploration to engineering realization.
+Despite persistent hardware challenges—such as noise, decoherence, and limited qubit scalability—ongoing engineering advancements are steadily improving the usability of Noisy Intermediate-Scale Quantum (NISQ) devices.
+These improvements have enabled NISQ systems to demonstrate computational advantages over classical systems in specific problem domains.
+This evolution not only broadens the scope of near-term quantum applications but also lays the groundwork for scalable and fault-tolerant quantum computing in the future.
+However, in contrast to the steady progress in hardware, most existing quantum programming frameworks exhibit a lack of engineering maturity, which poses significant barriers to the practical deployment of quantum–classical hybrid applications.
+In particular, the computational and communication overhead on the classical side often introduces substantial runtime latency, undermining the scheduling efficiency of quantum programs—especially in scenarios involving frequent quantum–classical interactions.
+These challenges are especially pronounced during quantum program compilation and in the implementation of quantum–classical communication protocols.
+Many mainstream frameworks remain at a research-oriented stage, with limited attention paid to compilation efficiency, scalability for large-scale circuits, and deep integration with hardware architectures.
+For example, widely used plaintext-style quantum intermediate representations (QIRs), such as OpenQASM [2] and OriginIR [3], while offering human readability, suffer from overly redundant instruction formats and lack structured expressiveness.
+These limitations complicate backend optimization and hinder efficient quantum–classical co-execution, exacerbating communication latency and creating scheduling bottlenecks.
+Moreover, plaintext QIRs introduce additional obstacles in real-world deployments, including increased parsing overhead, inefficient transmission, and security vulnerabilities.
+These issues collectively restrict the development of efficient, reliable, and scalable quantum compilation and execution infrastructures.
+Addressing these limitations requires a fundamental rethinking of quantum programming frameworks from an engineering perspective.
+This includes optimizing the compilation pipeline, designing efficient intermediate representations, and enabling high-performance execution tailored to the demands of practical quantum–classical systems.
+QPanda3, as a high-performance quantum programming framework, adopts an engineering-centric design philosophy aimed at enhancing the practical usability of quantum computing systems.
+Its core objectives include:
+  • Superior Compilation Performance:
+    QPanda3 employs a customized data structure tailored to the quantum compilation workflow, effectively reducing intermediate-stage information redundancy in instruction reordering, topology-constrained mapping, and gate fusion optimization.
+This design enables efficient and scalable compilation for quantum circuits at the scale of thousands of qubits.
+Benchmark results from Benchpress [4] demonstrate that, in fully connected topologies, QPanda3 achieves an average compilation speed 7.03$\times$ faster than Qiskit [5], with peak speedups up to 123.33$\times$.
+For square, heavy-hexagon, and linear topologies, average and peak speedups reach 15.87$\times$/597.41$\times$, 12.45$\times$/313.86$\times$, and 7.43$\times$/374.48$\times$, respectively.
+  • Quantum-Classical Protocol Optimization:
+    By introducing a concise and efficient binary stream format to represent quantum programs, QPanda3 significantly improves encoding and decoding speeds as well as compression ratios compared to traditional plaintext intermediate representations.
+Specifically, the encoding speed of OriginBIS surpasses that of OriginIR and OpenQASM 2.0 by factors of 18.7$\times$ and 86.9$\times$, respectively; decoding speed is enhanced by 18.9$\times$ and 35.6$\times$.
+This advancement effectively reduces communication latency in quantum hyper-converged applications.
+  • Fast Simulation of Hybrid Quantum Programs:
+    QPanda3 provides modular support for efficient simulation of variational quantum circuits and quantum-classical hybrid algorithms, addressing the rapid iteration and feedback-driven development characteristic of the NISQ era.
+Comparative experiments with Qiskit, PennyLane [6], and DeepQuantum [7] confirm that QPanda3 achieves the shortest gradient computation times and overall superior performance.
+  • Quantum Software and Hardware Profiling:
+    The framework incorporates a comprehensive performance analysis toolchain supporting benchmarking, calibration, and cross-platform performance comparison of both quantum software and underlying hardware, facilitating the validation and advancement of emerging quantum technologies.
+The remainder of this paper is organized as follows: Section 2 reviews related work on quantum compilation frameworks.
+Section 3 outlines the architecture and key features of QPanda3.
+Section 4 presents experimental validation of QPanda3’s performance advantages.
+In addition, the Appendix provides a modular breakdown of QPanda3’s internal components along with illustrative code examples, enabling readers to gain a deeper understanding of its usage and extensibility in practical quantum programming scenarios.
 II Related Work
@@
 However, such representations tend to be relatively complex.
-For example, with the increasing complexity of quantum computing systems and application demands, IBM has continuously extended QASM to develop OpenQASM 2.0 and OpenQASM 3.0.
-As widely adopted versions, OpenQASM 2.0[4] and OpenQASM 3.0[6] not only extend quantum gate operations and qubit management but also introduce more complex syntactic structures and functional features, such as classical control flow and modular programming support.
+For example, with the increasing complexity of quantum computing systems and application demands, IBM has continuously extended QASM to develop OpenQASM 2.0[2] and OpenQASM 3.0 [9].
+As widely adopted versions, OpenQASM 2.0 and OpenQASM 3.0 not only extend quantum gate operations and qubit management but also introduce more complex syntactic structures and functional features, such as classical control flow and modular programming support.
 They are more akin to the style of the classical programming language C compared to the original QASM.
@@
 Similarly, application-oriented compilers also include Twoqan [25], which is dedicated to QAOA circuits.
-CaQR focuses on the generation of dynamic circuits[23], is a compiler specifically for pulse-based devices.
+CaQR[26] focuses on the generation of dynamic circuits, is a compiler specifically for pulse-based devices.
 AutoComm[27] and QuComm [28], on the other hand, provide support for distributed devices.
@@
 In this paper, QPanda3 is introduced, which applies process analysis methods to the interconnectivity analysis of quantum logic gates and qubit utilization analysis.
+II-D Variational Quantum Circuit
+Variational quantum algorithms [41] and quantum machine learning [42], as two important research fields, represent two key directions and milestones in the practical application of quantum computing.
+The Variational Quantum Eigensolver (VQE) has been widely applied in quantum chemistry, condensed matter physics, and optimization problem-solving.
+The Quantum Approximate Optimization Algorithm (QAOA) [37] can be employed for combinatorial optimization problems such as the Max-Cut problem [43], the Traveling Salesman Problem, portfolio optimization, and scheduling.
+The Variational Quantum Classifier [44] is a significant application of variational quantum algorithms in the field of machine learning.
+Meanwhile, other quantum machine learning techniques, such as quantum neural networks, may also utilize non-variational quantum algorithms.
+In quantum computing, most of these critical algorithms rely on parameterized quantum circuits.
+Therefore, parameterized quantum circuits serve as essential components for the practical implementation of quantum computing.
+Variational quantum algorithms adopt the term Ansatz [45].
+In algorithms such as VQE [36], QAOA, and VQC, the term Ansatz can unambiguously refer to a parameterized quantum circuit while also implying the optimization of parameters to solve a target problem.
+However, Ansatz models that are not based on quantum circuits also exist.
+QPanda2 [46] explicitly uses variational quantum circuit to refer to parameterized quantum circuits employed in variational algorithms [47].
+It is worth noting that parameterized quantum circuits designed for variational algorithms can also be repurposed as general parameterized circuits and applied in quantum machine learning algorithms that do not rely on variational methods.
+QPanda3 retains the term *variational quantum circuit* from QPanda2, with parameterized quantum circuits at its core.
+It provides functionalities such as quantum circuit generation, Hamiltonian expectation value calculation, and gradient computation, thereby supporting the use of parameterized quantum circuits in both variational quantum algorithms and quantum machine learning algorithms.
+This differs from frameworks like Qiskit [5], PennyLane [6], and MindQuantum [48], which do not explicitly distinguish between standard quantum circuits (without tunable parameters) and parameterized quantum circuits.
+In contexts related to QPanda3, where no ambiguity arises, the terms Ansatz, parameterized quantum circuit, and variational quantum circuit are used interchangeably without explicit distinction.
+QPanda3 is a foundational quantum programming framework designed for cross-disciplinary research.
+While quantum machine learning-focused frameworks such as PennyLane, MindQuantum, and pyVQNet [49] have been optimized for various aspects of machine learning, they often struggle to support quantum computing tasks in other domains.
+A similar limitation exists with ChemiQ [50], which provides comprehensive support for quantum chemistry but is specialized for that field.
+QPanda3 maintains good compatibility with pyVQNet and ChemiQ.
+However, unlike these frameworks-which incorporate extensive application-specific features tailored to their respective domains-QPanda3’s variational quantum circuits do not extend many application-dependent functionalities.
+Instead, they focus on providing efficient implementations of essential core operations.
+By offering more fundamental building blocks, QPanda3 enables researchers to use a single software platform to customize algorithmic implementations across different application areas, thereby facilitating cross-disciplinary research.
+These essential core functionalities include: Constructing parameterized quantum circuits, Obtaining parameterized quantum states for given parameter values, Reusing variational quantum circuit structures for modular programming, Computing Hamiltonian expectation values, Calculating gradients, and Supporting expressions as parameters in quantum gates.
+This design philosophy allows QPanda3 to serve as a versatile and adaptable tool for researchers exploring diverse quantum computing applications.
+QPanda3 is a quantum programming framework designed for high-performance computing.
+While parameterized quantum circuits can be implemented by combining function definitions in programming languages with the general quantum circuit construction mechanisms provided by programming frameworks, this approach imposes limitations on certain high-performance extensions.
+For instance, when computing Hamiltonian expectation values and gradients, more efficient implementations exist on classical computers-some of which rely on direct manipulation of quantum states stored in classical memory [51].
+Currently, quantum programming SDKs such as Qiskit, PennyLane, pyQuil, MindQuantum, and QPanda all offer Python-based interfaces.
+However, constructing parameterized circuits using Python functions and generic circuit-building mechanisms prevents fine-grained control over memory operations, introduces redundant instructions, and ultimately restricts performance improvements in variational quantum circuit functionalities.
+To address this, QPanda3 provides dedicated data structures and processing logic specifically optimized for variational quantum circuits, aiming to enhance the performance of variational quantum computing tasks.
 III Overview of QPanda3
-QPanda3 is an advanced quantum programming framework that provides comprehensive support for users in quantum computing across both software and hardware aspects.
-As illustrated in Figure 2, QPanda3 not only offers abstractions for quantum programs and computing devices but also provides multiple related components and tools.
-III-A Oriented Towards Quantum Computing
-The core of quantum computing software lies in quantum programs, for which QPanda3 provides multi-level abstractions.
-Specifically, QPanda3 describes quantum programs using a quantum gate-based circuit model, while distinguishing between circuits without measurements and those with measurements to emphasize the difference between measurement operations and quantum gate operations.
-This distinction is directly reflected in QProg and QCircuit, which are part of the high-level programming language mechanism provided by QPanda3.
-The high-level programming language mechanism of QPanda3 uses instruction sequences within Python scripts to represent quantum programs.
-Users can design such quantum programs using Python and the APIs provided by QPanda3.
-Although high-level programming languages like Python are more user-friendly, they also present issues such as complex mapping relationships with hardware machine instructions and low transmission efficiency.
-QPanda3 offers a streamlined and efficiently transmittable intermediate representation of quantum programs.
-Furthermore, QPanda3 can compile these programs into a machine-level instruction stream format for quantum programs.
-QPanda3 abstracts quantum computing hardware to meet users’ needs for executing quantum programs and conducting hardware-software co-design.
-In terms of computing resources, QPanda3 abstracts real quantum processors, quantum simulators, and classical processors, thereby facilitating users to efficiently utilize these resources to complete computational tasks.
-Regarding the physical attributes of quantum systems, QPanda3 provides noise simulation and Hamiltonian simulation, supporting users in conducting related experimental research.
-Centered around quantum programs, QPanda3 offers a rich set of components.
-A significant portion of these components is designed to meet the needs of hardware-software co-design, including modules specifically for quantum circuit optimization and compilation, modules tailored for device-based quantum program analysis, and application interfaces that facilitate local-cloud collaboration.
-Other components provide various supports for quantum programs, primarily including a quantum program translation module to ensure program portability, a quantum program visualization module to enhance design efficiency, and a variational quantum circuit module for batch generation of circuits and post-processing.
-Figure 2: Framework
-QPanda3 aims to fully satisfy users’ requirements for quantum computing-related work and therefore provides numerous useful tools.
-These tools pertain to quantum information, operators, and testing convenience.
-The tools related to quantum information offer multiple representations of quantum states and quantum channels, along with several information analysis utilities.
-The tools concerning operators provide corresponding transformation and simulation tools.
-The testing tools include interfaces for randomly generating quantum circuits, which can produce a large number of random circuits based on different configurations, thereby meeting users’ diverse testing needs.
-III-B From QPanda2
-QPanda3 fully leverages the successful experience of QPanda2 in terms of design and functionality.
-It continues the support for the Python language, enabling researchers and developers to quickly implement quantum algorithms using Python’s concise syntax and rich ecosystem.
-This design not only lowers the threshold for quantum programming but also enhances development efficiency, making it particularly suitable for beginners in quantum computing and cross-disciplinary researchers.
-QPanda3 introduces OriginIR as an efficient intermediate representation for quantum programs and provides conversion tools between OriginIR and OpenQASM.
-This design facilitates seamless migration of quantum programs across different frameworks and hardware platforms, while also providing convenience for the optimization and analysis of quantum circuits.
-QPanda3 supports multiple types of quantum simulators (such as full amplitude simulators and partial amplitude simulators), quantum noise simulation, Hamiltonian simulation, and management of classical registers.
-This comprehensive hardware abstraction capability allows researchers to verify and optimize quantum algorithms in various scenarios.
-To reduce the learning curve for users, QPanda3 retains the quantum circuit construction methods consistent with QPanda2.
-Users can quickly get started with familiar interfaces and syntax, while enjoying the performance improvements and functional extensions offered by the new framework.
-III-C Enhanced Performance
-QPanda3 demonstrates significant performance improvements, with notable advantages primarily manifesting in efficient quantum program transmission and high-performance quantum circuit compilation.
+QPanda3 is an advanced quantum programming framework that provides comprehensive support for quantum computing across both software and hardware domains.
+As illustrated in Figure 2, QPanda3 offers abstractions for quantum programs and computing devices, along with a wide array of supporting components and tools.
+III-A Quantum Computing Orientation
+At the core of quantum software lies the quantum program, for which QPanda3 provides multi-level abstractions.
+Specifically, QPanda3 adopts a quantum gate-based circuit model, distinguishing between circuits with and without measurements.
+This distinction emphasizes the differing roles of measurement operations and quantum gate operations and is directly embodied in the QProg and QCircuit structures provided by QPanda3’s high-level programming interface.
+Quantum programs are expressed as instruction sequences in Python, allowing users to design quantum circuits using QPanda3’s APIs.
+While Python enhances usability, it introduces complexities in hardware instruction mapping and transmission inefficiency.
+QPanda3 addresses these issues by introducing a streamlined intermediate representation (IR) that supports efficient encoding and transmission, and by compiling programs into hardware-level instruction formats.
+For hardware abstraction, QPanda3 encompasses real quantum processors, quantum simulators, and classical computing units, enabling users to leverage diverse resources.
+Additionally, it supports noise modeling and Hamiltonian simulation to facilitate physical-system-level experimentation.
+Surrounding the quantum program, QPanda3 incorporates a rich set of components designed to enable hardware–software co-design.
+These include modules for circuit optimization and compilation, device-specific quantum program analysis, and cloud-based execution interfaces.
+Other auxiliary tools include quantum program translation for portability, visualization tools for circuit design, and a variational circuit module for efficient circuit generation and post-processing.
+Variational quantum circuits—key elements of quantum-classical hybrid computing—are also natively supported in QPanda3.
+Figure 2: Architecture of the QPanda3 Framework
+To further support quantum development, QPanda3 provides tools for quantum state and channel representation, operator simulation, and testing.
+These include utilities for state transformation, information-theoretic analysis, and random circuit generation, catering to diverse testing and benchmarking needs.
+III-B Advancements from QPanda2
+QPanda3 builds on the successful design principles of QPanda2.
+It continues to support Python, offering a concise syntax and rich ecosystem, thereby lowering the entry barrier and accelerating development—particularly for interdisciplinary researchers.
+A key innovation in QPanda3 is the introduction of OriginIR, an efficient intermediate representation, along with conversion tools for interoperability with OpenQASM.
+This enables seamless migration across platforms and supports program optimization and analysis.
+The framework also supports multiple simulator types, noise and Hamiltonian simulations, and classical register management, thereby offering a robust hardware abstraction layer.
+To ease the transition from QPanda2, QPanda3 maintains compatibility with its circuit construction paradigm, allowing users to reuse familiar interfaces while benefiting from enhanced performance and functionality.
+III-C End-to-End Software-Hardware Co-Design in QPanda3
+To address the various challenges encountered during the execution of quantum algorithms on real hardware, QPanda3 adopts a software-hardware co-design paradigm that spans the entire workflow including program design, transmission, compilation, analysis, and execution.
+Figure 3 illustrates the framework’s full-stack integration process in both local and cloud environments.
+Figure 3: Workflow of QPanda3 emphasizing software-hardware co-design.
+The framework realizes an integrated closed-loop encompassing program construction, cloud-based transmission, device-oriented compilation, and adaptive performance analysis.
+III-C1 Cloud Execution and Cross-Device Program Transmission
+With the increasing adoption of the “local design, cloud execution” paradigm in quantum computing, QPanda3 offers the following capabilities to support this scenario:
+  • Employs a compact intermediate representation (IR) that supports efficient, compressed, and encrypted transmission;
+  • Provides APIs and CLI tools to enable seamless interoperability with cloud backend systems;
+  • Supports device introspection interfaces capable of retrieving target quantum device topology, calibration data, and error rates.
+These capabilities empower users to efficiently deploy quantum programs across heterogeneous platforms, thereby enabling truly cloud-native quantum computing operations.
+III-C2 Hardware-Oriented Collaborative Optimization Compilation Mechanism
+QPanda3 utilizes a dynamic feedback-driven compilation architecture that constructs quantitative evaluation models based on real-time hardware performance metrics (e.g., gate fidelity, decoherence times, crosstalk noise) to drive multi-level compilation optimizations.
+The core innovation lies in establishing a closed-loop optimization system of compilation–execution–feedback:
+Hardware-Aware Real-Time Tuning Mechanism
+  • Dynamically monitors chip topology, gate error rates, and other parameters at runtime to generate adaptive mapping strategies;
+  • Supports dynamic instruction set redirection for heterogeneous quantum chips (e.g., superconducting, trapped ion).
+Performance-Driven Layered Compilation
+  • Physical Layer: Gate decomposition strategies dynamically adjusted based on chip T1/T2 parameters, optimizing parallel gate scheduling;
+  • Logical Layer: Incorporates backend feedback on chip information to optimize layout and routing;
+  • System Layer: Orchestrates pipeline scheduling of hybrid classical-quantum instructions to enhance quantum resource utilization.
+Heterogeneous Computing Adaptation Framework
+  • Constructs a unified intermediate representation layer (QIR) supporting cross-platform instruction translation;
+  • Continuously updates a chip characteristics database (QCDB) including gate sets, topology, and noise models;
+  • Employs adaptive backend code generators optimized respectively for control timing on superconducting and trapped ion chips.
+III-C3 Quantum Program Analysis and Optimization
+QPanda3 provides a comprehensive quantum program analysis toolchain supporting multi-dimensional program characteristic evaluation and optimization from the logical layer down to the physical layer.
+The main analysis functions include:
+Program Structure Analysis
+  • Quantum gate-level dependency analysis;
+  • Identification of parallelism and critical paths;
+  • Control flow analysis of quantum-classical hybrid programs.
+Performance Evaluation
+  • Supports fidelity evaluation methods such as cross-entropy benchmarking (XEB) and randomized benchmarking;
+  • Complexity analysis including quantum circuit depth and gate count.
+Resource Bottleneck Analysis
+  • Statistics on gate-level parallelism and qubit utilization;
+  • Detection of routing conflicts under qubit topology constraints;
+  • Pulse-level scheduling and timing analysis.
+III-D Enhanced Performance
+QPanda3 delivers notable performance enhancements, especially in quantum program transmission and high-efficiency circuit compilation.
 (1) Efficient Quantum Program Transmission
-QPanda3 has designed a set of efficient protocols specifically for quantum program transmission, aiming to optimize the process of transmitting quantum programs between terminal and cloud devices, as well as among other devices.
-By standardizing the representation format and transmission process of quantum programs, this protocol significantly enhances the efficiency of users submitting tasks to quantum computing devices.
-Specifically, the protocol supports encapsulating and transmitting quantum programs in a compact intermediate representation form, avoiding the inefficiencies caused by non-uniform formats or excessive redundant information in traditional transmission methods.
-Additionally, the protocol employs efficient compression and encoding techniques, effectively reducing the amount of data during transmission, thereby saving communication bandwidth between devices and accelerating interaction speed.
-In practical applications, this protocol is not only suitable for task submission between user terminals and cloud-based quantum computing devices but also for program sharing and collaborative computing among different quantum computing devices.
-For example, in distributed quantum computing scenarios, multiple quantum devices can quickly exchange quantum programs and data through this protocol, achieving more efficient parallel computing and resource utilization.
-Furthermore, the design of this protocol takes into account the compatibility and scalability of quantum programs, adapting to the evolution needs of future quantum hardware and algorithms.
-(2) Efficient Quantum Circuit Optimization and Compilation
-The implementation and optimization of quantum algorithms still face numerous challenges, particularly in the construction and compilation of quantum circuits.
-As an advanced quantum computing framework, QPanda3 is dedicated to addressing these issues and providing efficient quantum circuit compilation tools.
-Through optimized implementation, QPanda3 significantly accelerates the process of constructing quantum logic circuits from Python code, enhancing the development efficiency of quantum algorithms.
-Secondly, the framework introduces the Transpiler lass, which enables device-oriented quantum circuit compilation.
-It dynamically adjusts compilation strategies based on the characteristics of the target hardware, thereby improving the execution efficiency and portability of quantum programs.
-Furthermore, QPanda3 optimizes the quantum circuit compilation process from multiple aspects, significantly enhancing the efficiency of quantum circuit compilation.
-With efficient quantum circuit compilation technology, QPanda3 provides strong support for the implementation and optimization of quantum algorithms.
-III-D Software-Hardware Co-Design Oriented
-QPanda3 is a modern quantum computing software development kit oriented towards software-hardware co-design.
-This is manifested in various key scenarios including quantum program design, quantum program transmission, quantum circuit compilation, quantum program analysis, and quantum program execution.
-As illustrated in Figure 3, QPanda3, running on both terminal and cloud devices, fully meets the critical requirements of quantum computing based on software-hardware co-design.
-Figure 3: Workflow
-(1) Cloud Execution and Inter-device Transmission of Quantum Programs
-Due to the unique characteristics of quantum computers, the paradigm of terminal design for quantum programming coupled with cloud execution has been widely accepted and has found numerous practical applications.
-A hallmark of this paradigm is the separation of devices used for program design from those used for program execution, with the two interconnected and collaborating through communication links.
-Quantum programs are transmitted from the design end to the computation end via these communication links.
-QPanda3 is not merely a tool for quantum programming; rather, it is a collaborative tool that adheres to the aforementioned paradigm.
-On one hand, QPanda3 provides convenient access to the quantum cloud, allowing quantum programs designed on QPanda3 to be directly executed on various computing devices within the quantum cloud.
-On the other hand, to optimize the efficiency of quantum program transmission between devices, QPanda3 has designed a compact intermediate representation for quantum programs.
-Additionally, QPanda3 offers tools for converting between quantum programs expressed in high-level programming languages and this intermediate representation.
-Furthermore, serving as an interactive interface for users to access quantum computing devices, QPanda3 enables users to obtain relevant information about the computing devices.
-(2) Hardware-Software Co-Design for Optimized Compilation of Quantum Circuits
-Quantum circuit compilation is the process of transforming quantum logic circuits into sequences of machine instructions.
-This process, which is a software-based procedure, necessitates a comprehensive consideration of the actual connection topology of physical qubits and the physical operations that the device employs to control these qubits, with such operations typically corresponding to a limited set of quantum gates.
-Clearly, quantum circuit compilation represents a typical scenario of hardware-software co-design.
-The extent of various optimizations applied during the quantum circuit compilation process influences the degree to which quantum programs utilize device performance, and directly impacts the correctness of quantum computation results.
-QPanda3 decouples key steps in quantum circuit compilation, including qubit mapping, qubit routing, and machine instruction generation, thereby enabling users to compile and generate machine instruction sequences that can be efficiently executed on specified devices, based on different physical qubit layouts and native quantum gate sets.
-During the circuit compilation process, QPanda3 implements multiple software-level optimization strategies, some of which accelerate compilation speed, while others enhance the execution performance of the compiled circuits.
-(3) Hardware-Software Co-Design for Quantum Program Analysis and Design
-The goal of quantum programming is to develop quantum programs that can solve target problems, fully utilize computational resources, and exhibit acceptable quantum computing performance.
-For the quantum computing component, the key to quantum programming lies in obtaining efficient quantum logic circuits.
-QPanda3 enables users to design quantum programs using a high-level programming language (Python) and a user-friendly intermediate representation (OriginIR).
-QPanda3 provides a variety of quantum simulators, which can be deployed on the quantum cloud as services for users.
-These simulators are also built into QPanda3, allowing users to run them using the computing resources of their personal computers (CPUs, GPUs).
-Quantum simulators on the cloud and on personal computers can be used to verify whether the theoretical execution results of quantum programs match expectations, thereby guiding the further design of quantum programs.
-Collaborating with local or cloud computing resources to validate the correctness of quantum programs represents a classic hardware-software co-design scenario.
-There may be multiple implementations of quantum programs that solve the same problem, and different quantum programs may exhibit varying performance on the same device.
-It is of great significance to select an appropriate implementation (or subroutine) of a quantum program that is tailored to the target problem and the available computing device.
-Such task- and device-adapted quantum programs can obtain more accurate results more efficiently, aligning more closely with the goals of quantum programming.
-This type of quantum program design relies on device-based quantum program analysis.
-QPanda3 provides this analysis capability, facilitating users in designing more efficient quantum programs.
-Specifically, QPanda3 offers quantum program analysis tools that are oriented towards quantum computing performance.
-IV Representation and Transmission of Quantum Programs
-IV-A User-Friendly Intermediate Representation – OriginIR
+To address the data representation and communication bottlenecks inherent in large-scale and distributed quantum–classical hybrid computing workflows, QPanda3 introduces a unified, high-efficiency quantum program transmission protocol.
+This protocol enables robust interaction between user clients, cloud-based orchestration layers, and quantum backends.
+At its core lies the Unified Quantum Program Representation Module, which abstracts quantum instructions and metadata into a compact, binary-encoded intermediate representation (IR).
+This structure-preserving design facilitates compatibility across heterogeneous hardware and execution environments, while also enabling fast serialization and deserialization.
+Figure 4: Architecture of the compilation module with lightweight data structures and module-level transformations.
+As illustrated in Figure 4, the IR undergoes multi-level compression, including gate dictionary encoding, instruction fusion at the bitstream level, and entropy-aware payload optimization.
+These strategies collectively reduce the size of the transmitted data by up to 80% compared to text-based formats such as OpenQASM, thereby enhancing throughput and minimizing latency in cloud-device quantum computing pipelines.
+Optional encryption and integrity verification mechanisms—via hashing and secure payload encapsulation—are incorporated to safeguard against tampering in multi-tenant or federated quantum systems.
+(2) High-Performance Quantum Circuit Compilation
+A customized data structure design method tailored to the characteristics of the quantum compilation workflow is proposed to address the issue of redundant intermediate information management in instruction reordering, topology-constrained mapping, and gate fusion optimization.
+The core of this approach lies in the construction of the quantum circuit’s directed acyclic graph (DAG), where lightweight operation units and contextual annotation mechanisms are introduced.
+Logical gate operations are organized into pipelined fragments with localized context, enabling structured phase buffers and operation dependency chains to support efficient access and dynamic updates of critical dependencies throughout gate-level optimization, scheduling, and topology mapping.
+This method significantly enhances compilation speed while preserving compilation quality, thereby meeting the rapid deployment requirements for large-scale quantum systems with 1000+ qubits.
+Figure 5: Architecture of the compilation module with lightweight data structures and module-level transformations.
+The Figure 5 introduces a contextual annotation mechanism that segments the circuit into pipelined fragments with localized context.
+Each segment is annotated with metadata such as segment identifiers, topological constraints, and temporal execution windows, enabling fine-grained control over optimization boundaries.
+These annotated fragments are managed within a stage buffer, which acts as an optimization context pool for subsequent compilation stages.
+A dependency-linked list structure is maintained to support cross-segment dependency tracking.
+This linked representation allows for low-latency traversal and update of gate dependencies during optimization.
+The integrated design of the DAG, stage buffer, and dependency tracking supports structured construction of operation dependency chains, which facilitates efficient updates and localized scheduling decisions throughout various optimization passes.
+The optimization layer leverages this structured data representation to enable a sequence of specialized transformations, including:
+  • Pipeline-stage layout optimization, which improves spatial coherence of gate execution,
+  • Parallel gate fusion, which merges compatible gates across segments to reduce circuit depth,
+  • Constraint-driven topology mapping, which ensures hardware compliance, and
+  • Dependency-aware gate translation, which refines gate translating under strict causal constraints.
+This architecture demonstrates significant improvements in compilation throughput, supporting compilation for systems with 1000+ qubits.
+By integrating operation structure, temporal context, and dependency linkage into a unified data-driven framework, this approach not only accelerates the compilation pipeline but also ensures high-quality output circuits that are compliant with the constraints of realistic NISQ hardware.
+III-D1 Gate Fusion
+Gate fusion is performed in a two-stage iterative pipeline.
+First, adjacent two-qubit gates are fused in parallel across pipeline segments to form localized composite gate blocks.
+Each fused block is then analyzed for potential decomposition into a tensor product of single-qubit gates.
+If such decomposition is feasible, the resulting single-qubit gates are merged into the surrounding context, enabling further compression of circuit depth.
+This fusion–decomposition–merging loop, driven by explicit dependency tracking and localized segment annotations, yields a structurally optimized two-qubit gate layer that significantly reduces overall circuit depth.
+III-D2 Layout Optimization
+Layout optimization begins with a subgraph reconstruction algorithm that attempts to map logical circuit substructures onto hardware topologies within a bounded runtime.
+This mapping process considers spatial gate alignment, dependency constraints, and SWAP minimization.
+If an exact subgraph match is not found within the allowed time frame, the compiler switches to a noise-aware mapping strategy.
+In this fallback mode, qubit blocks are selected based on backend calibration data, prioritizing hardware regions with higher gate fidelity to improve the robustness and execution fidelity of the final circuit.
+III-D3 Topology Mapping
+Topology mapping is built upon the LightSabre algorithm, which offers superior scalability for deep circuits and large hardware graphs.
+To improve mapping efficiency, a lightweight data structure is designed to represent qubit connectivity and cost metrics.
+This structure maintains sparse neighborhood graphs augmented with local path caches and dynamic edge costs, allowing fast constraint evaluation and route selection without the overhead of global matrix traversal.
+The result is a high-throughput mapping process that scales effectively to devices with 1000+ qubits.
+III-D4 Gate Translation
+Gate translation enables backend-agnostic circuit generation by converting logical gate sets into native hardware instructions.
+A graph-based dependency model is used to capture gate relationships, allowing flexible decomposition of complex operations into target-compatible primitives.
+To mitigate the depth overhead introduced by redundant single-qubit gates—often a side effect of two-qubit gate decomposition—the compiler adopts a fused translation–optimization strategy, where gate synthesis and simplification are co-executed in a single pass.
+For large-scale quantum programs, this process is further accelerated via parallelism, allowing multiple gate segments to be translated and optimized concurrently.
+(a) OriginIRs’ Encode Time
+(b) OriginIRs’ Decode Time
+(c) OriginIRs’ Post-encoding Size
+(d) OriginBIS’ Encode Time
+(e) OriginBIS’ Decode Time
+(f) OriginBIS’ Post-encoding Size
+Figure 6: For Circuit Count
+IV Experiments
+IV-A Performance Experiments of OriginBIS
+QPanda3 has specifically designed OriginBIS for quantum program transmission.
+A typical application scenario involves users designing quantum programs on local devices and executing them using abundant computing resources on quantum clouds.
+In this scenario, quantum programs exist as memory objects in high-level programming languages on the source device.
+After transmission, they exist as directly executable machine instructions on the target device.
+This scenario comprises three stages: firstly, the source device converts the quantum program in the form of a high-level programming language memory object into an intermediate representation of the quantum program; secondly, the intermediate representation is transmitted to the target device using a communication link; and finally, the target device converts the intermediate representation into directly executable machine instructions.
+It should be noted that the scope of the concept of quantum program transmission may vary depending on specific business requirements.
+If the business only requires the quantum program on the source device to appear on the target device, then transmitting the data storing the intermediate representation of the quantum program from the source device to the target device completes the quantum program transmission, which only includes the second stage.
+If the business requires quantum program transmission from a memory object on the source device to executable machine instructions on the target device, then the quantum program transmission should encompass all three stages.
+For convenience of description, in the experiments of this paper, the quantum program transmission process refers to the second stage, the first stage is termed the Encode stage, and the third stage is termed the Decode stage.
+The experiments in this paper primarily focus on testing and comparing the performance of OriginBIS across these three stages.
+This paper tests the performance of OriginBIS in quantum program transmission and quantum program format conversion through multiple comparative experiments.
+The experimental results demonstrate that OriginBIS significantly improves the processing efficiency of the corresponding processes.
+(a) OriginIRs’ Encode Time
+(b) OriginIRs’ Decode Time
+(c) OriginIRs’ Post-encoding Size
+(d) OriginBIS’ Encode Time
+(e) OriginBIS’ Decode Time
+(f) OriginBIS’ Post-encoding Size
+Figure 7: For Circuit Depth
+(a) OriginIRs’ Encode Time
+(b) OriginIRs’ Decode Time
+(c) OriginIRs’ Post-encoding Size
+(d) OriginBIS’ Encode Time
+(e) OriginBIS’ Decode Time
+(f) OriginBIS’ Post-encoding Size
+Figure 8: For Qubit Count
+IV-A1 Evaluation Metrics
+(1) Encode Time and Decode Time
+For the Encode and Decode stages, the speed at which the device completes the entire processing operations of the respective stages reflects the processing efficiency of the corresponding processes.
+Considering that it is difficult to quantify the amount of processing operations performed by the device into specific numerical values, this paper uses the time taken to complete the same conversion task to reflect the impact of the intermediate representation on processing efficiency during the Encode and Decode stages.
+Specifically, for the same conversion task, the longer the time taken, the lower the processing efficiency; the shorter the time taken, the higher the processing efficiency.
+(2) Post-encoding Size
+Unlike the Encode and Decode stages, where time is used as a metric, this paper employs the data size of the quantum program after encoding (Post-encoding Size) to measure the impact of the intermediate representation on transmission efficiency during the quantum program transmission stage.
+The quantum program transmission stage constitutes a classical communication process, where latency is a crucial performance evaluation metric.
+However, in practical communication environments, latency is influenced by various factors, including the bandwidth and distance of physical communication links and devices, as well as the classical network protocols used for data transmission.
+Quantum program transmission primarily involves transmitting the intermediate representation of the quantum program as pure data.
+The data size of the intermediate representation is the only component of the quantum program intermediate representation that is directly related to communication latency, excluding the influences of physical communication links, devices, and classical network protocols.
+Specifically, for the same quantum program transmission task, a larger Post-encoding Size results in lower transmission efficiency, while a smaller Post-encoding Size leads to higher transmission efficiency.
+IV-A2 Experimental Setup
+This paper conducts comparative experiments from three aspects: the number of circuits, circuit depth, and the number of qubits.
+Batch transmission of a large number of circuits is one of the real and essential practical application requirements.
+Circuit depth is a crucial factor that affects the execution results of quantum circuits.
+The number of qubits measures the amount of quantum computing resources occupied by quantum circuits.
+These three metrics serve as important references for users when submitting quantum programs and for devices when processing quantum programs for execution.
+The experiments in this paper are carried out using the scheme shown in Table I.
+For the experiment on the number of quantum circuits, the number of quantum circuits is taken as the independent variable, while the circuit depth is controlled at a fixed value of 500 and the number of qubits is controlled at a fixed value of 72.
+For the experiment on circuit depth, the circuit depth is taken as the independent variable, with the number of circuits controlled at a fixed value of 500 and the number of qubits controlled at a fixed value of 72.
+For the experiment on the number of qubits, the number of qubits is taken as the independent variable, with the circuit depth controlled at a fixed value of 500 and the number of circuits controlled at a fixed value of 500.
+It should be noted that the parameters related to circuits are not limited to the number of circuits, circuit depth, and the number of qubits.
+Moreover, there may be numerical constraints among these parameters.
+For example, when the number of quantum gates is fixed, there is a certain negative correlation between circuit depth and the number of qubits.
+To minimize the influence of other parameters and the relationships among them, this paper adopts a randomized technical approach, generating random circuits based on the conventions shown in Table I to test the performance of quantum program intermediate representations in quantum program transmission.
+Relevant experiments are conducted on OriginBIS, OriginIR, and QASM to compare the quantum program transmission performance of these intermediate representations.
+TABLE I: Experimental Setup For Quantum Program Transmission
+Experiment Circuit Count Circuit Depth Qubit Count
+For Circuit Count Independent Variable 500 72
+For Circuit Depth 500 Independent Variable 72
+For Qubit Count 500 500 Independent Variable
+IV-A3 Experimental Results and Analysis
+(1) For Circuit Count
+The experimental results for the number of circuits are shown in Figure 6.
+The grouped columns chart in subplots (a), (b), and (c) demonstrates that OriginBIS is significantly smaller than OriginIR and QASM in terms of Encode Time, Decode Time, and Post-encoding Size.
+This indicates that in these experiments, OriginBIS outperforms OriginIR and QASM in the efficiency of quantum program transmission.
+Through simple numerical estimation, it can be found that for Encode Time, OriginIR is approximately 17 times that of OriginBIS, and QASM is approximately 80 times that of OriginBIS; for Decode Time, OriginIR is approximately 20 times that of OriginBIS, and QASM is approximately 40 times that of OriginBIS; for Post-encoding Size, OriginIR is approximately 3.5 times that of OriginBIS, and QASM is approximately 6.5 times that of OriginBIS.
+The point-line charts in subplots (d), (e), and (f) show that for OriginBIS, Encode Time, Decode Time, and Post-encoding Size increase with the growth in the number of circuits.
+(2) For Circuit Depth
+The experimental results regarding circuit depth are presented in Figure 7.
+The grouped columns chart in subplots (a), (b), and (c) demonstrate that OriginBIS exhibits significantly lower values in terms of Encode Time, Decode Time, and Post-encoding Size compared to both OriginIR and QASM.
+This indicates that in these experiments, OriginBIS outperforms OriginIR and QASM in terms of efficiency for quantum program transmission.
+Upon simple numerical estimation, it can be observed that for Encode Time, OriginIR is approximately 18 times that of OriginBIS, while QASM is about 88 times; for Decode Time, OriginIR is roughly 21 times that of OriginBIS, and QASM is approximately 39 times; for Post-encoding Size, OriginIR is around 3.5 times that of OriginBIS, and QASM is about 6.5 times.
+The point-line charts in subplots (d), (e), and (f) reveal that for OriginBIS, Encode Time, Decode Time, and Post-encoding Size increase with the growth of circuit depth.
+(a) OriginIRs’ Encode Time
+(b) OriginIRs’ Decode Time
+(c) OriginIRs’ Post-encoding Size
+(d) OriginBIS’ Encode Time
+(e) OriginBIS’ Decode Time
+(f) OriginBIS’ Post-encoding Size
+Figure 9: For Gate Count
+(3) For Qubit Count
+The experimental results concerning the number of qubits are illustrated in Figure 8.
+The grouped columns chart in subplots (a), (b), and (c) indicates that OriginBIS exhibits significantly lower Encode Time and Decode Time compared to both OriginIR and QASM.
+Additionally, for qubit counts of 50 or more, OriginBIS also demonstrates lower Post-encoding Time compared to OriginIR and QASM.
+This suggests that in the majority of the samples tested, OriginBIS outperforms OriginIR and QASM in terms of efficiency for quantum program transmission.
+Excluding the 90-qubit data in subplot (b) and the 40-bit data in subplot (c), upon simple numerical estimation, it can be observed that for Encode Time, OriginIR is approximately 21 times that of OriginBIS, while QASM is about 93 times; for Decode Time, OriginIR is roughly 20 times that of OriginBIS, and QASM is approximately 39 times; for Post-encoding Size, OriginIR is around 3.5 times that of OriginBIS, and QASM is about 6.5 times.
+The point-line charts in subplots (d), (e), and (f) reveal that for OriginBIS, there is a trend of increasing Encode Time and Decode Time with the growth in the number of qubits, albeit with considerable data fluctuation.
+However, Post-encoding Time consistently increases with the increase in the number of qubits.
+(4) Gate Count
+During the experiments examining circuit count, circuit depth, and qubit count, it was observed that the number of quantum gates varies with changes in these parameters.
+Figure 10 illustrates the trend in the number of quantum gates, demonstrating a strong positive linear correlation with the value of the third parameter when the values of any two of these three quantities are held constant.
+The number of gates is also an important metric for evaluating quantum programs.
+By combining the gate count data from Figure 10 with the corresponding Encode Time, Decode Time, and Post-encoding Size data, the results presented in Figure 9 were obtained.
+Subplots (a), (b), and (c) indicate that the Encode Time, Decode Time, and Post-encoding Size for OriginBIS are significantly lower than those for OriginIR and QASM.
+Subplots (d), (e), and (f) reveal that there is a trend of increasing Encode Time and Decode Time with the growth in the number of quantum gates, albeit with noticeable data fluctuations.
+However, Post-encoding Time increases with the increase in the number of quantum gates, exhibiting a strong linear characteristic.
+The experimental results concerning the number of qubits are illustrated in Figure 8.
+The grouped columns chart in subplots (a), (b), and (c) indicates that OriginBIS exhibits significantly lower Encode Time and Decode Time compared to both OriginIR and QASM.
+Additionally, for qubit counts of 50 or more, OriginBIS also demonstrates lower Post-encoding Time compared to OriginIR and QASM.
+This suggests that in the majority of the samples tested, OriginBIS outperforms OriginIR and QASM in terms of efficiency for quantum program transmission.
+Excluding the 90-qubit data in subplot (b) and the 40-qubit data in subplot (c), upon simple numerical estimation, it can be observed that for Encode Time, OriginIR is approximately 21 times that of OriginBIS, while QASM is about 93 times; for Decode Time, OriginIR is roughly 20 times that of OriginBIS, and QASM is approximately 39 times; for Post-encoding Size, OriginIR is around 3.5 times that of OriginBIS, and QASM is about 6.5 times.
+The point-line charts in subplots (d), (e), and (f) reveal that for OriginBIS, there is a trend of increasing Encode Time and Decode Time with the growth in the number of qubits, albeit with considerable data fluctuation.
+However, Post-encoding Time consistently increases with the increase in the number of qubits.
+Figure 10: Gate Count for Circuit Count, Circuit Depth and QBit Count
+Figure 11: Experimental Results About Circuit Compilation
+IV-B Compilation Efficiency Comparison Experiment
+IV-B1 Experiment Setup
+In this paper, a comparative experiment related to compilation efficiency was conducted on multiple mainstream quantum programming frameworks based on Benchpress.
+Benchpress is an evaluation suite for multi-quantum computing software development kits, containing thousands of test cases.
+It allows for the uniform testing of multiple quantum software packages across various performance and functional indicators, with the results reflecting the cost of processing quantum circuits on quantum computing devices using different software packages.
+All relevant experiments for each software package mentioned in the Benchpress literature were replicated in this study, with the addition of tests for QPanda3.
+All software packages tested are listed in Table II.
+It should be noted that a timeout limit of 180 seconds was imposed for executing each test case.
+The testing of these software packages was conducted in the same software and hardware environments. The environmental parameters used in our experiments are presented in Table III.
+TABLE II: SDKs
+SDK Version
+amazon-braket-sdk(braket) 1.91.0
+bqskit 1.2.0
+cirq 1.4.1
+pytket(tket) 2.0.1
+qiskit 2.0.1
+qiskit_ibm_transpiler 0.39.0
+pystaq 3.5
+pyqpanda3 0.3.0
+TABLE III: Software and Hardware Environments
+Type Info
+CPU Intel(R) Xeon(R) Platinum 8336C CPU @ 2.30GHz
+Memory 2.0TB
+Operating System Ubuntu 20.04.5 LTS
+Python 3.11.11
+IV-B2 Evaluation Metrics
+This paper adopts the evaluation metrics introduced by Benchpress. They are Standard Pytest Output Type, circuit construction time, manipulate time, and circuit compilation time.
+IV-B3 Experimental Results And Analysis
+TABLE IV: Experimental Results Evaluated Based on The Criterion of Standard Pytest Output Type
+SDK Passed Failed XFailed Skipped Total
+braket 7 2 0 1057 1066
+cirq 10 2 0 1054 1066
+qiskit 1044 0 0 22 1066
+QPanda3 1016 28 0 22 1066
+In our experiments, the results of the test_status_counts are presented in Table IV.
+It was observed that many SDKs failed to satisfy the requirements of numerous test cases in Benchpress.
+Specifically, bqskit was only capable of performing tests related to circuit construction.
+pytket and pystaq exhibited compatibility issues with the environments of other SDKs, making it impossible to conduct fair testing and comparison using a unified test suite and environment.
+Furthermore, qiskit_ibm_transpiler demonstrated strong dependency on the IBM computing platform, rendering it untestable under our experimental environment.
+As evident from Table IV, both braket and cirq skipped a significant number of test cases.
+According to the Benchpress literature, qiskit skipped 22 test cases and passed 1,044 test cases.
+Tabel IV indicates that our experiments actually tested multiple SDKs using 1,066 test cases.
+The performance of qiskit in our experiments was consistent with the reports in the Benchpress literature.
+QPanda successfully passed $95.3\%$ of all test cases, with the number of skipped cases equal to that of qiskit.
+This table showcases the excellent versatility of QPanda3 within the unified test suite, Benchpress.
+The performance of various SDKs in circuit construction is illustrated in Figure 12.
+It can be observed that QPanda3, qiskit, and cirq have successfully passed all seven test cases.
+For the four test cases, namely test_multi_control_circuit, test_bigint_qasm2_import, test_param_circSU2_100_build, and test_clifford_build, QPanda3 demonstrated the shortest execution time.
+In the case of the three test cases test_QV100_qasm2_import, test_param_circSU2_100_build, and test_QV100_buildQPanda3 took longer than qiskit but significantly less time compared to cirq, braket, and bqskit.
+For the test case test_DTC100_set_build, QPanda3’s execution time was shorter than that of qiskit but longer than cirq’s, ranking as the second shortest.
+This figure highlights the advantages of QPanda3 in circuit construction.
+Figure 12: Experimental Results About Circuit Construction(shorter is better)
+Figure 13: Experimental Results About Circuit Manipulation(shorter is better)
+The Benchpress literature indicates that Manipulate is a crucial metric for evaluating SDKs.
+As illustrated in Figure 13, only QPanda3 and qiskit have successfully passed all four test cases.
+In the test_QV100_basis_change test case, QPanda3 demonstrated the shortest execution time.
+For the three test cases: test_random_clifford_decompose, test_multi_control_decompose, and test_DTC100_twirling, QPanda3’s execution time slightly exceeded that of qiskit, with a minimal difference.
+This figure suggests that QPanda3 holds an advantage in terms of the Manipulate metric.
+In our experiments, only QPanda3 and qiskit passed the majority of the test cases related to compilation.
+We have plotted the corresponding experimental results into multiple scatter plots, as shown in Figure 11, which illustrate the differences between QPanda3 and qiskit in terms of 2Q Gate Count, 2Q Gate Depth, and compilation time for various topological structures.
+In each plot, the blue dashed line represents the point where the values for QPanda3 and qiskit are equal for the corresponding metric.
+Points above the blue dashed line indicate that the corresponding value for QPanda3 is less than that for qiskit.
+As observed from Figure 11, for both 2Q Gate Count and 2Q Gate Depth, the majority of the scatter points are densely distributed along or very close to the blue dashed line, suggesting that the performance of QPanda3 and qiskit is essentially the same.
+However, regarding compilation time, all scatter points in the four subplots of Figure 11 are located above the blue dashed line, indicating that QPanda3 exhibits higher compilation efficiency than qiskit.
+This characteristic is not limited by topological structure and is effective for a large number of different circuits.
+Additionally, it can be observed that only a very few test cases result in compilation times exceeding 10 seconds when using QPanda3.
+These observations collectively demonstrate the high compilation efficiency of QPanda3.
+IV-C Performance Comparison Experiment for Gradient Computation in Variational Quantum Circuits
+Analysis of Gradient Computation Performance and Circuit Depth in 12-Qubit Systems
+Figure 14: Circuit Depth Impact (Fixed 12 Qubits)
+The performance analysis of gradient computation in 12-qubit systems corresponding to Figure 14 demonstrates that quantum framework selection critically determines computational efficiency.
+PyQPanda3 exhibits significant time advantages, requiring only 0.0088 seconds average time at 2-layer circuit depth compared to Qiskit’s 0.232 seconds - a 26-fold difference.
+When depth increases to 6 layers, PyQPanda3’s time merely grows to 0.0241 seconds (2.7x increase), substantially lower than Qiskit’s 0.6097 seconds (2.6x) and DeepQuantum’s 0.1680 seconds (2.9x).
+Notably, frameworks maintain a consistent performance hierarchy: PyQPanda3 consistently delivers optimal performance, with computation times 40%-50% lower than second-ranked PennyLane-Lightning (0.0241s vs 0.0362s at 6 layers), while Qiskit persistently trails.
+During depth scaling, PyQPanda3 shows the smallest absolute time increment (+0.0153s), equivalent to just 4% of Qiskit’s increment (+0.3777s), enabling 97.6% time savings in deep circuits.
+Experimental data confirms PyQPanda3’s superior depth-scaling stability, with significantly lower time-complexity growth rates than comparable frameworks, providing an effective computational efficiency solution for deep quantum circuits.
+Comparative Analysis for Fixed Qubit Count and Circuit Depth
+Figure 15: Framework Comparison 8qubits 4layers
+Figure 16: Framework Comparison 16qubits 6layers
+Based on the comparative analysis of Figure 15 and Figure 16, a significant and stable statistical pattern emerges in gradient computation times across quantum frameworks.
+Under 8-qubit/4-layer conditions, PyQPanda3 demonstrates exceptional performance (0.0031s), requiring only 21.7% of the time needed by second-ranked PennyLane-Lightning (0.0143s), while Qiskit’s duration (0.3090s) reaches 99.7x that of PyQPanda3.
+When scaling to 16-qubit/6-layer configurations, the performance ranking remains strictly consistent: PyQPanda3 (0.1116s) maintains the lowest time, outperforming PennyLane-Lightning (0.2134s) by 47.7%, with Qiskit (2.0085s) requiring 18x more time than PyQPanda3.
+Notably, the performance differences exhibit systematic hierarchical characteristics: PyQPanda3 and PennyLane-Lightning consistently form an efficient cluster, collectively accounting for 16.1% and 7.5% of total computation time under the two respective scales; DeepQuantum and PennyLane-Default comprise a mid-tier cluster; while Qiskit independently constitutes an inefficient unit, dominating 82.9% (8-qubit) and 68.2% (16-qubit) of total time.
+As computational scale increases, PyQPanda3’s absolute advantage over Qiskit expands from 0.3059s to 1.8969s, though its relative advantage multiplier decreases from 99x to 18x, indicating nonlinear evolution of performance differentials with scaling.
+These statistical patterns confirm that quantum framework performance disparities maintain intrinsic stability regardless of computational scale.
+Time Analysis of Gradient Computation for Fixed 4-Layer Quantum Circuits
+Figure 17: Qubit Scalability 4layers
+Analysis of gradient computation time statistics for fixed 4-layer quantum circuits reveals the following patterns: As qubit count increases from 4 to 16, all frameworks exhibit exponential time growth (manifesting as linear trends in logarithmic coordinates), yet with significantly different growth slopes.
+PyQPanda3 demonstrates the mildest growth (approximately 100x increase from 4 to 16 qubits), while Qiskit shows the steepest escalation (approximately 10x increase).
+PennyLane-Default and DeepQuantum exhibit intermediate growth slopes with stable differentials, maintaining a consistent time ratio of $\approx$1:1.2 throughout the scaling process.
+The performance differentials between frameworks exhibit dual-scale dependency: Regarding absolute differences, the maximum time gap continuously widens with increasing qubits (from 0.099s between PyQPanda3 and Qiskit at 4 qubits to 0.9s at 16 qubits); whereas in relative terms, the advantage of efficient frameworks diminishes with scale (the PyQPanda3/Qiskit time ratio decreases from 1:100 at 4 qubits to 1:10 at 16 qubits).
+Comparison with PennyLane-Lightning
+Figure 18: Speedup Heatmap vs Pennylane-lightning
+Based on the analysis of the speedup heatmap 18, QPanda3 exhibits significant systematic performance advantages over Pennylane Lightning.
+In shallow small-scale quantum circuits (depth $\leq$6 layers, qubits $\leq$8), QPanda3 demonstrates prominent acceleration effects, reaching up to 26.0x (2-layer/4-qubit configuration).
+However, this advantage exhibits dual-decay characteristics with parameter variations: when fixing qubit count, the speedup ratio decreases sharply with increasing depth (e.g., for 4-qubit circuits: from 26.0 at 2 layers to 1.2 at 12 layers); when fixing depth, the speedup ratio decays progressively with qubit scaling (e.g., at 2-layer depth: from 26.0 at 4 qubits to 6.5 at 16 qubits).
+The impact of circuit depth on acceleration effects is particularly significant: a depth of 8 layers constitutes a critical turning point, where speedup ratios are generally $>$4.0 below this threshold (e.g., 5.3 for 6-layer/4-qubit circuits) but typically $<$2.0 above it (e.g., only 1.6 for 10-layer/4-qubit circuits).
+A qubit count of 12 emerges as another demarcation point-below this scale, speedup ratios exhibit drastic fluctuations (maximum 73% decrease from 4 to 8 qubits), while above this scale they tend to stabilize ($<$5% variation between 14-16 qubits).
+The spatial distribution of performance advantages exhibits a clear hierarchical structure: strong acceleration zones ($>$5.0) are concentrated in the upper-left triangular region (depth $\leq$6 layers & qubits $\leq$8), weak acceleration zones (1.2-2.0) occupy the lower-right triangular region (depth $\geq$10 layers or qubits $\geq$14), and transition zones (1.6-4.9) are distributed along the diagonal.
+This distribution pattern indicates that the inhibitory effect of circuit depth on acceleration surpasses that of qubit scale, converging to a 1.2-2.0x performance gap between the two frameworks when depth $\geq$10 layers or qubits $\geq$14.
+Comparison with Qiskit
+Figure 19: Speedup Heatmap vs Qiskit
+Analysis based on heatmap 19 data indicates that QPanda3 demonstrates significant performance advantages compared to Qiskit, with its acceleration effects exhibiting clear parameter-dependent patterns.
+In medium-depth small-scale quantum circuits (48 layers depth, 610 qubits), QPanda3 achieves super-strong acceleration, peaking at 157.3 times (6 layers depth/6 qubits).
+This advantage shows asymmetric attenuation with circuit parameter changes: the attenuation intensity caused by increasing depth (up to 83%) is significantly higher than the impact of increasing qubit count (up to 45%).
+Notably, a depth of 8 layers constitutes a critical threshold-below this value, the speedup ratio is generally $>$68.6 times (e.g., 99.2 times for 6 layers depth/8 qubits), while above this value it sharply decreases to $<$27.8 times (e.g., 24.6 times for 10 layers depth/4 qubits).
+The impact of qubit scale on acceleration exhibits phased characteristics: when the number of qubits $\leq$10, the speedup ratio fluctuates drastically with increasing depth (e.g., from 70.4 to 24.6 for 4-qubit circuits when depth increases from 2 to 10 layers); whereas when qubits $\geq$12, the speedup ratio shows blunted responses to parameter variations (e.g., 16-qubit circuits maintain only 17.118.0x acceleration across different depths).
+This differential response forms three characteristic spatial distribution zones: super-strong acceleration zone ($>$80x) concentrates in the rectangular region with depth $\leq$6 layers and qubits $\leq$10; strong acceleration zone (40-99x) extends to depth 8 layers / qubits 14; stable acceleration zone (17-28x) covers the lower-right triangular region where depth $\geq$10 layers or qubits $\geq$12.
+Even under deep-layer large-qubit-scale conditions (16 layers depth/16 qubits), QPanda3 maintains stable acceleration advantages exceeding 17x, with fluctuation amplitude $<$20%.
+This phenomenon indicates that while circuit depth exerts stronger inhibitory effects on acceleration than qubit scale, QPanda3 consistently demonstrates significant performance advantages across the entire parameter space, particularly exhibiting groundbreaking acceleration capabilities in shallow-to-medium quantum circuits.
+V Conclusion
+This paper focuses on efficient compilation and execution in quantum computing, introducing OriginIR and OriginBIS as intermediate representations to enhance programmability, transmission efficiency, and execution performance.
+Experimental results demonstrate that OriginBIS significantly outperforms OpenQASM 2.0 in encoding and decoding speed as well as information capacity.
+Additionally, QPanda3 surpasses Qiskit in quantum circuit construction, operation execution, and compilation speed, exhibiting exceptional performance improvements, particularly in large-scale quantum circuit compilation tasks.
+QPanda3 provides fundamental and efficient functionalities related to variational quantum circuits, offering robust support for quantum-classical hybrid computation.
+These advancements lay a solid foundation for the engineering applications of future quantum computing.
+VI Future Work
+The quantum programming framework QPanda3 is poised to enhance its capabilities for future ”quantum-HPC-AI” integrated computing by focusing on several key advancements.
+Building on its existing strength of compiling quantum circuits 14.97 times faster than Qiskit in Benchpress tests, QPanda3 will develop hybrid task schedulers to enable seamless coordination between quantum computations and classical supercomputing resources.
+It aims to integrate AI-driven optimization tools, such as AutoML modules, to automatically reduce quantum gate counts and circuit depths while mitigating noise impacts.
+The framework will expand its hybrid programming interfaces to support classical HPC technologies like MPI and CUDA, facilitating efficient quantum-classical algorithm interoperability.
+Additionally, QPanda3 plans to implement a unified resource management platform for dynamic allocation of quantum hardware, supercomputing clusters, and AI accelerators.
+To strengthen ecosystem integration, it will promote standardization through open-source collaboration, enabling deeper compatibility with classical AI libraries (e.g., TensorFlow) while maintaining its low-latency compilation advantages.
+These upgrades will position QPanda3 as a core enabler for cross-paradigm applications in finance, drug discovery, and materials science.
+Appendix A Representation and Transmission of Quantum Programs
+A-A User-Friendly Intermediate Representation-OriginIR
 The significant role of high-level intermediate representations of quantum programs in the quantum software stack cannot be overstated.
 QPanda3 employs OriginIR as a lower-level representation of quantum programs compared to high-level programming languages.
-IV-A1 Compatibility with QPanda2
+A-A1 Compatibility with QPanda2
 QPanda3 largely retains all the features of OriginIR from QPanda2, while also introducing appropriate extensions.
@@
 OriginIR enables seamless portability of quantum programs between QPanda2 and QPanda3.
-IV-A2 Geared Towards Researchers
+A-A2 Geared Towards Researchers
 (1) Complexity and Conversion
@@
 The improvement in readability will greatly facilitate researchers in conducting quantum computing-related research based on OriginIR.
-IV-B Intermediate Representation for Quantum Program Transmission - OriginBIS
+A-B Intermediate Representation for Quantum Program Transmission - OriginBIS
 Quantum program transmission is a crucial component of the quantum computing software stack.
@@
 It applies limited compression to the quantum program.
-V Optimized Compilation of Quantum Circuits
+Appendix B Optimized Compilation of Quantum Circuits
 The compilation process of quantum circuits involves the transformation from high-level abstract descriptions to hardware-executable instructions.
@@
 The practical application of quantum software is inseparable from efficient quantum circuit compilation schemes.
-V-A Transpiler of QPanda3
+B-A Transpiler of QPanda3
 QPanda3 features a device-oriented modern quantum circuit transpiler.
@@
 Whether it is the merging of quantum gates, simplification of circuits, or allocation of resources, QPanda3 can accomplish these tasks at a faster pace, thereby improving overall compilation efficiency.
-V-B Starting with an Example
+B-B Starting with an Example
 This section demonstrates the use of the built-in transpiler in QPanda3 through a simple example.
@@
 The generate_topology function is used for randomly generating topology data.
-from pyqpanda3.core import QProg, draw_qprog, RZ, X1, CZ, SWAP, set_print_options
-from pyqpanda3.transpilation import Transpiler, generate_topology
+1from pyqpanda3.core import QProg, draw_qprog, RZ, X1, CZ, SWAP, set_print_options
+2from pyqpanda3.transpilation import Transpiler, generate_topology
+Listing 1: An example Python program that evaluates the properties of quantum states.
 (2) Constructing Quantum Logic Circuits
-prog_1 = QProg()
-prog_1 << RZ(3,1.57) << RZ(4,3.14) << RZ(5,6.28) << CZ(3,4) << CZ(4,5) << CZ(3,5)
-set_print_options(2)
-[…]
-The output is
+1prog_1 = QProg()
+2prog_1 << RZ(3,1.57) << RZ(4,3.14) << RZ(5,6.28) << CZ(3,4) << CZ(4,5) << CZ(3,5)
+3set_print_options(2)
+Listing 2: An example Python program that evaluates the properties of quantum states.
+The output is show as Fig 21
 Figure 21: Result of Constructing Quantum Logic Circuits
@@
 Users can also generate similar data using the generate_topology interface.
-topo = [[0,1],[1,2]]
+1topo = [[0,1],[1,2]]
+Listing 3: An example Python program that evaluates the properties of quantum states.
 (4) Compiling with Different Optimization Levels
@@
 In the subsequent sections, some key output results of this program will be extracted and analyzed.
-transpiler = Transpiler()
-prog_level_0 = transpiler.transpile(prog_1, topo, {}, 0)
+1transpiler = Transpiler()
+2prog_level_0 = transpiler.transpile(prog_1, topo, {}, 0)
 […]
-V-C Qubit Mapping and Routing
+Listing 4: An example Python program that evaluates the properties of quantum states.
+B-C Qubit Mapping and Routing
 (1) Qubit Mapping
@@
 When compiling this quantum logic circuit, QPanda3 may not always use this specific layout and routing scheme, as other layout and routing solutions can still satisfy the target requirements.
-V-D Optimization
+B-D Optimization
 QPanda3 adopts a strategy that combines local and global optimization.
@@
 Global optimization, on the other hand, considers the optimization of the entire quantum circuit by adjusting the order of quantum gates, reallocating resources, and other means to further improve the execution efficiency of the circuit.
-V-D1 Sabre
+B-D1 Sabre
 Sabre[29] is an advanced quantum circuit compilation algorithm.
@@
 The Sabre algorithm is closely integrated with the characteristics of quantum hardware, enabling the compiled circuit to better adapt to the hardware’s execution environment, thereby ensuring efficient execution of the algorithm.
-V-D2 Independent Optimization Steps
+B-D2 Independent Optimization Steps
 This step involves isolating certain optimization strategies as a standalone procedure, intended for pre-layout and post-routing optimization.
@@
 Figure 26: Eliminating Redundant Gates
-V-D3 Pre-placement Optimization and Post-routing Optimization
+B-D3 Pre-placement Optimization and Post-routing Optimization
 Prior to the placement stage, QPanda3 conducts a series of preprocessing optimizations on the quantum circuit.
@@
 After the routing stage is completed, QPanda3 further optimizes the circuit to enhance its execution efficiency even more.
-VI Device-Based Quantum Program Analysis
+Appendix C Device-Based Quantum Program Analysis
 Whether a quantum program can be executed efficiently on a specific computing device and yield desirable results determines its suitability for performance-sensitive tasks and its advantage in terms of computational resource costs.
@@
 This section focuses on the latter, introducing the tools provided by QPanda3 for device-based quantum program analysis.
-VI-A Runtime State Information and Quantum Program Analysis
+C-A Runtime State Information and Quantum Program Analysis
 In classical computing, runtime state information of various hardware components can be directly obtained through dedicated counters, registers, and debug-level interfaces during program execution.
@@
 Currently, QPanda3 offers quantum circuit analysis tools tailored to device performance and quantum program performance analysis tools based on workflow and device information.
-VI-B Quantum Circuit Analysis for Device Performance
-VI-B1 Overview
+C-B Quantum Circuit Analysis for Device Performance
+C-B1 Overview
 QPanda3 extends the quantum computing performance benchmarks to the quantum circuit analysis context.
@@
 Finally, the draw_circuit_features function is utilized to perform the analysis.
-from pyqpanda3.profiling import draw_circuit_features
-from pyqpanda3.core import QCircuit, H, CNOT, X
-circuit = QCircuit(4)
-circuit << H(0) << CNOT(0, 1) << CNOT(1, 2) << X(3)
-draw_circuit_features(circuit, True)
+1from pyqpanda3.profiling import draw_circuit_features
+2from pyqpanda3.core import QCircuit, H, CNOT, X
+3circuit = QCircuit(4)
+4circuit << H(0) << CNOT(0, 1) << CNOT(1, 2) << X(3)
+5draw_circuit_features(circuit, True)
+Listing 5: An example Python program that evaluates the properties of quantum states.
 The corresponding visualization results are shown in Figure 27.
 Figure 27: Performance Metrics of A Quantum Program
-VI-C Performance Analysis of Quantum Programs Based on Flow and Device Information
+C-C Performance Analysis of Quantum Programs Based on Flow and Device Information
 Inspired by classical program analysis methods, QPanda3 provides users with a tool for analyzing quantum programs based on flow and device information.
@@
 Then, centering around this example, we will introduce the differences between this feature of QPanda3 and classical program analysis, as well as the other software components of this functionality.
-VI-C1 An Example
+C-C1 An Example
 QPanda3 provides the functionality for performance analysis of quantum programs based on flow and device information through the draw_circuit_profile interface in the pyqpanda3.profiling package.
@@
 This information is passed as the second parameter to the draw_circuit_profile function in the form of a Python dictionary.
-from pyqpanda3.profiling import draw_circuit_profile
-from pyqpanda3.core import QCircuit, H, CNOT, X
-cir1 = QCircuit(4)
-cir1 << H(0) << CNOT(0, 1) << CNOT(1, 2) << X(3) << X(3)
-cir2 = QCircuit(4)
-cir2 << X(1) << cir1
-draw_circuit_profile(cir2, {’CNOT’: 200, ’H’: 60, ’X’: 30}, True)
+1from pyqpanda3.profiling import draw_circuit_profile
+2from pyqpanda3.core import QCircuit, H, CNOT, X
+3cir1 = QCircuit(4)
+4cir1 << H(0) << CNOT(0, 1) << CNOT(1, 2) << X(3) << X(3)
+5cir2 = QCircuit(4)
+6cir2 << X(1) << cir1
+7draw_circuit_profile(cir2, {’CNOT’: 200, ’H’: 60, ’X’: 30}, True)
+Listing 6: An example Python program that evaluates the properties of quantum states.
 In this code, the third parameter of the draw_circuit_profile function controls whether visualization is performed. The corresponding visualization result is shown in Figure 28.
 Figure 28: Profile of Circuit
-VI-C2 Comparison with Classical Program Analysis
+C-C2 Comparison with Classical Program Analysis
 Quantum sub-circuits bear similarities to called functions (subroutines) in classical programs.
@@
 Through flow-based and device-informed quantum program performance analysis, users can ascertain the resource usage of various quantum sub-circuits and quantum gates during execution, thereby adjusting the implementation schemes of corresponding sub-circuits accordingly.
-VI-C3 Output Analysis Results
+C-C3 Output Analysis Results
 The analysis results provided by draw_circuit_profile offer the frequency and time consumption percentage of each sub-circuit and quantum gate within the quantum circuit.
@@
 Figure 29: Gprof-Style Report
-(a) IRs’ Encode Time
-(b) IRs’ Decode Time
-(c) IRs’ Post-encoding Size
-(d) OriginBIS’ Encode Time
-(e) OriginBIS’ Decode Time
-(f) Visualization of the circuit
-Figure 14: For Circuit Count
-VII Experiments
-VII-A Performance Experiments of OriginBIS
-QPanda3 has specifically designed OriginBIS for quantum program transmission.
-A typical application scenario involves users designing quantum programs on local devices and executing them using abundant computing resources on quantum clouds.
-In this scenario, quantum programs exist as memory objects in high-level programming languages on the source device.
-After transmission, they exist as directly executable machine instructions on the target device.
-This scenario comprises three stages: firstly, the source device converts the quantum program in the form of a high-level programming language memory object into an intermediate representation of the quantum program; secondly, the intermediate representation is transmitted to the target device using a communication link; and finally, the target device converts the intermediate representation into directly executable machine instructions.
-It should be noted that the scope of the concept of quantum program transmission may vary depending on specific business requirements.
-If the business only requires the quantum program on the source device to appear on the target device, then transmitting the data storing the intermediate representation of the quantum program from the source device to the target device completes the quantum program transmission, which only includes the second stage.
-If the business requires quantum program transmission from a memory object on the source device to executable machine instructions on the target device, then the quantum program transmission should encompass all three stages.
-For convenience of description, in the experiments of this paper, the quantum program transmission process refers to the second stage, the first stage is termed the Encode stage, and the third stage is termed the Decode stage.
-The experiments in this paper primarily focus on testing and comparing the performance of OriginBIS across these three stages.
-This paper tests the performance of OriginBIS in quantum program transmission and quantum program format conversion through multiple comparative experiments.
-The experimental results demonstrate that OriginBIS significantly improves the processing efficiency of the corresponding processes.
-(a) IRs’ Encode Time
-(b) IRs’ Decode Time
-(c) IRs’ Post-encoding Size
-(d) OriginBIS’ Encode Time
-(e) OriginBIS’ Decode Time
-(f) OriginBIS’ Post-encoding Size
-Figure 15: For Circuit Depth
-(a) IRs’ Encode Time
-(b) IRs’ Decode Time
-(c) IRs’ Post-encoding Size
-(d) OriginBIS’ Encode Time
-(e) OriginBIS’ Decode Time
-(f) OriginBIS’ Post-encoding Size
-Figure 16: For Qubit Count
-VII-A1 Evaluation Metrics
-(1) Encode Time and Decode Time
-For the Encode and Decode stages, the speed at which the device completes the entire processing operations of the respective stages reflects the processing efficiency of the corresponding processes.
-Considering that it is difficult to quantify the amount of processing operations performed by the device into specific numerical values, this paper uses the time taken to complete the same conversion task to reflect the impact of the intermediate representation on processing efficiency during the Encode and Decode stages.
-Specifically, for the same conversion task, the longer the time taken, the lower the processing efficiency; the shorter the time taken, the higher the processing efficiency.
-(2) Post-encoding Size
-Unlike the Encode and Decode stages, where time is used as a metric, this paper employs the data size of the quantum program after encoding (Post-encoding Size) to measure the impact of the intermediate representation on transmission efficiency during the quantum program transmission stage.
-The quantum program transmission stage constitutes a classical communication process, where latency is a crucial performance evaluation metric.
-However, in practical communication environments, latency is influenced by various factors, including the bandwidth and distance of physical communication links and devices, as well as the classical network protocols used for data transmission.
-Quantum program transmission primarily involves transmitting the intermediate representation of the quantum program as pure data.
-The data size of the intermediate representation is the only component of the quantum program intermediate representation that is directly related to communication latency, excluding the influences of physical communication links, devices, and classical network protocols.
-Specifically, for the same quantum program transmission task, a larger Post-encoding Size results in lower transmission efficiency, while a smaller Post-encoding Size leads to higher transmission efficiency.
-VII-A2 Experimental Setup
-This paper conducts comparative experiments from three aspects: the number of circuits, circuit depth, and the number of qubits.
-Batch transmission of a large number of circuits is one of the real and essential practical application requirements.
-Circuit depth is a crucial factor that affects the execution results of quantum circuits.
-The number of qubits measures the amount of quantum computing resources occupied by quantum circuits.
-These three metrics serve as important references for users when submitting quantum programs and for devices when processing quantum programs for execution.
-The experiments in this paper are carried out using the scheme shown in Table I.
-For the experiment on the number of quantum circuits, the number of quantum circuits is taken as the independent variable, while the circuit depth is controlled at a fixed value of 500 and the number of qubits is controlled at a fixed value of 72.
-For the experiment on circuit depth, the circuit depth is taken as the independent variable, with the number of circuits controlled at a fixed value of 500 and the number of qubits controlled at a fixed value of 72.
-For the experiment on the number of qubits, the number of qubits is taken as the independent variable, with the circuit depth controlled at a fixed value of 500 and the number of circuits controlled at a fixed value of 500.
-It should be noted that the parameters related to circuits are not limited to the number of circuits, circuit depth, and the number of qubits.
-Moreover, there may be numerical constraints among these parameters.
-For example, when the number of quantum gates is fixed, there is a certain negative correlation between circuit depth and the number of qubits.
-To minimize the influence of other parameters and the relationships among them, this paper adopts a randomized technical approach, generating random circuits based on the conventions shown in Table I to test the performance of quantum program intermediate representations in quantum program transmission.
-Relevant experiments are conducted on OriginBIS, OriginIR, and QASM to compare the quantum program transmission performance of these intermediate representations.
-TABLE I: Experimental Setup For Quantum Program Transmission
-Experiment Circuit Count Circuit Depth Qubit Count
-For Circuit Count Independent Variable 500 72
-For Circuit Depth 500 Independent Variable 72
-For Qubit Count 500 500 Independent Variable
-VII-A3 Experimental Results and Analysis
-(1) For Circuit Count
-The experimental results for the number of circuits are shown in Figure 14.
-The grouped columns chart in subplots (a), (b), and (c) demonstrates that OriginBIS is significantly smaller than OriginIR and QASM in terms of Encode Time, Decode Time, and Post-encoding Size.
-This indicates that in these experiments, OriginBIS outperforms OriginIR and QASM in the efficiency of quantum program transmission.
-Through simple numerical estimation, it can be found that for Encode Time, OriginIR is approximately 17 times that of OriginBIS, and QASM is approximately 80 times that of OriginBIS; for Decode Time, OriginIR is approximately 20 times that of OriginBIS, and QASM is approximately 40 times that of OriginBIS; for Post-encoding Size, OriginIR is approximately 3.5 times that of OriginBIS, and QASM is approximately 6.5 times that of OriginBIS.
-The point-line charts in subplots (d), (e), and (f) show that for OriginBIS, Encode Time, Decode Time, and Post-encoding Size increase with the growth in the number of circuits.
-(2) For Circuit Depth
-The experimental results regarding circuit depth are presented in Figure 15.
-The grouped columns chart in subplots (a), (b), and (c) demonstrate that OriginBIS exhibits significantly lower values in terms of Encode Time, Decode Time, and Post-encoding Size compared to both OriginIR and QASM.
-This indicates that in these experiments, OriginBIS outperforms OriginIR and QASM in terms of efficiency for quantum program transmission.
-Upon simple numerical estimation, it can be observed that for Encode Time, OriginIR is approximately 18 times that of OriginBIS, while QASM is about 88 times; for Decode Time, OriginIR is roughly 21 times that of OriginBIS, and QASM is approximately 39 times; for Post-encoding Size, OriginIR is around 3.5 times that of OriginBIS, and QASM is about 6.5 times.
-The point-line charts in subplots (d), (e), and (f) reveal that for OriginBIS, Encode Time, Decode Time, and Post-encoding Size increase with the growth of circuit depth.
-(a) IRs’ Encode Time
-(b) IRs’ Decode Time
-(c) IRs’ Post-encoding Size
-(d) OriginBIS’ Encode Time
-(e) OriginBIS’ Decode Time
-(f) OriginBIS’ Post-encoding Size
-Figure 17: For Gate Count
-(3) For Qubit Count
-The experimental results concerning the number of qubits are illustrated in Figure 16.
-The grouped columns chart in subplots (a), (b), and (c) indicates that OriginBIS exhibits significantly lower Encode Time and Decode Time compared to both OriginIR and QASM.
-Additionally, for qubit counts of 50 or more, OriginBIS also demonstrates lower Post-encoding Time compared to OriginIR and QASM.
-This suggests that in the majority of the samples tested, OriginBIS outperforms OriginIR and QASM in terms of efficiency for quantum program transmission.
-Excluding the 90-qubit data in subplot (b) and the 40-bit data in subplot (c), upon simple numerical estimation, it can be observed that for Encode Time, OriginIR is approximately 21 times that of OriginBIS, while QASM is about 93 times; for Decode Time, OriginIR is roughly 20 times that of OriginBIS, and QASM is approximately 39 times; for Post-encoding Size, OriginIR is around 3.5 times that of OriginBIS, and QASM is about 6.5 times.
-The point-line charts in subplots (d), (e), and (f) reveal that for OriginBIS, there is a trend of increasing Encode Time and Decode Time with the growth in the number of qubits, albeit with considerable data fluctuation.
-However, Post-encoding Time consistently increases with the increase in the number of qubits.
-(4) Gate Count
-During the experiments examining circuit count, circuit depth, and qubit count, it was observed that the number of quantum gates varies with changes in these parameters.
-Figure 18 illustrates the trend in the number of quantum gates, demonstrating a strong positive linear correlation with the value of the third parameter when the values of any two of these three quantities are held constant.
-The number of gates is also an important metric for evaluating quantum programs.
-By combining the gate count data from Figure 18 with the corresponding Encode Time, Decode Time, and Post-encoding Size data, the results presented in Figure 17 were obtained.
-Subplots (a), (b), and (c) indicate that the Encode Time, Decode Time, and Post-encoding Size for OriginBIS are significantly lower than those for OriginIR and QASM.
-Subplots (d), (e), and (f) reveal that there is a trend of increasing Encode Time and Decode Time with the growth in the number of quantum gates, albeit with noticeable data fluctuations.
-However, Post-encoding Time increases with the increase in the number of quantum gates, exhibiting a strong linear characteristic.
-The experimental results concerning the number of qubits are illustrated in Figure 16.
-The grouped columns chart in subplots (a), (b), and (c) indicates that OriginBIS exhibits significantly lower Encode Time and Decode Time compared to both OriginIR and QASM.
-Additionally, for qubit counts of 50 or more, OriginBIS also demonstrates lower Post-encoding Time compared to OriginIR and QASM.
-This suggests that in the majority of the samples tested, OriginBIS outperforms OriginIR and QASM in terms of efficiency for quantum program transmission.
-Excluding the 90-qubit data in subplot (b) and the 40-qubit data in subplot (c), upon simple numerical estimation, it can be observed that for Encode Time, OriginIR is approximately 21 times that of OriginBIS, while QASM is about 93 times; for Decode Time, OriginIR is roughly 20 times that of OriginBIS, and QASM is approximately 39 times; for Post-encoding Size, OriginIR is around 3.5 times that of OriginBIS, and QASM is about 6.5 times.
-The point-line charts in subplots (d), (e), and (f) reveal that for OriginBIS, there is a trend of increasing Encode Time and Decode Time with the growth in the number of qubits, albeit with considerable data fluctuation.
-However, Post-encoding Time consistently increases with the increase in the number of qubits.
-Figure 18: Gate Count for Circuit Count, Circuit Depth and QBit Count
-Figure 19: Experimental Results About Circuit Compilation
-VII-B Compilation Efficiency Comparison Experiment
-VII-B1 Experiment Setup
-In this paper, a comparative experiment related to compilation efficiency was conducted on multiple mainstream quantum programming frameworks based on Benchpress.
-Benchpress is an evaluation suite for multi-quantum computing software development kits, containing thousands of test cases.
-It allows for the uniform testing of multiple quantum software packages across various performance and functional indicators, with the results reflecting the cost of processing quantum circuits on quantum computing devices using different software packages.
-All relevant experiments for each software package mentioned in the Benchpress literature were replicated in this study, with the addition of tests for QPanda3.
-All software packages tested are listed in Table II.
-It should be noted that a timeout limit of 180 seconds was imposed for executing each test case.
-The testing of these software packages was conducted in the same software and hardware environments. The environmental parameters used in our experiments are presented in Table III.
-TABLE II: SDKs
-SDK Version
-amazon-braket-sdk(braket) 1.91.0
-bqskit 1.2.0
-cirq 1.4.1
-pytket(tket) 2.0.1
-qiskit 1.4.2
-qiskit_ibm_transpiler 0.10.3
-pystaq 3.5
-pyqpanda3 0.2.0
-TABLE III: Software and Hardware Environments
-Type Info
-CPU Intel(R) Xeon(R) Platinum 8336C CPU @ 2.30GHz
-Memory 2.0TB
-Operating System Ubuntu 20.04.5 LTS
-Python 3.11.11
-VII-B2 Evaluation Metrics
-This paper adopts the evaluation metrics introduced by Benchpress. They are Standard Pytest Output Type, circuit construction time, manipulate time, and circuit compilation time.
-VII-B3 Experimental Results And Analysis
-TABLE IV: Experimental Results Evaluated Based on The Criterion of Standard Pytest Output Type
-SDK Passed Failed XFailed Skipped Total
-braket 7 2 0 1057 1066
-cirq 10 2 0 1054 1066
-qiskit 1044 0 0 22 1066
-QPanda3 1016 28 0 22 1066
-In our experiments, the results of the test_status_counts are presented in Table IV.
-It was observed that many SDKs failed to satisfy the requirements of numerous test cases in Benchpress.
-Specifically, bqskit was only capable of performing tests related to circuit construction.
-pytket and pystaq exhibited compatibility issues with the environments of other SDKs, making it impossible to conduct fair testing and comparison using a unified test suite and environment.
-Furthermore, qiskit_ibm_transpiler demonstrated strong dependency on the IBM computing platform, rendering it untestable under our experimental environment.
-As evident from Table IV, both braket and cirq skipped a significant number of test cases.
-According to the Benchpress literature, qiskit skipped 22 test cases and passed 1,044 test cases.
-Tabel IV indicates that our experiments actually tested multiple SDKs using 1,066 test cases.
-The performance of qiskit in our experiments was consistent with the reports in the Benchpress literature.
-QPanda successfully passed $95.3\%$ of all test cases, with the number of skipped cases equal to that of qiskit.
-This table showcases the excellent versatility of QPanda3 within the unified test suite, Benchpress.
-The performance of each SDK in circuit construction is illustrated in Figure 20.
-It can be observed that QPanda3, qiskit, and cirq passed all seven test cases.
-For the five test cases, namely test_multi_control_circuit, test_bigint_qasm2_import, test_param_circSU2_100_build, test_clifford_build, and test_QV100_build, QPanda3 demonstrated the shortest execution time.
-For the test case test_param_circSU2_100_bind, QPanda3’s execution time was almost on par with qiskit, ranking as the second fastest.
-In the test case test_QV100_qasm2_import, although QPanda3 lagged behind qiskit, its execution time was significantly less than that of cirq and bqskit.
-Despite QASM not being the native intermediate representation for quantum programs in QPanda3, it still exhibited a circuit construction speed close to that of qiskit in these two test cases.
-This suggests that the QASM conversion tool in QPanda3 may be highly efficient, or that QPanda3’s circuit construction efficiency is so high that it can offset any inefficiencies in QASM conversion.
-This figure highlights the advantages of QPanda3 in circuit construction.
-Figure 20: Experimental Results About Circuit Construction(shorter is better)
-Figure 21: Experimental Results About Circuit Manipulation(shorter is better)
-The Benchpress literature highlights that Manipulate is a crucial metric for evaluating SDKs.
-As shown in Figure 21, only QPanda3 and qiskit passed all four test cases.
-QPanda3 demonstrated the shortest execution time in three tests: test_multi_control_decompose, test_QV100_basis_change, and test_random_clifford_decompose.
-For the test_DTC100_twirling, QPanda3’s execution time was only slightly longer than that of qiskit, with a minimal difference between the two.
-This figure indicates that QPanda3 has a clear advantage in the Manipulate metric.
-In our experiments, only QPanda3 and qiskit passed the majority of the test cases related to compilation.
-We have plotted the corresponding experimental results into multiple scatter plots, as shown in Figure 19, which illustrate the differences between QPanda3 and qiskit in terms of 2Q Gate Count, 2Q Gate Depth, and compilation time for various topological structures.
-In each plot, the blue dashed line represents the point where the values for QPanda3 and qiskit are equal for the corresponding metric.
-Points above the blue dashed line indicate that the corresponding value for QPanda3 is less than that for qiskit.
-As observed from Figure 19, for both 2Q Gate Count and 2Q Gate Depth, the majority of the scatter points are densely distributed along or very close to the blue dashed line, suggesting that the performance of QPanda3 and qiskit is essentially the same.
-However, regarding compilation time, all scatter points in the four subplots of Figure 19 are located above the blue dashed line, indicating that QPanda3 exhibits higher compilation efficiency than qiskit.
-This characteristic is not limited by topological structure and is effective for a large number of different circuits.
-Additionally, it can be observed that only a very few test cases result in compilation times exceeding 10 seconds when using QPanda3.
-These observations collectively demonstrate the high compilation efficiency of QPanda3.
-VIII Conclusion
-This paper focuses on efficient compilation and execution in quantum computing, introducing OriginIR and OriginBIS as intermediate representations to enhance programmability, transmission efficiency, and execution performance.
-Experimental results demonstrate that OriginBIS significantly outperforms OpenQASM 2.0 in encoding and decoding speed as well as information capacity.
-Additionally, QPanda3 surpasses Qiskit in quantum circuit construction, operation execution, and compilation speed, exhibiting exceptional performance improvements, particularly in large-scale quantum circuit compilation tasks.
-These advancements lay a solid foundation for the engineering applications of future quantum computing.
-IX Future Work
-The quantum programming framework QPanda3 is poised to enhance its capabilities for future ”quantum-HPC-AI” integrated computing by focusing on several key advancements.
-Building on its existing strength of compiling quantum circuits 14.97 times faster than Qiskit in Benchpress tests, QPanda3 will develop hybrid task schedulers to enable seamless coordination between quantum computations and classical supercomputing resources.
-It aims to integrate AI-driven optimization tools, such as AutoML modules, to automatically reduce quantum gate counts and circuit depths while mitigating noise impacts.
-The framework will expand its hybrid programming interfaces to support classical HPC technologies like MPI and CUDA, facilitating efficient quantum-classical algorithm interoperability.
-Additionally, QPanda3 plans to implement a unified resource management platform for dynamic allocation of quantum hardware, supercomputing clusters, and AI accelerators.
-To strengthen ecosystem integration, it will promote standardization through open-source collaboration, enabling deeper compatibility with classical AI libraries (e.g., TensorFlow) while maintaining its low-latency compilation advantages.
-These upgrades will position QPanda3 as a core enabler for cross-paradigm applications in finance, drug discovery, and materials science.
+Appendix D Variational Quantum Circuits
+[…]
+1from pyqpanda3.vqcircuit import VQCircuit
+Listing 7: An example Python program that update and modify quantum states.
+D-A Constructing Variational Quantum Circuits
+D-A1 Setting Formal Parameters $\theta$
+Our recommendation is that formal parameters (or placeholders) with the same name or symbol (e.g., $\theta$) should refer to the same parameter object (or variable) within the same context, while in different contexts they should refer to distinct parameter objects (or variables).
+It is possible for two different variational quantum circuits to both use $\theta$ as the name for their formal parameters, which may introduce certain issues.
+On one hand, naming conflicts commonly occur across various domains.
+During programming, such conflicts may lead to logical confusion in the code and consequently cause computational errors.
+On the other hand, larger-scale software programs may involve deeply nested parameter passing, making maintenance challenging due to difficulties in determining parameter meanings.
+These issues are particularly pronounced in weakly typed programming languages like Python.
+QPanda3 binds each variational quantum circuit to a formal parameter (corresponding to $\theta$ in $U(\theta)$), and this formal parameter is only valid for that specific variational quantum circuit.
+Through this enforced constraint, QPanda3 alleviates the aforementioned problems to some extent.
+Additionally, since $U(\theta)$ is a unary function of $\theta$, the process of explicitly naming the formal parameter (which effectively corresponds to $\theta$) can be omitted during programming.
+This not only avoids ambiguity but also simplifies the programming workflow.
+[…]
+(1) Scalar $\theta\in\mathbb{R}$
+1from pyqpanda3.vqcircuit import VQCircuit
+2vqc = VQCircuit()
+3vqc.set_Param([1])
+Listing 8: An example Python program that update and modify quantum states.
+[…]
+1from pyqpanda3.vqcircuit import VQCircuit
+2vqc = VQCircuit()
+3vqc.set_Param([3])
+Listing 9: An example Python program that update and modify quantum states.
+[…]
+1from pyqpanda3.vqcircuit import VQCircuit
+2vqc = VQCircuit()
+3vqc.set_Param([3,2])
+Listing 10: An example Python program that update and modify quantum states.
+[…]
+1from pyqpanda3.vqcircuit import VQCircuit
+2vqc = VQCircuit()
+3vqc.set_Param([3,2,4])
+Listing 11: An example Python program that update and modify quantum states.
+D-A2 Adding Quantum Logic Gates
+In QPanda3, general quantum programs are abstracted using the QProg class, while quantum circuits without measurements are abstracted using QCircuit.
+Quantum programs or measurement-free quantum circuits are constructed by adding quantum logic gates to QProg or QCircuit objects.
+For QProg, QCircuit, and VQCircuit, QPanda3 employs a unified approach to adding quantum gates.
+To facilitate differentiation, we refer to gate parameters in constant value form as fixed parameters, and gate parameters whose values are determined by $\theta$ as variable parameters.
+Quantum gates without variable parameters are called non-parameterized quantum gates, while those containing variable parameters are called parameterized quantum gates.
+(1) Adding a single non-parametric quantum gate or quantum gate containing only fixed parameters
+1from pyqpanda3.vqcircuit import VQCircuit
+2from pyqpanda3.core import X,RX
+3vqc = VQCircuit()
+4vqc.set_Param([3])
+5vqc << X(0)
+6vqc << RX(0,3.14)
+Listing 12: An example Python program that update and modify quantum states.
+(2) Adding a measurement-free subcircuit
+1from pyqpanda3.vqcircuit import VQCircuit
+2from pyqpanda3.core import QCircuit,X,RX
+3vqc = VQCircuit()
+4vqc.set_Param([3])
+5cir = QCircuit()
+6cir << X(0)
+7cir << RX(0,3.14)
+8vqc << cir
+Listing 13: An example Python program that update and modify quantum states.
+(3) Adding parameterized quantum gates
+The following code first defines the formal parameters $\theta=(\theta_{0},\theta_{1},\theta_{2})$, then adds a parameterized RX gate to the variational quantum circuit vqc.
+This RX gate operates on the qubit with index 0, and its parameter corresponds to $\theta_{0}$.
+1from pyqpanda3.vqcircuit import VQCircuit
+2from pyqpanda3.core import X,RX,RY
+3vqc = VQCircuit()
+4vqc.set_Param([3])
+5vqc << RX(0,vqc.Param([0]))
+Listing 14: An example Python program that update and modify quantum states.
+QPanda3 allows assigning names to specific elements of the formal parameter $\theta=(\theta_{0},\theta_{1},\theta_{2})$.
+If an element of $\theta$ is assigned a name, that name can subsequently be used to reference the element directly.
+Note that each element can only be named once.
+The following code first defines the formal parameters $\theta=(\theta_{0},\theta_{1},\theta_{2})$, then adds a parameterized RX gate to the variational quantum circuit vqc.
+This RX gate operates on the qubit with index 0, and its parameter corresponds to $\theta_{0}$.
+During the addition of the RX gate, $\theta_{0}$ is assigned the name ”gamma”.
+Finally, a parameterized RY gate is added to the variational quantum circuit vqc.
+This RY gate operates on the qubit with index 1, and its parameter corresponds to the element of $\theta$ named ’gamma’, i.e., $\theta_{0}$.
+1from pyqpanda3.vqcircuit import VQCircuit
+2from pyqpanda3.core import X,RX,RY
+3vqc = VQCircuit()
+4vqc.set_Param([3])
+5vqc << RX(0,vqc.Param([0],’gama’))
+6vqc << RY(1,vqc.Param(’gama’))
+Listing 15: An example Python program that update and modify quantum states.
+D-A3 Reusing Variational Quantum Circuit Structures
+[…]
+A concrete example is shown in the code.
+Since this corresponds to a single-qubit system, the distinction and explanation of qubit indices can be omitted here.
+For clarity, we denote the variational quantum circuit obtained by get_vqc1 as $U^{{1}}(\theta^{(1)})=RX(\theta_{1}^{(1)})\cdot X,\quad\theta^{(1)}=(\theta^{(1)}_{0},\theta^{(1)}_{1})$, and the variational quantum circuit obtained by get_vqc2 as $U^{{2}}(\theta^{(2)})=RZ(\theta_{2}^{{(2)}})\cdot U^{{1}}([\theta_{0}^{{2}},\theta_{1}^{{2}}]),\quad\theta^{(2)}=(\theta^{(2)}_{0},\theta^{(2)}_{1},\theta^{(2)}_{2})$.
+Thus, $U^{{2}}(\theta^{(2)})=RZ(\theta_{2}^{{(2)}})\cdot RX(\theta_{1}^{(2)})\cdot X$.
+Although this code below is relatively simple, it demonstrates how reusing variational quantum circuit structures can better achieve modularity in program code, improve code reusability, and facilitate the construction and maintenance of larger-scale program systems.
+1from pyqpanda3.vqcircuit import VQCircuit
+2from pyqpanda3.core import QCircuit,X,RX,RZ
+3
+4def get_vqc1():
+5 vqc = VQCircuit()
+6 vqc.set_Param([2])
+7 vqc << X(0)
+8 vqc << RX(0,vqc.Param([1]))
+9 return vqc
+10def get_vqc2():
+11 vqc2 = VQCircuit()
+12 vqc2.set_Param([3])
+13 vqc1= get_vqc1()
+14 vqc2.append(
+15 vqc1
+16 ,[
+17 (vqc1.Param([0]),vqc2.Param([0]))
+18 ,(vqc1.Param([1]),vqc2.Param([1]))
+19 ]
+20 )
+21 vqc2 << RZ(2,vqc2.Param([2]))
+22 return vqc2
+23
+24vqc = get_vqc2()
+Listing 16: An example Python program that update and modify quantum states.
+D-A4 Expressions as Parameters for Parameterized Quantum Gates
+QPanda3 supports the construction of variational quantum circuits in the form of formula U_F.
+Here, the expression refers to one derived from the formal parameter $\theta$ through mathematical operations.
+Although this formula does not explicitly mention subcircuits without variable parameters, it can clearly be easily extended to variational quantum circuits containing subcircuits without variable parameters.
+It should be noted that this case must also satisfy the requirement of deep binding between formal parameters and the variational quantum circuit.
+Therefore, an expression should not be composed of formal parameters (or their elements) from multiple variational quantum circuits.
+Otherwise, when such an expression serves as a gate parameter in a variational quantum circuit, it would violate this constraint.
+Currently, for expressions serving as parameters of parameterized quantum gates, QPanda3 supports addition, multiplication, and scalar multiplication operations, providing functionality for expression evaluation and expression printing.
+The following code demonstrates the basic usage.
+For the formal parameter $\theta=(p_{0},p_{1})$ of the variational quantum circuit in the code, the expression is $e=3.14p_{0}p_{1}+p_{1}+4$.
+1from pyqpanda3.vqcircuit import VQCircuit
+2from pyqpanda3.core import RX
+3vqc = VQCircuit()
+4vqc.set_Param([2])
+5p0 = vqc.Param([0],’p0’)
+6p1 = vqc.Param([1],’p1’)
+7e = 3.14*p0*p1+p1+4
+[…]
+9vqc << RX(0,e)
+Listing 17: An example Python program that update and modify quantum states.
+The corresponding output is:
+1e:
+2 ((3.14*p0*p1+p1)+4)
+Listing 18: An example Python program that update and modify quantum states.
+D-A5 Displaying the Structure of Variational Quantum Circuits
+Currently, QPanda3 provides functionality to display the structure of variational quantum circuits in a format similar to OriginIR [3], which is implemented through the member method display_ansatz of VQCircuit.
+The output of this method, as observed from the code results, excludes the qubit and classical bit declarations required by OriginIR but otherwise largely aligns with OriginIR’s format except for variable quantum gate parameters.
+For parameter positioning, variable parameters follow the same convention as fixed parameters.
+In terms of output format, fixed parameters are displayed as numerical values, while variable parameters are represented by identifier strings.
+These identifier strings for variable parameters use assigned names if available for specific elements of the formal parameter; otherwise, they adopt a fixed format ”Param(list of indices of the element in $\theta$)” to represent the element.
+1from pyqpanda3.vqcircuit import VQCircuit
+2from pyqpanda3.core import RX,RZ
+3vqc = VQCircuit()
+4vqc.set_Param([2,2])
+5p0 = vqc.Param([0,0],’p0’)
+6p1 = vqc.Param([0,1],’p1’)
+7e = (3.14*p0*p1+p1+4)*vqc.Param([1,1])
+8vqc << RX(0,e)
+9vqc << RZ(0,vqc.Param([1,0]))
+[…]
+11vqc.display_ansatz()
+Listing 19: An example Python program that update and modify quantum states.
+The corresponding output is:
+1vqc:
+2
+3RX q[0],(((3.14*p0*p1+p1)+4)*Param([1,1]))
+4RZ q[0],(Param([1,0]))
+Listing 20: An example Python program that update and modify quantum states.
+D-B Obtaining Quantum States, Expectations, and Gradients
+D-B1 Acquiring Parameterized Quantum States for Given Parameters
+[…]
+In QPanda3, a concrete quantum circuit represented by QCircuit can be obtained by assigning actual values to the formal parameter $\theta$, and the parameterized quantum state with given parameter values can be acquired by evolving the resulting quantum circuit.
+An example usage is provided in the code (Appendix E).
+D-B2 Computing Hamiltonian Expectation Value
+[…]
+QPanda3 provides the functionality to construct Hamiltonian operators using linear combinations of Pauli operations.
+1from pyqpanda3.hamiltonian import Hamiltonian # Import the corresponding package name
+2# Represent a linear combination of Pauli operations as a list of tuples
+3# The first tuple indicates that the coefficient of this term is 12.36525580995888 + 14.85172018664403j,
+4# and the operation applies Pauli Y on the qubit with index 0 and Pauli X on the qubit with index 1
+5paulis = [
+6 (’YX’, [0, 1], (12.36525580995888 + 14.85172018664403j)), # (Pauli basis, qubit indices on which the Pauli basis acts, coefficient)
+7 (’YY’, [0, 1], (12.920260765526914 + 26.29613065236354j)) # The qubit indices used should be consistent with those used in the quantum circuit
+8 ]
+9ham = Hamiltonian(paulis) # ham is the constructed Hamiltonian operator
+Listing 21: An example Python program that update and modify quantum states.
+In QPanda3, two methods are available for computing Hamiltonian expectations, corresponding to get_hamiltonian_expectation and get_hamiltonian_expectation2 in the code (Appendix F).
+D-B3 Computing Gradient Values
+In the field of quantum computing, the concept of gradient extends from the classical rate of function change to the sensitivity of quantum states to parameterized operations [64], with its general form defined as the partial derivative of the measurement expectation value $\left\langle H_{\theta}\right\rangle$ with respect to circuit parameter $\theta$, $\nabla\langle\psi(\theta)|H|\psi(\theta)\rangle$[65].
+This definition incorporates the unitary constraints of quantum evolution and the characteristics of non-commuting operators, introducing statistical uncertainty from quantum measurements compared to classical gradients[66], serving as the optimization engine for variational quantum algorithms.
+The significance of variational quantum gradient computation manifests in multiple dimensions: it directly determines algorithmic feasibility, for example, reducing convergence iterations by 60% when overcoming barren plateaus[67]; in quantum machine learning, it supports weight updates for tasks like image classification, improving accuracy by 23%[68]; simultaneously, it advances the practicality of NISQ devices, expanding solvable problem scales by 10 times[63].
+Quantum gradient computation methods have evolved into four mainstream techniques: finite difference estimates gradients through minor parameter perturbations[69], simple but noise-amplifying and requiring $O(k)$ circuit evaluations; parameter shift rules leverage quantum gate properties for analytical solutions[64][65] [70], more precise but requiring $O(2k)$ executions; automatic differentiation propagates gradients backward through computational graphs in hybrid frameworks[6], $O(1)$ efficient but constrained by quantum no-cloning [71]; adjoint methods[51] enable efficient gradient computation through quantum state operations.
+The adjoint method leverages the time-reversal symmetry in quantum mechanics to compute gradients through the synergy of forward evolution and adjoint evolution: first execute the parameterized circuit $U(\theta)$ to generate the final state, then apply the Hamiltonian $H$ and reversely execute $U^{\dagger}(\theta)$ to revert to the initial state.
+The core advantage of this method lies in its revolutionary improvement in resource efficiency-regardless of the parameter scale, only a constant 2 circuit executions (1 forward + 1 adjoint) are required to obtain the full gradient.
+The set of rotation gates $G=\{RX,RY,RZ,CRX,CRY,CRZ\}$ constitutes the core parameterized gates for variational quantum circuits.
+In variational quantum circuits within QPanda3, these rotation gates can all function as parameterized quantum gates.
+Currently, QPanda3 combines classical backpropagation and the adjoint method to support gradient computation for variational quantum circuits where expressions serve as variable parameters.
+A complete application example is shown in the code (Appendix G).
+D-C Application Examples
+A QAOA application example is provided in Appendix H.
+Appendix E Obtaining Parameterized Quantum States for Given Parameter Values
+1from pyqpanda3.vqcircuit import VQCircuit
+2from pyqpanda3.core import QCircuit,X,RX,RY,Y,CPUQVM,QProg
+3def get_vqc():
+4 # Prepare the variational quantum circuit U(theta)
+5 vqc = VQCircuit()
+6 vqc.set_Param([2]) # 1> Agree that the parameter theta is a one-dimensional vector with a length of 2, 2> The parameter theta is bound to the variational quantum circuit U(theta) and cannot be used for other variational quantum circuits
+7 vqc << RX(0, vqc.Param([0])) # Add a parameterized quantum gate RX, whose parameter corresponds to the first element of the vector theta
+8 vqc << RY(1, vqc.Param([1]))
+9 return vqc
+10def get_param_val():
+11 # Prepare the actual numerical values thetaval for the parameter theta
+12 param_val = [5.14, 6.14] # The number of elements should be consistent with the number of elements in the parameter theta (vector)
+13 return param_val
+14
+15def get_qstate(vqc, param_val):
+16 # Obtain |psi(thetaval))
+17 cir = vqc(param_val).at([0]) # Obtain U(thetaval)
+18 qvm = CPUQVM()
+19 qvm.run(QProg(cir), 1) # Perform the evolution
+20 stv = qvm.result().get_state_vector()
+21 return stv
+22
+23vqc = get_vqc() # Prepare the variational quantum circuit U(theta)
+24param_val = get_param_val() # Prepare the actual numerical values thetaval for the parameter theta
+25stv = get_qstate(vqc, param_val) # Obtain |psi(thetaval))
+26print(’|psi(thetaval)):’,stv)
+Listing 22: An example Python program that update and modify quantum states.
+Appendix F Example Code for Computing Hamiltonian Expectation
+1from pyqpanda3.vqcircuit import VQCircuit,DiffMethod
+2from pyqpanda3.core import QCircuit,X,RX,RY,Y,CPUQVM,QProg
+3from pyqpanda3.hamiltonian import Hamiltonian
+4def get_vqc():
+5 # Prepare the variational quantum circuit U(theta)
+6 vqc = VQCircuit()
+7 vqc.set_Param([2]) # 1> Agree that the parameter theta is a one-dimensional vector with a length of 2.
+2> The parameter theta is bound to the variational quantum circuit U(theta) and cannot be used for other variational quantum circuits.
+8 vqc << RX(0, vqc.Param([0])) # Add a parameterized quantum gate RX, whose parameter corresponds to the first element of the vector theta
+9 vqc << RY(1, vqc.Param([1]))
+10 return vqc
+11
+12def get_param_val():
+13 # Prepare the actual numerical values thetaval for the parameter theta
+14 param_val = [2.14, 3.14] # The number of elements should be consistent with the number of elements in the parameter theta (vector)
+15 return param_val
+16
+17def get_hamiltonian():
+18 # Prepare the Hamiltonian operator H
+19 paulis = [
+20 (’YY’, [0, 1], (12.36525580995888 + 14.85172018664403j)), # (Pauli basis, qubit indices on which the Pauli basis acts, coefficient)
+21 (’YX’, [0, 1], (12.920260765526914 + 26.29613065236354j)) # The qubit indices used should be consistent with those used in the quantum circuit
+22 ]
+23
+24 return Hamiltonian(paulis)
+25
+26def get_hamiltonian_expectation(vqc, param_val, ham):
+27 # Obtain the Hamiltonian expectation value (Method 1)
+28 res = vqc(param_val).expval_hamiltonian(
+29 ham
+30 , [0]
+31 , used_threads=4
+32 , backend=’CPU’
+33 )
+34 return res
+35
+36def get_hamiltonian_expectation2(vqc: VQCircuit, param_val, ham):
+37 # Obtain the Hamiltonian expectation value (Method 2)
+38 # The interface get_gradients_and_expectation can obtain both the gradient values and the expectation values simultaneously. Here, only the expectation value is returned.
+39 res = vqc.get_gradients_and_expectation(
+40 params=param_val
+41 , observable=ham
+42 , diff_method=DiffMethod.ADJOINT_DIFF
+43 ).expectation_val()
+44 return res
+45
+46vqc = get_vqc() # Prepare the variational quantum circuit U(theta)
+47param_val = get_param_val() # Prepare the actual numerical values thetaval for the parameter theta
+48ham = get_hamiltonian() # It is recommended not to use ’H’ as the variable name to avoid naming conflicts with the H gate.
+49ham_expectation = get_hamiltonian_expectation(
+50 vqc
+51 , param_val
+52 , ham
+53 ) # Obtain the Hamiltonian expectation value (Method 1)
+54ham_expectation2 = get_hamiltonian_expectation2(
+55 vqc
+56 , param_val
+57 , ham
+58 ) # Obtain the Hamiltonian expectation value (Method 2)
+59
+60print(’ham_expectation:’,ham_expectation)
+61print(’ham_expectation2:’,ham_expectation2)
+Listing 23: An example Python program that update and modify quantum states.
+Appendix G Example Code for Computing Gradients of Variational Quantum Circuits
+1
+2from pyqpanda3.vqcircuit import VQCircuit,DiffMethod
+3from pyqpanda3.core import QCircuit,X,RX,RY,Y,CPUQVM,QProg
+4from pyqpanda3.hamiltonian import Hamiltonian
+5
+6def get_vqc():
+7 # Prepare the variational quantum circuit U(theta)
+8 vqc = VQCircuit()
+9 vqc.set_Param([2]) # 1> Agree that the parameter theta is a one-dimensional vector with a length of 2.
+2> The parameter theta is bound to the variational quantum circuit U(theta) and cannot be used for other variational quantum circuits.
+10 vqc << RX(0, vqc.Param([0])) # Add a parameterized quantum gate RX, whose parameter corresponds to the first element of the vector theta
+11 vqc << RY(1, vqc.Param([1]))
+12 return vqc
+13
+14def get_param_val():
+15 # Prepare the actual numerical values thetaval for the parameter theta
+16 param_val = [2.14, 3.14] # The number of elements should be consistent with the number of elements in the parameter theta (vector)
+17 return param_val
+18
+19def get_hamiltonian():
+20 # Prepare the Hamiltonian operator H
+21 paulis = [
+22 (’YY’, [0, 1], (12.36525580995888 + 14.85172018664403j)), # (Pauli basis, qubit indices on which the Pauli basis acts, coefficient)
+23 (’YX’, [0, 1], (12.920260765526914 + 26.29613065236354j)) # The qubit indices used should be consistent with those used in the quantum circuit
+24 ]
+25
+26 return Hamiltonian(paulis)
+27
+28def get_gradient(vqc, param_val, ham):
+29 # Obtain the gradient values (Method 1)
+30 return vqc.get_gradients(
+31 params=param_val
+32 , observable=ham
+33 , diff_method=DiffMethod.ADJOINT_DIFF
+34 )
+35
+36def get_gradient2(vqc: VQCircuit, param_val, ham):
+37 # Obtain the gradient values (Method 2)
+38 # The interface get_gradients_and_expectation can obtain both the gradient values and the expectation values simultaneously.
+39 return vqc.get_gradients_and_expectation(
+40 params=param_val
+41 , observable=ham
+42 , diff_method=DiffMethod.ADJOINT_DIFF
+43 )
+44
+45vqc = get_vqc() # Prepare the variational quantum circuit U(theta)
+46param_val = get_param_val() # Prepare the actual numerical values thetaval for the parameter theta
+47ham = get_hamiltonian() # It is recommended not to use ’H’ as the variable name to avoid naming conflicts with the H gate.
+48res1 = get_gradient(vqc, param_val, ham) # Obtain the gradient values (Method 1)
+49res2 = get_gradient2(vqc, param_val, ham) # Obtain the gradient values (Method 2)
+50
+51print(’res1:’,res1)
+52print(’res1 gradient:’,res1.gradients())
+53print(’res2:’,res2)
+54print(’res2 gradient:’,res2.gradients())
+Listing 24: An example Python program that update and modify quantum states.
+Appendix H A QAOA Application Example (Max Cut Problem)
+H-A Problem Definition
+Given an undirected graph $G=(V,E)$, where:
+  • $V$ is the vertex set ($|V|=n$)
+  • $E$ is the edge set, with each edge having weight $w_{ij}$
+[…]
+  • Quantum encoding of this problem
+    • Assign qubit to each vertex $i$, with state $|0\rangle$ or $|1\rangle$ indicating subset membership
+    • Problem Hamiltonian: $H_{C}=\frac{1}{2}\sum_{(i,j)\in E}w_{ij}(1-Z_{i}Z_{j})$
+    • Minimizing $\langle H_{C}\rangle$ is equivalent to maximizing cut value
+H-B A Specific Problem Instance (Max Cut Problem for a Square Graph)
+  • Vertex set V: {0, 1, 2, 3}
+  • Edge set E: see Table V
+    TABLE V: Max Cut Edge Set E
+    Edge Weight Note
+    (0,1) w=1.0
+    (1,2) w=1.0
+    (2,3) w=1.0
+    (3,0) w=1.0
+    (0,2) w=0.5 Diagonal
+  • Theoretical optimal solution: vertex partition {0,2} and {1,3}
+H-C Solution Using VQCircuit
+H-C1 Step List
+  1. 1.
+    Construct the problem Hamiltonian $H_{C}$
+  2. 2.
+    Design the QAOA parameterized quantum circuit
+  3. 3.
+    Optimize parameters using a classical optimizer
+  4. 4.
+    Measure the optimal quantum state to obtain the solution
+H-C2 Example Code Implementation
+1
+2from pyqpanda3.hamiltonian import Hamiltonian
+3from pyqpanda3.vqcircuit import VQCircuit, DiffMethod
+4from pyqpanda3.core import H, RZ, RX, CNOT, QProg, CPUQVM, measure
+5import numpy as np
+6import matplotlib.pyplot as plt
+7import networkx as nx
+8
+9# =====================
+10# 1 Problem definition and Hamiltonian construction
+11# =====================
+12
+13# Create graph structure (vertices 0-3)
+14edges = [
+15 (0, 1, 1.0), # (vertex i, vertex j, weight)
+16 (1, 2, 1.0),
+17 (2, 3, 1.0),
+18 (3, 0, 1.0),
+19 (0, 2, 0.5) # diagonal
+20]
+21
+22# Correctly construct Max Cut Hamiltonian H_C = 1/2 SUM(w_ij (I - Z_i Z_j))
+23ham_terms = []
+24for i, j, w in edges:
+25 term = "ZZ"
+26 # Coefficient should be positive since we need to maximize the cut value
+27 ham_terms.append((term, [i, j], 0.5 * w))
+28
+29maxcut_hamiltonian = Hamiltonian(ham_terms)
+30print("Problem Hamiltonian:", maxcut_hamiltonian)
+31
+32
+33# =====================
+34# 2. Construct QAOA quantum circuit
+35# =====================
+36
+37def build_qaoa_circuit(num_qubits, p=1):
+38 """
+39 Build p-layer QAOA circuit
+40 :param num_qubits: Number of qubits (vertices)
+41 :param p: QAOA layers
+42 :return: VQCircuit object
+43 """
+44 vqc = VQCircuit()
+45 # Set parameters: p layers, 2 parameters per layer (gamma, beta)
+46 vqc.set_Param([p, 2])
+47
+48 # Initial Hadamard gates
+49 for q in range(num_qubits):
+50 vqc << H(q)
+51
+52 # QAOA layers
+53 for layer in range(p):
+[…]
+55 for (i, j, w) in edges:
+56 gamma = vqc.Param([layer, 0]) # gamma parameter
+57 vqc << CNOT(i, j) << RZ(j, 2 * gamma * w) << CNOT(i, j)
+58
+[…]
+[…]
+61 for q in range(num_qubits):
+[…]
+63
+64 return vqc
+65
+66
+67# Build 2-layer QAOA circuit
+68num_vertices = 4
+69qaoa_layers = 2
+70vqc = build_qaoa_circuit(
+71 num_vertices
+72 , p=qaoa_layers
+73 )
+[…]
+75vqc.display_ansatz()
+76
+77
+78# =====================
+79# 3. Parameter optimization
+80# =====================
+81
+82def optimize_parameters(vqc, hamiltonian):
+83 """
+84 Optimize QAOA parameters
+85 :return: Optimized parameter values
+86 """
+87 # Random initial parameters
+88 init_params = np.random.uniform(
+89 0
+90 , np.pi
+91 , size=qaoa_layers * 2
+92 )
+93
+94 # Use BFGS optimizer
+95 from scipy.optimize import minimize
+96
+97 def objective(params):
+98 """Objective function: Minimize Hamiltonian expectation"""
+99 params_2d = params.reshape((qaoa_layers, 2))
+100 result = vqc.get_gradients_and_expectation(
+101 params_2d,
+102 hamiltonian,
+103 diff_method=DiffMethod.ADJOINT_DIFF
+104 )
+105 return result.expectation_val(), np.array(result.gradients()).flatten()
+106
+107 # Optimization process
+108 res = minimize(
+109 objective,
+110 init_params,
+111 jac=True, # Use gradient information
+112 method=’L-BFGS-B’,
+113 bounds=[(0, 2 * np.pi)] * len(init_params),
+114 options={’maxiter’: 50}
+115 )
+116
+117 if res.success:
+[…]
+119 return res.x.reshape((qaoa_layers, 2))
+120 else:
+121 raise RuntimeError("Parameter optimization failed: " + res.message)
+122
+123
+124# Execute parameter optimization
+125optimal_params = optimize_parameters(vqc, maxcut_hamiltonian)
+126print("Optimal parameter values:")
+127print(optimal_params)
+128
+129
+130# =====================
+131# 4. Result measurement and analysis
+132# =====================
+133
+134def measure_solution(vqc, params):
+135 """Measure optimal quantum state to obtain solution"""
+136 # Get optimized quantum circuit
+137 prog = QProg()
+138 prog << vqc(params).at([0])
+139
+140 # Quantum virtual machine execution
+141 qvm = CPUQVM()
+142
+143 # Add measurement gates
+144 for i in range(num_vertices):
+145 prog << measure(i, i)
+146
+147 # Sample 1000 times
+148 qvm.run(prog, shots=5000)
+149 results = qvm.result().get_prob_dict()
+150
+151 # Parse results
+152 solutions = []
+153 for bitstr, prob in results.items():
+154 bitstr_full = bitstr.zfill(num_vertices)
+155 # Reverse bit string to match vertex order
+156 bitstr_full = bitstr_full[::-1]
+157
+158 cut_value = 0
+159 # Calculate cut value
+160 for i, j, w in edges:
+161 if bitstr_full[i] != bitstr_full[j]:
+162 cut_value += w
+163 solutions.append((bitstr_full, cut_value, prob))
+164
+165 # Sort by cut value
+166 solutions.sort(key=lambda x: x[1], reverse=True)
+167 return solutions
+168
+169
+170# Get and print results
+171solutions = measure_solution(
+172 vqc
+173 , optimal_params
+174 )
+[…]
+176print("Bitstring | Cut Value | Probability")
+177print("-----------------------------------")
+178for bitstr, cut_val, prob in solutions[:5]:
+179 print(f"{bitstr} | {cut_val:.1f} | {prob:.4f}")
+180
+181# =====================
+182# 5. Visualization of results
+183# =====================
+184
+185# Extract cut value distribution
+186tmp_dict = {}
+187for _,cut_val,prob in solutions:
+188 if cut_val in tmp_dict.keys():
+189 tmp_dict[cut_val]+=prob
+190 else:
+191 tmp_dict[cut_val]=prob
+192cut_values = []
+193probabilities = []
+194
+195for val,prob in tmp_dict.items():
+196 cut_values.append(val)
+197 probabilities.append(prob)
+198
+199plt.figure(figsize=(10, 6))
+200plt.bar(cut_values, probabilities, width=0.1)
+201plt.xlabel(’Cut Value’)
+202plt.ylabel(’Probability’)
+203plt.title(’Max Cut Solution Distribution’)
+204plt.xticks(sorted(set(cut_values)))
+205plt.grid(alpha=0.2)
+206plt.savefig(’cut_distribution.png’) # Save image
+207plt.show()
+208
+209# Visualize optimal solution
+210optimal_solution = solutions[0][0]
+211# Assign colors using reversed bit string
+212node_colors = [’red’ if bit == ’0’ else ’blue’ for bit in optimal_solution]
+213
+214# Draw graph structure
+215G = nx.Graph()
+216G.add_edges_from([(i, j, {’weight’: w}) for i, j, w in edges])
+217
+218pos = nx.spring_layout(G)
+219nx.draw(
+220 G
+221 , pos
+222 , node_color=node_colors
+223 , with_labels=True
+224 ,node_size=800
+225 , font_size=16
+226 , edge_color=’gray’
+227 )
+228edge_labels = {(i, j): f"w={w}" for i, j, w in edges}
+229nx.draw_networkx_edge_labels(
+230 G
+231 , pos
+232 , edge_labels=edge_labels
+233 )
+234plt.title(f’Optimal Cut: {solutions[0][1]}’)
+235plt.savefig(’optimal_cut.png’) # Save image
+236plt.show()
+Listing 25: An example Python program that evaluates the properties of quantum states.
+H-D Explanation of Running Results
+H-D1 One Run Result with 700 shots
+1Problem Hamiltonian: { qbit_total = 4, pauli_with_coef_s = { ’Z0 Z1 :0.5 + 0j, ’Z1 Z2 :0.5 + 0j, ’Z2 Z3 :0.5 + 0j, ’Z0 Z3 :0.5 + 0j, ’Z0 Z2 :0.25 + 0j, } }
+2
+3QAOA circuit structure:
+4
+5
+6H q[0]
+7H q[1]
+8H q[2]
+9H q[3]
+10CNOT q[0],q[1]
+11RZ q[1],(1*2*Param([0,0]))
+12CNOT q[0],q[1]
+13CNOT q[1],q[2]
+14RZ q[2],(1*2*Param([0,0]))
+15CNOT q[1],q[2]
+16CNOT q[2],q[3]
+17RZ q[3],(1*2*Param([0,0]))
+18CNOT q[2],q[3]
+19CNOT q[3],q[0]
+20RZ q[0],(1*2*Param([0,0]))
+21CNOT q[3],q[0]
+22CNOT q[0],q[2]
+23RZ q[2],(0.5*2*Param([0,0]))
+24CNOT q[0],q[2]
+25RX q[0],(2*beta_0)
+26RX q[1],(2*beta_0)
+27RX q[2],(2*beta_0)
+28RX q[3],(2*beta_0)
+29CNOT q[0],q[1]
+30RZ q[1],(1*2*Param([1,0]))
+31CNOT q[0],q[1]
+32CNOT q[1],q[2]
+33RZ q[2],(1*2*Param([1,0]))
+34CNOT q[1],q[2]
+35CNOT q[2],q[3]
+36RZ q[3],(1*2*Param([1,0]))
+37CNOT q[2],q[3]
+38CNOT q[3],q[0]
+39RZ q[0],(1*2*Param([1,0]))
+40CNOT q[3],q[0]
+41CNOT q[0],q[2]
+42RZ q[2],(0.5*2*Param([1,0]))
+43CNOT q[0],q[2]
+44RX q[0],(2*beta_1)
+45RX q[1],(2*beta_1)
+46RX q[2],(2*beta_1)
+47RX q[3],(2*beta_1)
+48
+49Optimization successful! Expectation value: -0.3581
+50Optimal parameter values:
+51[[3.68782485 2.28451761]
+52 [0.68693742 1.32432699]]
+53
+54Top 5 solutions:
+55Bitstring&Cut Value&Probability
+56-----------------------------------
+571010 &4.0 &0.1400
+580101 &4.0 &0.1443
+591000 &2.5 &0.0986
+601100 &2.5 &0.0100
+610010 &2.5 &0.0814
+Listing 26: An example Python program that update and modify quantum states.
+Figure 30: Optimal Cut Result for 700 Shots
+Figure 31: Optimal Cut Distribution for 700 Shots
+H-D2 One Run Result with 5000 shots
+1Problem Hamiltonian: { qbit_total = 4, pauli_with_coef_s = { ’Z0 Z1 :0.5 + 0j, ’Z1 Z2 :0.5 + 0j, ’Z2 Z3 :0.5 + 0j, ’Z0 Z3 :0.5 + 0j, ’Z0 Z2 :0.25 + 0j, } }
+2
+3QAOA circuit structure:
+4H q[0]
+5H q[1]
+6H q[2]
+7H q[3]
+8CNOT q[0],q[1]
+9RZ q[1],(1*2*Param([0,0]))
+10CNOT q[0],q[1]
+11CNOT q[1],q[2]
+12RZ q[2],(1*2*Param([0,0]))
+13CNOT q[1],q[2]
+14CNOT q[2],q[3]
+15RZ q[3],(1*2*Param([0,0]))
+16CNOT q[2],q[3]
+17CNOT q[3],q[0]
+18RZ q[0],(1*2*Param([0,0]))
+19CNOT q[3],q[0]
+20CNOT q[0],q[2]
+21RZ q[2],(0.5*2*Param([0,0]))
+22CNOT q[0],q[2]
+23RX q[0],(2*beta_0)
+24RX q[1],(2*beta_0)
+25RX q[2],(2*beta_0)
+26RX q[3],(2*beta_0)
+27CNOT q[0],q[1]
+28RZ q[1],(1*2*Param([1,0]))
+29CNOT q[0],q[1]
+30CNOT q[1],q[2]
+31RZ q[2],(1*2*Param([1,0]))
+32CNOT q[1],q[2]
+33CNOT q[2],q[3]
+34RZ q[3],(1*2*Param([1,0]))
+35CNOT q[2],q[3]
+36CNOT q[3],q[0]
+37RZ q[0],(1*2*Param([1,0]))
+38CNOT q[3],q[0]
+39CNOT q[0],q[2]
+40RZ q[2],(0.5*2*Param([1,0]))
+41CNOT q[0],q[2]
+42RX q[0],(2*beta_1)
+43RX q[1],(2*beta_1)
+44RX q[2],(2*beta_1)
+45RX q[3],(2*beta_1)
+46
+47Optimization successful! Expectation value: -0.3581
+48Optimal parameter values:
+49[[1.20130278 0.61696788]
+50 [1.97099634 2.90861432]]
+51
+52Top 5 solutions:
+53Bitstring&Cut Value&Probability
+54-----------------------------------
+551010 &4.0 &0.4126
+560101 &4.0 &0.4328
+571000 &2.5 &0.0050
+581100 &2.5 &0.0312
+590010 &2.5 &0.0054
+Listing 27: An example Python program that update and modify quantum states.
+Figure 32: Optimal Cut Result for 5000 Shots
+Figure 33: Optimal Cut Distribution for 5000 Shots
+The theoretical formula $H_{C}=\frac{1}{2}\sum_{(i,j)\in E}w_{ij}(I-Z_{i}Z_{j})$ is exactly equivalent to the code implementation.
+The theoretical formula can be decomposed into a constant term $\frac{1}{2}\sum w_{ij}I$ and a ZZ operator term $-\frac{1}{2}\sum w_{ij}Z_{i}Z_{j}$, where the constant term is reasonably omitted in quantum optimization because it only shifts the energy value globally without changing the relative relationships of the eigenstates.
+In the code, the ZZ term $\frac{1}{2}w_{ij}Z_{i}Z_{j}$ is implemented with a positive coefficient $0.5*w$, which is equivalent to the negative term $-\frac{1}{2}w_{ij}Z_{i}Z_{j}$ in the theoretical formula, because minimizing the expectation value of $\frac{1}{2}w_{ij}Z_{i}Z_{j}$ is equivalent to maximizing the theoretical objective.
+TABLE VI: Correspondence between Vertex States and Energy Contributions
+Vertex Relationship $Z_{i}Z_{j}$ Theoretical Cut Contribution Code Hamiltonian Contribution
+Same Side $+1$ 0 $+\frac{1}{2}w_{ij}$
+Opposite Sides $-1$ $w_{ij}$ $-\frac{1}{2}w_{ij}$
+Physically, essentially, when vertices are on opposite sides (cut contribution $w_{ij}$), $Z_{i}Z_{j}=-1$ causes the code term $\frac{1}{2}w_{ij}Z_{i}Z_{j}$ to become $-\frac{1}{2}w_{ij}$ (negative value); when on the same side (no contribution), it yields a positive value $\frac{1}{2}w_{ij}$.
+Therefore, minimizing the expectation value of the code Hamiltonian naturally drives the system toward the state with maximum cut contribution, which is completely consistent with the theoretical objective.
+The key code implementation is as follows:
+1for i, j, w in edges:
+2 ham_terms.append(("ZZ", [i, j], 0.5 * w)) # Positive coefficient achieves physical equivalence to the negative term
+Listing 28: An example Python program that update and modify quantum states.
+Using the QAOA algorithm, we successfully solved the maximum cut problem for a 4-vertex weighted graph featuring a ring structure (vertices 0-1-2-3-0) with an additional diagonal edge (0-2).
+The ring edges carried a weight of 1.0, while the diagonal edge had a weight of 0.5.
+Two independent experiments demonstrated that the algorithm consistently identified the theoretical maximum cut value of 4.0, corresponding to two optimal solutions: ‘1010‘ and ‘0101‘.
+These solutions represent the optimal partitioning scheme where vertices are divided into groups 0,2 and 1,3.
+Increasing the number of quantum measurements (shots) significantly enhanced result precision, raising the probability of obtaining the optimal solution from 29.7% to 82.58%, thereby verifying the probabilistic nature of quantum algorithms.
+TABLE VII: QAOA Experimental Results Comparison
+Evaluation Metric Shots=700 Shots=5000 Change Analysis
+Expectation Value -0.3581 -1.5075 More negative values indicate better optimization
+Optimal Parameters [[3.69 2.28],[0.69 1.32]] [[1.20 0.62],[1.97 2.91]] Multiple equivalent optima in parameter space
+Total Optimal Solution Probability 28.4% 84.54% Increased sampling enhances accuracy
+Suboptimal Solution Probability 40.0% 15% Probability distribution concentrates on optima
+Optimal Cut Value 4.0 4.0 Consistently achieves theoretical optimum
+As shown in Table VII, different parameter sets consistently converge to the same optimal solution, demonstrating the algorithm’s robustness.
+Increasing shots from 700 to 5000 elevates the probability of obtaining the optimal solution from 28.4% to 84.54% (row 3 of Table VII), consistent with the law of large numbers.
+The diversity in parameter space (row 2 of Table VII) reveals multiple equivalent optimal parameter combinations in QAOA, all leading to identical optimal solutions.
+These results fully validate QAOA’s effectiveness for combinatorial optimization problems, achieving the theoretical optimum at p=2 layers with over 80% probability and demonstrating quantum computing’s potential for optimization challenges.